### v1

# Don’t Pay Attention

Abstract
The Transformer has become the de facto standard for large language models and a wide range of downstream tasks across various domains. Despite its numerous advantages like inherent training parallelism, the Transformer still faces key challenges due to its inability to effectively process sequences beyond a fixed context window and the quadratic complexity of its attention mechanism. These challenges have renewed interest in RNN-like architectures, which offer linear scaling with sequence length and improved handling of long-range dependencies, albeit with limited parallelism due to their inherently recurrent nature. In this paper, we propose Avey, a new neural foundational architecture that breaks away from both attention and recurrence. Avey comprises a ranker and an autoregressive neural processor, which collaboratively identify and contextualize only the most relevant tokens for any given token, regardless of their positions in the sequence. Specifically, Avey decouples sequence length from context width, thus enabling effective processing of arbitrarily long sequences. Experimental results show that Avey compares favorably to the Transformer across a variety of standard short-range NLP benchmarks, while notably excelling at capturing long-range dependencies.
1 Introduction
The Transformer [98] has emerged as one of the most influential AI innovations in recent years, profoundly impacting various aspects of modern life, including work, science, and art, to mention just a few. Notably, Large Language Models (LLMs) are almost universally based on the Transformer [25], which has demonstrated remarkable performance in natural language processing (NLP) [66, 55, 75] and various other fields [32, 56, 2].
The Transformer’s state-of-the-art performance is largely driven by its recurrence-free self-attention mechanism, which enables parallel processing of entire token sequences. Nevertheless, the computational and memory costs of self-attention grow quadratically with sequence length, making it inefficient for handling arbitrarily long sequences. Extensive research has been conducted over the years to address this limitation [94], with a noticeable emphasis on linearizing attention [48, 9, 106, 99]. These linear approaches aim at approximating self-attention in a more computationally efficient manner, without considerably compromising performance.
Nonetheless, linear attention mechanisms have generally underperformed the original self-attention mechanism, often by a significant margin in language modeling tasks [102, 47]. While recent linear models such as RWKV [70] and RetNet [91] have shown promising results, substantial progress is still needed before they can reliably and consistently surpass the Transformer [52]. In addition, these models have yet to demonstrate definitive empirical effectiveness at scale [25]. This persistent performance gap between quadratic and linear approaches has spurred renewed interest in RNN-based architectures, which offer linear scalability with sequence length but limit parallelism due to their inherently cyclical nature.
To exemplify, state space models (SSMs) [44, 28], which are viewed as extensions of RNNs, have recently emerged as a compelling class of architectures. Unlike traditional RNNs, SSMs can parameterize their state transition matrices in a structured manner (e.g., via using a diagonal plus low-rank decomposition) to improve computational efficiency and enhance gradient flow. A specialized subclass of these models, known as structured state space sequence (S4) models [27, 28], has garnered growing attention. Yet, despite their theoretical appeal, S4 models struggled with language modeling tasks, typically trailing Transformers by several points [20, 27].
Most recently, Mamba [25] advanced S4 models by enhancing their selectivity and effectiveness while enabling high training concurrency. It demonstrated strong performance on tasks involving long-range dependencies and compared favorably to Transformers in language modeling. However, training, scaling, and interpreting Mamba—and SSMs more broadly [87, 72, 31]—remain challenging, while continue to be promising [13].
We posit that the primary limitation of the Transformer lies in its inability to effectively model dependencies beyond its fixed context window. While its core self-attention mechanism is inherently parallelizable, this constraint makes its quadratic complexity a significant bottleneck at scale. This explains the surge of research aimed at reducing this complexity or exploring RNN-inspired alternatives. In this work, we propose a more viable approach by decoupling context width from sequence length, allowing models to scale to arbitrarily long sequences. Under this paradigm shift, the quadratic training complexity becomes less of a concern when small context windows are used, especially if the models maintain high parallelizability.
This paper introduces Avey [^1] Avey is not an acronym, but a name that the authors like., a new architecture for language modeling that entirely departs from Transformer-based and RNN-like designs. Avey is a flexible, sequence-length-invariant model that decouples sequence length from context width, thus enabling effective processing of long-range sequences. It preserves the influence of tokens that appear outside its context window, regardless of their positions in the sequence. This is achieved via a weighted-selective-split interaction mechanism, which systematically skips irrelevant tokens beyond the context window and ensures direct interactions with only relevant ones, retaining their contributions irrespective of sequence length.
Figure 1: Needle-in-a-Haystack test performance comparison between Transformer++, Mamba, RWKV-7, and Avey, all using 1.5B parameters. The x-axis denotes the lengths of haystacks (i.e., documents with distractor texts, varying from 2k to 64k tokens) and the y-axis refers to the position of the needle (i.e., a short sentence) within any of the haystacks. A green cell means that the model was able to recall the needle and a red one entails that it could not. Transformer++, Mamba, and RWKV-7 were all trained with 2k-token contexts, while Avey was trained with only a 512-token context and was able to extrapolate to the maximum needed modeling capacity.
Fig. 1 illustrates Avey’s ability to generalize beyond its training context. A popular benchmark for evaluating this capability is Needle-in-a-Haystack (NiaH) [45]. This benchmark measures a model’s capacity to recite a specific sentence (i.e., the needle) placed at an arbitrary position within a large body of distractor text (i.e., the haystack). Since its introduction, NiaH has become a widely used sandbox for probing the limits of long-context language models in capturing distant dependencies, and smaller models in generalizing beyond their trained context windows [22]. As shown in the figure, Transformer++ (i.e., the Transformer with an enhanced architecture and training recipe– see Section 3.1), which was trained with a 2k-token context window, could not generalize beyond that limit. In contrast, Mamba and RWKV-7 [70], also trained with 2k-token context windows, managed to generalize to nearly 8k and 16k tokens, respectively. Most notably, Avey, despite being trained on a context window of only 512 tokens, successfully generalized to the maximum tested sequence length of 64k tokens, demonstrating strong extrapolative capability far beyond its original training regime.
To elaborate on its technical aspects, Avey is a recurrence- and attention-free architecture comprising two principal components, a ranker and a neural processor. The ranker slices each input sequence into splits of consecutive tokens and selects the top k most relevant splits for each current split being processed by the neural processor. The neural processor consists of three core units, the enricher, contextualizer, and fuser. The enricher enhances the quality of token embeddings by expanding their learnable features using a position-wise neural network. The contextualizer is an embedding-wise neural network with dynamic parameterization, enabling interactions between relevant tokens across the current and top k splits. Lastly, the fuser learns a function that integrates the contextualized features produced by the contextualizer with some uncontextualized features bypassed by a partial-embedding bypassing mechanism.
To summarize, our main contributions in this paper are as follows:
  • We propose Avey, a new recurrence- and attention-free neural architecture that decouples context window from sequence length, thus enabling effective processing of long-range sequences.
  • We show that Avey performs comparably to the Transformer—outperforming it at two model sizes and underperforming it at one—across a range of popular zero-shot NLP benchmarks, thereby establishing an initial foundational architecture with potential for more scalable and effective language modeling.
  • In contrast to the Transformer, we demonstrate that Avey can scale far beyond its context window using the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER [38].
  • We show that Mamba (representing SSMs) and RWKV-7 (representing linear attention models) exhibit some ability to generalize beyond their training context windows, but their performance decline significantly as the sequence length increases far beyond them. By comparison, Avey consistently and substantially outperforms both Mamba and RWKV-7 on the S-NIAH benchmark suite.
  • We conduct extensive ablation studies to assess the impact of each design choice in Avey.
  • We provide a comprehensive survey of related work in Appendix O.
  • We open-source the code and pretrained checkpoints of Avey to facilitate reproducibility and foster future research [^2] https://github.avey.ai/avey-dpa..
2 Avey
As indicated earlier, Avey comprises two components, a ranker and an autoregressive neural processor. Below, we provide a detailed explanation of each component.
2.1 Ranker
Figure 2: The ranker (left) partitions each input sequence into equal-sized splits and identifies the top $k$ most relevant ones (e.g., splits 1 and 3 for $k=2$) with respect to the current split (e.g., split 4), using the MaxSim operator. These top-$k$ splits are then weighted by their normalized scores, where the normalized score (NS) of a split is computed as the ratio of its MaxSim value to the highest MaxSim score among the $k$ splits. Finally, the weighted top-$k$ splits are contextualized together with the current split by the neural processor (right).
Avey decouples sequence length from context width, enabling the processing of arbitrarily long sequences. The sequence length refers to the total number of tokens in a sequence, while the context width denotes the number of tokens that the neural processor can contextualize simultaneously. Importantly, the sequence length can be set to a value that is much larger than the context width. As such, the influence of global tokens (or tokens that fall outside the context window) may diminish as more tokens are successively processed. If such global tokens are semantically relevant to local tokens (or tokens that fall within the context window), the quality of token representations will decline, and the effectiveness of the model will degrade.
To this end, the ranker and neural processor jointly employ a weighted-selective-split interaction mechanism, which skips irrelevant global tokens and ensures direct interactions with relevant ones, preserving their impact regardless of sequence length. As demonstrated in Fig. 2, Avey divides each input sequence into equal-sized splits, each consisting of a list of contiguous token embeddings. Prior to predicting the next token in the sequence (e.g., token 9 in the figure), Avey involves the ranker to identify the top k (e.g., 2 in the figure) splits that are most relevant (e.g., splits 1 and 3) to the current split (i.e., split 4).
The current split is defined as the one that either contains the token to be predicted (e.g., split 4 may contain only embedding 7, and Avey will aim to predict token 8) or contributes to predicting the first token in its subsequent split (e.g., split 4 serves in predicting token 9, which will belong to split 5 once predicted). To determine relevance, the ranker computes a similarity score between the current split, say $S_{c}$, and each preceding split, say $S_{p}$, using the MaxSim operator [49], originally proposed and utilized in Information Retrieval. Specifically, pairwise similarities are calculated (e.g., using a cosine function) between each embedding in $S_{c}$ and all embeddings in $S_{p}$. For each embedding in $S_{c}$, the maximum similarity across all $S_{p}$’s embeddings is taken, and then the maxima of all $S_{c}$’s embeddings are added to yield the final MaxSim score (see Fig. 2). This score signifies how relevant $S_{p}$ is to $S_{c}$.
Subsequently, the preceding splits are ranked based on their MaxSim scores, and the top k most relevant ones are contextualized with the current split by the neural processor, while maintaining their original order in the sequence. Before contextualization, however, the MaxSim scores of the top k splits are normalized with respect to the highest MaxSim score among them (e.g., split 3’s MaxSim score of 0.8 in Fig. 2 is normalized via dividing it by the maximum score among the top k splits, i.e., 1.6, yielding 0.5). Each selected split is then weighted by its corresponding normalized MaxSim score, effectively scaling its contribution during contextualization.
As a result, the weighted-selective-split interaction mechanism does not only allow the ranker to rank splits based on relevance but also the neural processor to contextualize them accordingly, as each selected split is pre-weighted by its relevance score. This empowers the neural processor to judiciously leverage global information (i.e., splits beyond the context width) by focusing selectively on only relevant features, emphasizing informative ones and deemphasizing less useful ones, thus enhancing performance. We analyze the impact of weighting the top k splits in Appendix K.
Note that the ranker is invoked only once per full forward and backward pass. To elucidate, Avey’s depth can be increased by stacking multiple layers within the neural processor (see Fig. 3), thereby enabling the modeling of complex, hierarchical patterns. In contrast, only a single ranker is required before the stack of layers within the processor, regardless of their number. Once the ranker identifies the top k relevant splits of the current split, the neural processor contextualizes them all using one or more layers.
Consequently, during training, each current split is matched once against every preceding split. This results in a compute cost of $\frac{N/S\cdot(N/S+1)}{2}\cdot S^{2}d$ or a time complexity of $O(N^{2}d)$, where $N$ is the sequence length, $S$ is the split size, and $d$ is the embedding dimension. This complexity assumes that scalar multiply-add operations (e.g., those used in computing cosine similarity for MaxSim) and comparisons (e.g., those utilized to determine maximum scores) are constant-time. We next discuss the neural processor.
2.2 Neural Processor
The neural processor encompasses three key machineries, the enricher, contextualizer, and fuser (see Fig. 3). We describe each in detail below.
2.2.1 The Enricher
Figure 3: The neural processor (top) with its three major components, the enricher, contextualizer (Cx), and fuser. The processor is unfolded into two copies for illustrative purposes only, to show how different embeddings, (e.g., $e_{1}$ and $e_{2}$, or more precisely, parts of their tails, i.e., $e_{122}$ and $e_{222}$) are contextualized by Cx (i.e., in reality, all components are shared across all embeddings and many embeddings can be input to Cx simultaneously).
The enricher aims at enriching the quality of each token representation via expanding the quantity of its learnable features, thereby enabling the contextualizer to capture more nuanced distinctions between tokens. Concretely, it is a one-layer, position-wise neural network (i.e., the input to each neuron is a single scalar element from an embedding), thus operating on each embedding independently, without considering neighboring embeddings. As such, it allows intra-feature interactions within the context of each individual embedding, facilitating the learning of higher-order and more expressive representations. The enricher can expand each input embedding by an arbitrary factor. We study the effect of varying the expansion factor on Avey’s performance in Appendix D, and ablate the enricher’s contribution in Appendix K.
Equation 1 formalizes the enricher, where $\mathbf{X}\in\mathbb{R}^{C\times d}$ is a matrix of $C$ input embeddings ($C\leq N$, where $N$ is the sequence length), each of dimension $d$; $\sigma$ is an activation function; $\mathbf{U}\in\mathbb{R}^{d\times m}$ is a learnable weight matrix defining a linear projection from dimension $d$ to $m$, where $m>d$; and $\mathbf{b}\in\mathbb{R}^{C\times m}$ denotes biases.
$\mathbf{Z}=\sigma(\mathbf{X}\mathbf{U}+\mathbf{b})$ (1)
As demonstrated in Fig. 3, the enricher feeds both, the contextualizer and the fuser. In particular, it bypasses a portion of each expanded embedding directly to the fuser in a technique that we refer to as partial-embedding bypassing. More precisely, the output of the enricher, $\mathbf{Z}\in\mathbb{R}^{C\times m}$, is split into two parts: (1) the head $\mathbf{Z}_{h}\in\mathbb{R}^{C\times m_{h}}$, which is bypassed directly to the fuser, and (2) the tail $\mathbf{Z}_{t}\in\mathbb{R}^{C\times m_{t}}$, which is forwarded to the contextualizer, where $m=m_{h}+m_{t}$. Consequently, varying the tail size alters the head size, which can influence Avey’s performance. We investigate the impact of different tail sizes on Avey’s performance in Appendix E.
The partial-embedding bypassing technique allows preserving raw distinctive features of each embedding, thus inducing representations with inherent diversity. This diversity may serve in alleviating issues like entropy collapse [107], where the contextualizer increasingly focuses on a few tokens, and over-smoothing [109, 85, 110], where embeddings become increasingly similar, as Avey’s depth is increased. We analyze the significance of partial-embedding bypassing on Avey’s effectiveness in Appendix K.
Lastly, Equation 1 implies that each neuron performs a weighted sum of $d$ input features (i.e., the elements of an embedding), incurring $d-1$ multiply-add operations. Since $d$ is projected to a higher dimension $m$ [^3] In our case, we experiment with $m$ being a multiple of $d$, entailing that $m\geq 2d$ (see Appendix D)., the total computational cost is $m(d-1)$ per token. For a sequence of $N$ tokens, the cost becomes $Nm(d-1)$, or asymptotically $\mathcal{O}(Nmd)$.
2.2.2 The Contextualizer
The contextualizer is a one-layer, embedding-wise neural network (i.e., the input to each neuron is one embedding), thus operating in parallel on $C$ embeddings, where $C$ denotes the context width. More precisely, it enables inter-embedding, data-dependent interactions of only tail embeddings (i.e., $\mathbf{Z}_{t}\in\mathbb{R}^{C\times m_{t}}$, as defined in Section 2.2.1), after each enricher’s output embedding $m$ is split into a head part (i.e., $m_{h}$) and a tail part (i.e., $m_{t}$), and only the $m_{t}$ part (e.g., $e_{12}$ and $e_{22}$ in Fig. 3) is forwarded to the contextualizer.
The $m_{t}$ part of each enriched embedding is further divided into two equal portions, $m_{tl}$ (or left portion) and $m_{tr}$ (or right portion), to enable judicious control of information flow through the neural processor. Specifically, $m_{tl}$ serves as a gating mechanism for $m_{tr}$, regulating how much of its contextualized feature values are propagated forward. Both $m_{tl}$ and $m_{tr}$ are learnable by the model, hence, allowing $m_{tl}$ to dynamically capture the significance of each $m_{tr}$’s feature, and emphasize or deemphasize its influence accordingly. This gating mechanism was inspired from gMLP [53] and resembles that of Gated Linear Units [15, 84, 100].
More formally, $\mathbf{Z}_{t}\in\mathbb{R}^{C\times m_{t}}$ is partitioned into two equal parts, $\mathbf{Z}_{tl}\in\mathbb{R}^{C\times\left(m_{t}/2\right)}$, which is bypassed to a multiplicative element-wise operation as part of a gating mechanism, and $\mathbf{Z}_{tr}\in\mathbb{R}^{C\times\left(m_{t}/2\right)}$, which is contextualized via a neural network, where each neuron takes as input an embedding of dimension $m_{t}/2$. Equation 2 defines the overall process, where $\mathbf{V}\in\mathbb{R}^{C\times C}$ is a learnable weight matrix representing a linear cross-embedding transformation, $\odot$ denotes element-wise multiplication, $\mathbf{b}^{\prime}\in\mathbb{R}^{C\times(m_{t}/2)}$ refers to optional biases, and $\mathcal{N}(\mathbf{Z}_{tr})$ and $\mathcal{N}(\mathbf{Z}_{tr}^{\top})$ are row- and column-wise normalized versions of $\mathbf{Z}_{tr}$, respectively.
$\mathbf{c}(\mathbf{Z}_{t})=\mathbf{Z}_{tl}\odot\sigma\left(\left(\mathbf{V}\odot\mathcal{N}(\mathbf{Z}_{tr})\mathcal{N}(\mathbf{Z}_{tr}^{\top})\right)\mathbf{Z}_{tr}+\mathbf{b}^{\prime}\right)$ (2)
Equation 2 suggests that each neuron in the contextualizer’s network performs a weighted sum of the cosine similarities between embeddings (denoted by $\mathcal{N}(\mathbf{Z}_{tr})\mathcal{N}(\mathbf{Z}_{tr}^{\top})$) and the embeddings themselves (denoted by $\mathbf{Z}_{tr}$). This introduces a level of selectivity into the neural processor, as described and advocated by [25]. Specifically, it makes the parametrization of the neural processor dynamic, enabling it to disregard or focus on information during inference based on the input. We examine the influence of dynamic parametrization on Avey’s performance in Appendix K.
Finally, we note that the contextualizer inherently models the relationships between tokens, making the neural processor naturally aware of their positions in the sequence (i.e., positional encodings are not needed). In terms of complexity, as each neuron performs a weighted sum of $C$ embeddings, each of dimension $m_{t}/2$, it results in a cost of $(C-1)m_{t}/2$ multiply-add operations. With $C$ neurons, the cost becomes $C(C-1)m_{t}/2$. For a sequence of $N$ tokens, the contextualizer processes $N/S$ splits, each contextualized with $k$ relevant splits, making $C=S(k+1)$ and yielding a total cost of $(N/S)[C(C-1)m_{t}/2]=N(k+1)[(C-1)m_{t}/2]$ (after substituting $S$ with $C/(k+1)$), or asymptotically $\mathcal{O}(NkCm_{t})$.
2.2.3 The Fuser
The fuser is designed to learn an optimal function, referred to as fusion [^4] The name is inspired from the CNN literature [40]., that integrates uncontextualized features (i.e., those of dimension $m_{h}$, bypassed by the partial-embedding bypassing technique presented in Section 2.2.1) with contextualized features (i.e., those of dimension $m_{t}/2$, output by the contextualizer). Subsequently, it produces, for each input token, a contracted representation that matches the token’s original embedding dimension $d$ (see Fig. 3). Akin to the enricher, it is a one-layer, position-wise neural network, which operates on each embedding of dimension $m_{h}+m_{t}/2$ independently.
[…]
$f(\mathbf{Z})=[\mathbf{Z}_{h}\,\|\,\mathbf{c}(\mathbf{Z}_{t})]\mathbf{O}$ (3)
Equation 3 entails that each neuron performs a weighted sum of $m_{h}+m_{t}/2$ embedding elements, yielding a cost of $(m_{h}+m_{t}/2-1)$ multiply-add operations. For $d$ neurons (since the fuser projects $m_{h}+m_{t}/2$ to $d$), the cost is $d(m_{h}+m_{t}/2-1)$. For a sequence of $N$ tokens, the total cost is $Nd(m_{h}+m_{t}/2-1)$, or asymptotically $\mathcal{O}(Nmd)$.
Considering the aggregate computational costs of the ranker, enricher, contextualizer, and fuser, Avey exhibits a training time complexity of $\mathcal{O}(L(2Nmd+NkCm_{t})+N^{2}d)$, where $L$ denotes the number of neural processor layers. As the term $N^{2}d$ dominates asymptotically, the overall complexity simplifies to $\mathcal{O}(N^{2}d)$. During inference, the complexity reduces to $\mathcal{O}(N)$, or linear per token. We elaborate on Avey’s inference time complexity in Appendix N.
Figure 4: The Time to First Token (TTFT) for Avey, Transformer++, Mamba, and RWKV-7 using $\sim$1.5B parameters, across varying sequence lengths. See Section 3.1 for details on the experimental methodology.
To characterize Avey’s inference-time efficiency relative to other models, we benchmarked Time to First Token (TTFT)—a key latency metric for real-time applications [36, 54, 17]—on a single NVIDIA H200 GPU. The evaluation included Avey, Transformer++, Mamba, and RWKV-7 across a range of sequence lengths. As shown in Figure 4, Transformer++ demonstrates an approximately quadratic increase in TTFT as the sequence length $N$ grows, owing to its full self-attention mechanism, which must operate over the entire prompt before generating the first token. In contrast, Mamba and RWKV-7 scale linearly with $N$, as they require a full forward pass to construct their RNN-style hidden states prior to emitting the first token. While Avey is also theoretically expected to scale linearly, its empirical TTFT is significantly lower than that of Transformer++, Mamba, and RWKV-7. This efficiency gain stems from Avey’s architectural design, which invokes its dominant contributor to inference complexity (see Appendix N), namely, the ranker, only once per forward pass. As a result, the ranker imposes minimal computational overhead in practice, enabling Avey to achieve substantially lower TTFT and making it particularly well-suited for latency-sensitive, real-world applications (e.g., conversational AI and edge deployments).
3 Experiments
3.1 Experimental Setup
We compare Avey against three leading open-source models, namely, Mamba [42], RWKV-7 [43], and Transformer++, extended to the strongest known architectural recipe of the standard Transformer [46] (see Appendix A for details). All models were trained using their best-known hyperparameters under a fixed budget of 100 billion tokens drawn from the FineWeb dataset [41]. Complete training and model hyperparameters for all the baselines are provided in Appendix A.
To assess each model’s accuracy, we employed a suite of widely used NLP benchmarks, including ARC-E and ARC-C [11], HellaSwag [105], PIQA [5], OBQA [61], SIQA [82], and Winogrande [80]. Additionally, we evaluated the long-context retrieval capabilities of all the models using the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER [38]. Full details of all the benchmarks and additional experimental setups are included in Appendix A.
3.2 Design Choices and Ablations
Table 1: Summary of studies for key design choices and corresponding experimental references.
Question Answer Experiments
RMSNorm or LayerNorm? RMSNorm Appendix H
LR: to decay or not to decay? Yes, cosine decay with peak LR of $1\text{e}{-3}$ Appendix I
Best values for sequence length $N$, split size $S$, and top-$k$ splits? $N=512$, $S=64$, $k=7$ Appendix G
Activation in the enricher? Yes, ReLU2 Appendix B
Activation in the contextualizer? No Appendix C
Deeper model and narrower embeddings, or shallower model and wider embeddings? Deeper model and narrower embeddings Appendix F
Weight ranked splits? Yes, using normalized scores Appendix K
Enrich embeddings before contextualization? Yes, by 4x Appendix K, D
Bypass uncontextualized features to the fuser? Yes, 50% of each enriched embedding Appendix K, E
Static or dynamic parametrization for the contextualizer? Dynamic parametrization Appendix K
Replace the contextualizer with self-attention? No Appendix K
Besides enabling extrapolation past the context window, does the ranker improve performance? Yes Appendix K
We conducted over 200 experiments to investigate and ablate several key design choices in our architecture. These included, whether to: (1) utilize activation functions in the enricher and contextualizer; (2) trade off additional layers for reduced embedding dimensionality (or vice versa) under a fixed parameter budget; (3) apply weights to ranked splits; (4) enrich embeddings; (5) bypass uncontextualized features to the fuser; (6) use dynamic parameterization in the neural processor; and (7) replace the neural contextualizer with self-attention. Furthermore, we evaluated Avey without the ranker to assess its contribution to downstream task performance, beyond its primary role of enabling effective extrapolation past the trained context window. Table 1 summarizes our findings and provides references to the corresponding experiments that support each conclusion.
3.3 Short-Range Benchmark Results
Table 2: Zero-shot performance across multiple NLP tasks.
Model ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
Avey-153M 24.37 42.33 39.36 31.40 68.37 39.13 51.28 42.32
Transformer++-152M 23.63 43.17 39.32 29.80 67.01 38.89 50.22 41.72
Mamba-144M 24.17 43.53 40.55 30.40 68.32 39.41 52.72 42.73
RWKV-7-168M 24.17 43.01 41.55 29.67 68.72 39.17 51.09 42.48
Avey-496M 27.50 48.95 51.82 32.47 72.49 40.15 54.38 46.82
Transformer++-488M 26.73 48.09 52.66 31.73 72.13 39.93 55.25 46.65
Mamba-500M 28.64 51.02 54.15 34.47 73.03 40.84 55.49 48.23
RWKV-7-501M 27.13 49.37 54.54 36.27 73.58 39.40 55.72 48.00
Avey-1.52B 31.26 56.55 61.42 36.80 75.61 42.00 57.06 51.53
Transformer++-1.5B 30.00 56.29 63.87 38.00 76.01 42.24 61.38 52.54
Mamba-1.4B 32.13 57.74 63.74 36.85 76.19 42.00 60.40 52.72
RWKV-7-1.5B 32.94 59.05 64.43 37.13 76.84 41.71 60.06 53.17
We now evaluate Avey on standard autoregressive language modeling benchmarks, comparing it against Transformer++, Mamba, and RWKV-7 across three model sizes, small, medium, and large, as defined in Section 3.1. We utilize a range of widely used zero-shot downstream evaluation tasks, all detailed in Section 3.1. Table 2 summarizes the results. With small models, Avey, Mamba, and RWKV-7 outperformed Transformer++ by average margins of 1.43%, 2.41%, and 1.82%, respectively. Mamba and RWKV-7 slightly exceeded Avey’s performance, with average margins of 0.9% and 0.3%, respectively. With medium models, Avey, Mamba, and RWKV-7 again outperformed Transformer++ by averages of 0.3%, 3.4%, and 2.9%, respectively. Lastly, with large models, Avey underperformed Transformer++ by an average of 1.9%, while Mamba and RWKV-7 marginally outpaced it by 0.71% and 1.19%, respectively.
The results above assume a fixed training budget of 100B tokens. To better understand how Avey scales with increasing model size, we conducted additional experiments following the Chinchilla scaling laws [35], which recommend increasing the number of training tokens proportionally with model size. Consequently, we adjusted the number of training steps and tokens to align with these laws. Appendix J outlines the configurations of the trained models, including the numbers of layers, embedding dimensions, training steps, learning rates, and total training tokens. The methodology of these experiments closely follows that of [25], with slight modifications (e.g., to accommodate parameter budget constraints). As demonstrated in Appendix J, Avey exhibits more favorable scaling behavior than Transformer++, Mamba, and RWKV-7, particularly when both model size and token count are increased proportionally in accordance with the Chinchilla scaling laws.
Finally, to provide deeper insights into the behaviors of the compared models, we performed a detailed statistical analysis, presented in Appendix L. The analysis reports the standard deviation, standard error, and 95% confidence interval for each model, computed over its final three checkpoints—captured at 5-billion-token intervals (i.e., at 90B, 95B, and 100B tokens). The results revealed meaningful variance among models. For example, in the small model regime, although Avey does not surpass Mamba in mean performance, their confidence intervals overlap substantially within the range (42.06, 43.24), indicating that the two models are statistically comparable and that Avey may outperform Mamba in some runs. A similar, though narrower, overlap is observed between Avey and RWKV-7. In the large model regime, while Avey does not exceed Transformer++ in average performance, their confidence intervals also exhibit significant overlap, suggesting that Avey could potentially outperform Transformer++ in certain cases. A more limited overlap is seen between Avey and Mamba, implying that although Mamba generally performs better, Avey may outperform it in some runs.
3.4 Long-Range Benchmark Results
Figure 5: Performance comparison between Transformer++, Mamba, RWKV-7, and Avey on S-NIAH-1 and S-NIAH-2. The x-axis denotes the lengths of haystacks (i.e., documents with distractor texts, varying from 2k to 64k tokens). All models use 0.5B parameters. Similar results are shown in Appendix M for other model sizes.
In this section, we evaluate Avey, Transformer++, Mamba, and RWKV-7 on benchmarks designed to assess performance on tasks with long-range dependencies. Specifically, we use the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER[38], as described in Section3.1. The S-NIAH suite includes multiple variants, notably S-NIAH-1 (pass-key retrieval) and S-NIAH-2 (number in haystack). S-NIAH-1 involves retrieving the specific value associated with a given key (the pass-key) from a lengthy distractor text (the haystack) containing many key-value pairs. The pass-key, serving as the needle, appears only once, and the model must accurately recall its corresponding value regardless of its position in the haystack. S-NIAH-2 is similar to S-NIAH-1 but poses a greater challenge, whereby the value to be retrieved is numerical (e.g., a random 9-digit number). This task requires exact recall, where even a single-digit error is considered incorrect, thereby testing the model’s precision in extracting structured numerical information from long haystacks.
Fig. 5 demonstrates the results of Avey, Transformer++, Mamba, and RWKV-7 on both S-NIAH-1 and S-NIAH-2 benchmarks. As described in Section 3.1, Transformer++, Mamba, and RWKV-7 were all trained with a context window of 2,048 tokens. As shown, Transformer++ performs strongly on both benchmarks as long as the haystack length remains within its trained context window. Once the haystack’s length exceeds its context window width, Transformer++ fails to recall the correct values associated with the keys. In contrast, Mamba and RWKV-7 demonstrate some ability to generalize beyond their training windows, but their performance also declines significantly as the haystack length increases far beyond those limits. On the flip side, Avey achieves good performance across both benchmarks, despite being trained with a context window of only 512 tokens. For instance, on S-NIAH-2 with a 64k-token haystack, Avey outperforms Mamba and RWKV-7 by averages of 85.25% and 23.6%, respectively. In addition, on S-NIAH-1 under the same 64k-token setting, Avey achieves an accuracy of 97.8%, while Mamba and RWKV-7 drop to 0% and 0.8%, respectively.
Interestingly, Avey’s performance tends to improve as the haystack length increases, highlighting its strong extrapolative capability. This behavior can be attributed to the fact that as the haystack length (i.e., sequence length $N$) grows, the candidate pool from which the ranker selects the top-$k$ splits for contextualization also expands. As discussed in Appendix G, a larger $N$ enables the ranker to identify and retrieve more relevant splits while discarding less informative ones, thereby improving the overall quality of contextualization and potentially enhancing performance. This effect is further supported by the results in Appendix K, where the inclusion of the ranker led to measurable performance gains. Notably, embeddings containing a needle—whether in S-NIAH-1 or S-NIAH-2—are not processed in isolation but rather contextualized alongside other embeddings. As such, an improved quality of contextualization driven by the ranker may contribute to more accurate value recall. However, whether this mechanism fully explains Avey’s increasing performance with longer haystacks remains uncertain, and further interpretability studies are needed to better understand the underlying drivers of this behavior.
4 Related Work
Extensive research has focused on addressing the limitations of the Transformer and exploring alternative RNN-like architectures. Appendix O provides a comprehensive survey of these efforts and offers a comparison with Avey.
5 Conclusion
In this paper, we introduced Avey, a new foundational architecture for autoregressive language modeling. Unlike traditional models, Avey relies neither on recurrence nor attention. Instead, it employs a neural approach to enrich and contextualize embeddings. Additionally, it leverages a ranker that enables the model to flexibly and effectively handle sequences of arbitrary lengths, despite being trained with only a small context window. We hope this work lays the groundwork for future research and inspires further advances in scalable and effective language modeling.
Limitations
We define the scope of our work and discuss limitations in Appendix P.
Appendix A Experimental Methodology
In this section, we describe the experimental methodology employed throughout the paper. To begin with, we adopted a cascaded search process to identify the best configuration of Avey. Specifically, we began with a baseline version of the neural processor—excluding the ranker—and sequentially explored several architectural design choices. After each empirical conclusion regarding a specific architectural element, we integrated that element into the processor (one at a time) and resumed the search process from the updated configuration. This cascaded process is captured chronologically in Appendices B, C, D, E, and F.
To elaborate, we started with an expansion factor of 4×4$\times$4 × in the enricher, a tail size of 50% (i.e., half of each expanded embedding is forwarded to the contextualizer), RMSNorm [108] as a normalization technique, no activation functions in the enricher and contextualizer, global batch size of 0.5M, context width of 1024, and a constant learning rate of $1e{-3}$. As we empirically verified and decided upon each architectural element, we updated the processor accordingly. For example, after determining that ReLU2 was the most effective activation function for the enricher, we integrated it into the model and proceeded with the remaining search.
After finalizing the above exploratory set of experiments, we incorporated the ranker into the neural processor and conducted an extensive study—comprising over 138 training and inference runs—to identify the optimal sequence length (i.e., $N$), split size (i.e., $S$), and number of top-$k$ splits (i.e., $k$). The results of this sensitivity study are summarized in Appendix G. Following this, we evaluated the best normalization technique (Appendix H) as well as the optimal peak learning rate and learning rate schedule (Appendix I) for the full architecture.
All the experiments described above were conducted using a 145-million-parameter model trained on 10 billion tokens from the FineWeb dataset [^6] This dataset is released under the Open Data Commons Attribution License (ODC-By) v1.0. [41] (specifically, the sample-100BT subset of FineWeb). The results of these experiments informed the following final selection of training and model hyperparameters for Avey across three parameter scales, 153M (small), 496M (medium), and 1.52B (large).
Training and Model Hyperparameters of Avey:
  • Training hyperparameters:
    • Optimizer: AdamW
    • Betas: (0.9, 0.95)
    • Epsilon: $1\text{e}{-12}$
    • Peak learning rate: $1\text{e}{-3}$
    • Schedule: Cosine decay to 10% of the peak learning rate, with no warmup
    • Batch size: 0.5M for the small and medium models, and 1M for the large model
    • Gradient norm clip: 1.0
    • Weight decay: 0.1 (applied only to matrices)
  • Model hyperparameters:
    • All models:
      • Context width: 512
      • Split size ($S$): 64
      • Number of top-$k$ splits: 7
      • Vocabulary size: 50,304
      • Expansion factor: 4
      • Tail size: 0.5
    • Small model (153M parameters):
      • Embedding dimension: 768
      • Number of layers: 26
    • Medium model (496M parameters):
      • Embedding dimension: 768
      • Number of layers: 104
    • Large model (1.52B parameters):
      • Embedding dimension: 2048
      • Number of layers: 48
For baselines, we compared Avey against three leading open-source models, namely, Mamba [42], RWKV-7 [43], and Transformer++ [46]. For Transformer++, we implemented the strongest architectural recipe known to us, incorporating rotary positional encodings [89], SwiGLU MLPs [84], and RMSNorm in place of LayerNorm [108]. All models were trained using their best-known hyperparameters (see details below, assuming three model sizes, small, medium, and large) under a fixed budget of 100 billion tokens drawn from the aforementioned FineWeb dataset  [^7] More precisely, all models, including Avey, Transformer++, Mamba, and RWKV-7 were trained for 1 epoch over the sample-100BT subset of FineWeb.. For consistency and comparability, we used the p50k_base tokenizer [64] across all the models, as it aligns with the GPT-2-derived token counts reported for this dataset.
Training and Model Hyperparameters of Transformer++:
  • Training hyperparameters:
    • Optimizer: AdamW
    • Betas: (0.9, 0.95)
    • Epsilon: $1\text{e}{-12}$
    • Peak learning rates:
      • Small model: $3\text{e}{-3}$
      • Medium model: $1.5\text{e}{-3}$
      • Large model: $1.25\text{e}{-3}$
    • Schedule: A linear warmup for 10% of steps, followed by cosine decay to 10% of the peak learning rate
    • Batch size: 0.5M for the small and medium models, and 1M for the large model
    • Gradient norm clip: 1.0
    • Weight decay: 0.1 (only applied to matrices)
  • Model hyperparameters:
    • All models:
      • Context width: 2048
      • Vocabulary size: 50,304
      • Intermediate size in FFN: 4$\times$ the embedding dimension
    • Small model (152M parameters):
      • Embedding dimension: 768
      • Number of layers: 12
      • Number of heads: 12
    • Medium model (488M parameters):
      • Embedding dimension: 1024
      • Number of layers: 26
      • Number of heads: 16
    • Large model (1.5B parameters):
      • Embedding dimension: 1664
      • Number of layers: 32
      • Number of heads: 16
Training and Model Hyperparameters of Mamba:
  • Training hyperparameters:
    • Optimizer: AdamW
    • Betas: (0.9, 0.95)
    • Epsilon: $1\text{e}{-12}$
    • Peak learning rates:
      • Small model: $3\text{e}{-3}$
      • Medium model: $1.5\text{e}{-3}$
      • Large model: $1.0\text{e}{-3}$
    • Schedule: A linear warmup for 10% of steps, followed by cosine decay to 10% of the peak learning rate
    • Batch size: 0.5M for the small and medium models, and 1M for the large model
    • Gradient norm clip: 1.0
    • Weight decay: 0.1 (applied only to matrices)
  • Model hyperparameters:
    • All models:
      • Context width: 2048
      • Vocabulary size: 50,304
      • All hyperparameters other than the ones specified are left at their default values according to [42]
    • Small model (144M parameters):
      • Embedding dimension: 768
      • Number of layers: 28
    • Medium model (500M parameters):
      • Embedding dimension: 1280
      • Number of layers: 42
    • Large model (1.4B parameters):
      • Embedding dimension: 2048
      • Number of layers: 52
Training and Model Hyperparameters of RWKV-7:
  • Training hyperparameters:
    • Optimizer: AdamW
    • Betas: (0.9, 0.95)
    • Epsilon: $1\text{e}{-12}$
    • Peak learning rates:
      • Small model: $6\text{e}{-4}$
      • Medium model: $4\text{e}{-4}$
      • Large model: $4\text{e}{-4}$
    • Schedule: Cosine decay to 10% of the peak learning rate, with no warmup
    • Batch size: 1M for the small and medium models, and 2M for the large model
    • Gradient norm clip: 1.0
    • Weight decay: 0.1 (applied only to matrices)
  • Model hyperparameters:
    • All models:
      • Context width: 2048
      • Vocabulary size: 50,304
      • All hyperparameters other than the ones specified are left at their default values according to [43]
    • Small model (168M parameters):
      • Embedding dimension: 768
      • Number of layers: 12
    • Medium model (501M parameters):
      • Embedding dimension: 1024
      • Number of layers: 30
    • Large model (1.5B parameters):
      • Embedding dimension: 2048
      • Number of layers: 24
To compare all models, we employed a suite of widely used NLP benchmarks, including ARC-E and ARC-C (for scientific reasoning and reading comprehension) [11], HellaSwag (for commonsense inference) [105], PIQA (for physical reasoning) [5], OBQA (for open-book science reasoning) [61], SIQA (for social interaction understanding) [82], and Winogrande (for coreference and commonsense reasoning) [80]. In addition, we evaluated long-context retrieval capabilities using the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER [38], which measures a model’s ability to extract pass-keys from large distractor corpora, with sequence lengths ranging from 2k to 64k tokens. All evaluations were conducted using the widely adopted LM Evaluation Harness from EleutherAI [23], consistent with the prior work in the field.
For all models, we reported performance in terms of benchmark accuracy [^8] Additionally, throughout the paper, all reported perplexity values specifically refer to training perplexity.. Specifically, we used normalized accuracy (acc-norm) from the LM Evaluation Harness whenever available. For each model, the reported score on each NLP benchmark is the average accuracy across its final three checkpoints (taken at 90B, 95B, and 100B tokens) to account for variability due to training randomness. Complete benchmark results across these checkpoints, along with key summary statistics and discussions, are provided in Appendix L.
Finally, all training and evaluation runs were executed on 208 NVIDIA H200 GPUs, with mixed-precision (bfloat16) enabled for training. The total training time for all models—Avey, Transformer++, Mamba, and RWKV-7—across the three presented model sizes and 100B training tokens is estimated to be approximately 80–90 hours, assuming optimal parallelization across the 208 GPUs and using the implementations referenced above. To avoid potential sources of randomness and ensure consistency across results, we disabled Torch Compile during all design choice and sensitivity experiments. For the ablation studies and final training runs, however, Torch Compile was enabled whenever possible to accelerate training. Additionally, we fixed the random seed to 11 (arbitrarily chosen) for all training runs to further reduce variability due to stochastic effects.
Appendix B Activation or No Activation in the Enricher
Table 3: Avey’s performance without and with an activation function in the enricher. The study involves only the neural processor within Avey and trains it on 10B tokens.
Configuration Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
No Activation 37.65 22.53 35.19 28.95 27.40 60.45 36.13 48.70 37.05
GELU 30.10 23.04 37.88 31.37 26.00 63.55 36.95 51.85 38.66
ReLU 31.02 22.87 37.21 31.35 28.00 62.79 38.13 48.86 38.46
ReLU2 30.18 24.15 38.76 32.08 28.00 63.76 38.28 50.12 39.31
SiLU 30.81 22.18 38.43 31.30 28.20 62.30 37.05 53.20 38.95
In this study, we illustrate the performance of Avey with and without an activation function in the enricher. To this end, we trained a model with 145 million parameters using 10 billion tokens from the Fineweb dataset [41]. The model employs an expansion factor of 4×4$\times$4 × in the enricher (i.e., each embedding dimension is expanded fourfold by the enricher), a tail size of 50% (i.e., half of each expanded embedding is forwarded to the contextualizer), RMSNorm [108] as a normalization technique, and no activation function in the contextualizer. Additionally, the context width (i.e., the maximum number of tokens that can be input to and processed by the contexutalizer simultaneously) is set to 1024 and a constant learning rate of $1e{-3}$ is maintained throughout training.
The study excludes the ranker and focuses solely on the neural processor. Besides, it evaluates four activation functions, namely, GELU [33], ReLU [62], ReLU2 [10], and SiLU [76]. All other experimental settings follow the methodology detailed in Appendix A. Table 3 summarizes the results. As shown, ReLU2 yielded an improvement in performance versus a baseline with no activation, hence, was adopted as the default activation function for the enricher throughout our experiments presented in Sections 3.3 and 3.4. It is important to note, however, that the lowest perplexity was provided by GELU and not ReLU2 (although the difference in perplexity was very minimal). While perplexity quantifies how effectively the model predicts the next token in the training dataset, it remains a proxy for overall modeling capability and does not always precisely predict downstream task performance.
Appendix C Activation or No Activation in the Contextualizer
Table 4: Avey’s performance without and with an activation function in the contextualizer. The study involves only the neural processor within Avey, trains it on 10B tokens, and uses ReLU2 within the enricher, capitalizing on the results shown in Table 3.
Configuration Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
No Activation 30.18 24.15 38.76 32.08 28.00 63.76 38.28 50.12 39.31
GELU 30.64 22.01 37.54 31.23 27.20 64.74 37.21 50.67 38.66
ReLU 30.30 23.29 37.84 31.82 26.60 64.15 38.13 50.59 38.92
ReLU2 31.05 22.35 38.80 30.78 27.20 63.49 37.31 52.01 38.85
SiLU 30.92 23.63 36.74 31.51 27.40 64.20 36.18 50.04 38.53
We now evaluate Avey with and without an activation function in the contextualizer. We use the same experimental settings outlined in Appendix B and add to that ReLU2 as an activation function in the enricher, capitalizing on the findings therein. We also experiment with four activation functions, namely, GELU, ReLU, ReLU2, and SiLU. Results are summarized in Table 4. As illustrated, the best performance was achieved without any activation function in the contextualizer, thus was employed as the default configuration in all our experiments reported in Sections 3.3 and 3.4.
Appendix D What is the Best Expansion Factor?
Table 5: The effect of the expansion factor on Avey’s performance. The study involves only the neural processor within Avey, trains it on 10B tokens, adopts ReLU2 within the enricher, and does not use an activation function within the contextualizer, building upon the results portrayed in Tables 3 and 4.
Expansion Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
30.85 23.29 36.83 30.92 26.20 63.98 37.72 51.78 38.67
30.18 24.15 38.76 32.08 28.00 63.76 38.28 50.12 39.31
30.00 23.21 37.79 31.48 26.20 63.82 36.59 50.75 38.55
In this study, we vary the expansion factor in the enricher, defined as the degree to which each input embedding is expanded. Specifically, we evaluate several expansion factors, ranging from $2x$ to $8x$ as shown in Table 5, while keeping the total model parameter count constant (e.g., with a $2x$ expansion factor we use 34 layers, while with a $4x$ one we utilize 20 layers). The experimental setup follows the methodology outlined in Appendix B, but incorporates ReLU2 as an activation function in the enricher and omits any activation function in the contextualizer, aligning with the findings reported in Appendices B and C. As depicted in the table, an expansion factor of $4x$ yielded the best performance, hence, was set as the default configuration in the enricher throughout all our experiments presented in Sections 3.3 and 3.4.
Appendix E What is the Best Tail Size?
Table 6: The effect of the tail size on Avey’s performance. The study involves only the neural processor within Avey, trains it on 10B tokens, adopts ReLU2 within the enricher, does not use an activation function within the contextualizer, and utilizes an expansion factor of $4x$, as recommended by the findings demonstrated in Tables 34, and 5.
Tail Size Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
10% 34.23 21.59 36.78 30.34 27.0 63.60 37.15 50.28 38.11
30% 30.91 23.55 37.33 31.41 29.4 64.25 37.15 51.22 39.19
50% 30.18 24.15 38.76 32.08 28.0 63.76 38.28 50.12 39.31
70% 29.79 23.04 38.26 31.68 28.4 63.55 37.36 49.96 38.89
90% 30.20 23.04 38.38 32.13 27.6 64.04 37.67 50.91 39.11
We now examine the impact of forwarding a tail portion of each enriched embedding to the contextualizer. More precisely, we vary the size of this tail portion, referred to as the tail size, from 10% to 90% of each enriched embedding, as illustrated in Table 6. The study follows the experimental setup described in Appendix B and employs ReLU2 as an activation function in the enricher, no activation function in the contextualizer, and an expansion factor of $4x$, based on the findings presented in Appendices B, C, and D, respectively. As depicted in Table 6, the best performance was accomplished using a tail size of 50%, thus was adopted as the default configuration for Avey in all our experiments reported in Sections 3.3 and 3.4.
Appendix F Deeper Models and Narrower Embeddings, or Shallower Models and Wider Embeddings
Table 7: Avey’s performance across different model configurations, including wider embedding dimensions (e.g., 1536 under 0.5B-parameter model) with shallower layers (e.g., 24 layers under 0.5B-parameter model), or narrower embedding dimensions (e.g., 768 under 0.5B-parameter model) with deeper layers (e.g., 90 layers under 0.5B-paramter model). The models with 140M, 0.5B, and 1.5B parameters are referred to as small Avey, medium Avey, and large Avey in the text.
# Params Embed. # Layers Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
140 M 512 40 31.14 22.53 37.58 31.05 28.6 63.87 37.46 52.64 39.10
768 20 30.18 24.15 38.76 32.08 28.0 63.76 38.28 50.12 39.31
1024 11 31.46 23.46 38.30 30.78 27.0 63.93 37.72 48.86 38.58
0.5 B 768 90 23.02 23.55 42.05 38.61 30.2 66.10 39.20 51.78 41.64
1024 54 23.27 24.40 41.92 38.46 29.2 67.19 38.74 51.46 41.62
1536 24 23.51 23.98 42.85 37.79 29.4 66.97 38.89 51.78 41.67
1.5 B 1536 80 19.97 25.26 45.50 44.56 30.2 68.61 40.02 52.33 43.78
2048 48 19.84 25.77 46.55 44.99 31.6 69.42 40.17 52.17 44.38
2560 30 20.23 26.62 45.16 43.91 29.2 69.10 39.82 52.09 43.70
Table 8: Avey’s performance under different sequence lengths, $N$, split sizes, $S$, and top-$k$ values. All models (a total of 69) were trained on 10B tokens using 140 million parameters. The sweet spot in terms of downstream task performance was at $N=512$, $S=64$, and $k=7$; hence, it was adopted as Avey’s default configuration.
$N$ $S$ $k$ Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
256 16 1 43.59 21.16 37.67 30.86 27.20 64.20 37.72 51.85 38.66
3 31.83 22.18 39.81 32.23 26.80 63.76 38.02 51.07 39.12
5 27.14 24.15 39.02 32.57 26.00 64.20 38.02 53.20 39.59
7 25.65 22.27 38.59 31.89 28.60 64.80 37.97 50.43 39.22
9 24.28 23.72 38.55 32.44 28.20 64.15 37.36 51.54 39.42
11 22.89 22.53 39.27 32.45 27.40 65.07 38.02 51.30 39.43
13 23.85 22.35 38.97 31.04 26.80 64.91 37.51 49.72 38.76
15 22.04 22.61 38.80 32.66 28.60 65.61 38.54 52.25 39.87
32 1 36.58 23.98 39.18 32.57 26.80 64.25 37.36 50.83 39.28
3 31.51 22.78 39.56 33.15 25.00 64.91 38.74 52.25 39.48
5 30.06 23.38 38.76 33.51 28.60 65.34 37.87 52.41 39.98
7 30.58 24.23 38.97 33.14 26.80 65.23 37.82 51.14 39.62
64 1 32.76 23.89 39.18 33.26 27.60 66.05 38.69 50.99 39.95
3 30.90 23.04 40.11 33.70 27.40 64.91 38.08 50.83 39.72
128 1 31.30 22.27 39.65 33.07 28.20 65.45 39.20 51.78 39.95
512 16 1 42.51 22.27 37.42 31.38 27.80 63.87 37.31 51.07 38.73
3 29.28 22.87 38.76 31.99 28.00 64.96 36.49 52.17 39.32
5 24.64 22.01 38.05 32.69 26.80 63.98 36.44 52.01 38.85
7 23.49 23.46 38.13 31.72 26.40 64.58 37.82 49.41 38.79
9 20.79 23.38 39.02 31.74 27.00 64.04 38.23 51.14 39.22
11 19.52 23.63 37.88 32.13 27.20 64.91 38.02 53.67 39.63
13 19.45 22.44 37.54 31.34 26.80 63.33 36.80 50.83 38.44
15 17.95 21.84 36.66 31.39 27.80 64.09 37.82 51.07 39.87
32 1 35.47 22.87 39.94 32.49 28.40 64.47 38.18 51.70 39.72
3 29.49 22.95 39.18 33.22 25.60 65.13 39.00 50.51 39.37
5 27.99 22.78 37.71 33.46 28.20 65.02 38.33 52.72 39.75
7 28.07 22.01 40.07 33.45 29.20 64.80 37.67 50.75 39.71
9 27.17 23.89 39.02 33.46 28.40 64.91 38.59 50.20 39.78
11 26.77 22.87 39.65 32.55 27.20 64.09 38.33 51.14 39.40
13 25.72 23.55 38.97 33.52 29.00 65.67 37.56 51.85 40.02
15 26.29 22.70 39.23 32.53 29.40 64.80 38.08 50.04 39.54
64 1 31.67 23.46 39.35 33.15 27.80 65.02 38.13 51.70 39.80
3 29.51 23.38 37.92 33.12 28.40 65.72 39.10 50.83 39.78
5 29.31 24.23 39.77 33.17 27.60 64.58 38.33 52.09 39.97
7 28.02 24.49 39.98 33.77 29.80 65.13 38.08 51.30 40.36
128 1 29.25 23.72 39.90 33.76 28.20 64.09 37.10 50.99 39.68
3 29.77 22.70 39.10 33.38 28.80 65.23 38.74 51.62 39.94
256 1 29.26 22.70 39.02 33.49 27.00 64.25 37.51 52.41 39.48
1024 16 1 41.64 21.42 37.16 31.12 29.80 64.47 37.56 50.75 38.90
3 28.08 22.61 38.26 31.88 27.20 64.64 38.08 50.99 39.09
5 23.69 22.18 38.38 31.94 28.80 64.09 37.87 51.22 39.21
7 21.48 23.81 38.05 31.41 27.00 63.38 36.80 50.59 38.72
9 19.83 22.53 37.50 31.90 26.80 64.47 37.92 49.64 38.68
11 18.34 21.93 37.16 31.45 28.60 65.13 37.77 50.67 38.96
13 16.80 23.55 37.50 30.55 26.80 63.11 36.95 52.33 38.68
15 15.33 23.29 37.54 31.04 27.60 63.06 37.77 50.91 38.74
32 1 35.07 22.70 39.98 32.87 27.60 65.23 37.77 51.14 39.61
3 28.54 23.55 38.55 32.91 27.60 64.74 37.51 50.28 39.31
5 26.25 22.95 39.06 33.39 28.60 64.64 38.28 50.04 39.57
7 26.29 24.06 38.76 32.70 27.60 64.80 37.67 53.12 39.82
9 24.79 23.63 38.89 33.34 27.80 64.53 37.72 52.09 39.71
11 24.33 21.93 38.76 32.56 26.40 64.36 37.36 51.38 38.96
13 23.44 22.78 37.46 32.73 29.00 65.23 37.31 50.67 39.31
15 23.14 23.72 39.56 32.39 28.40 63.60 37.31 51.38 39.48
64 1 30.84 23.89 38.51 33.31 27.20 65.18 38.28 49.88 39.46
3 27.82 22.61 39.60 33.39 28.60 64.74 38.84 50.20 39.71
5 27.89 23.04 40.49 32.97 30.00 65.02 38.49 49.09 39.87
7 27.48 24.06 39.27 33.54 28.80 65.72 37.72 50.75 39.98
9 27.37 22.35 39.98 33.48 28.00 65.29 38.74 52.64 40.07
11 27.38 23.29 39.35 33.05 28.60 65.72 37.97 50.67 39.81
13 27.44 24.23 39.77 33.05 28.20 66.05 37.31 51.07 39.95
15 26.85 24.32 40.36 34.01 28.60 65.45 37.72 51.30 40.25
128 1 28.62 23.12 40.57 33.56 27.80 65.61 38.54 51.30 40.07
3 27.08 23.89 39.81 33.65 29.00 64.69 37.72 52.09 40.12
5 28.07 23.98 40.03 33.22 29.60 65.89 39.10 50.67 40.35
7 27.27 24.32 38.85 33.92 27.40 65.13 38.13 49.88 39.66
256 1 28.30 22.61 38.38 33.22 27.80 64.85 38.79 51.30 39.56
3 28.53 23.04 39.90 32.68 27.40 64.36 37.62 49.88 39.27
512 1 28.24 24.40 39.39 32.98 28.20 65.13 37.72 50.59 39.77
The objective of this study is to determine whether a narrower embedding dimension with a greater model depth yields better or worse performance than a wider embedding dimension with fewer layers. The study utilizes the experimental setup described in Appendix B and leverages the findings presented in Appendices BCD, and E. Consequently, it utilizes ReLU2 as an activation function in the enricher, no activation function in the contextualizer, an expansion factor of $4x$, and a tail size of 50%.
To begin with, we evaluated a small Avey model (referred to as small Avey) with 140 million parameters, using three different embedding dimensions, 512, 768, and 1024. These configurations resulted in 40, 20, and 11 layers, respectively, to maintain a constant parameter count. Table 7 summarizes the results. As illustrated, the configuration with an embedding dimension of 768 and a layer count of 20 outperformed the other two configurations.
Afterwards, we assessed a larger Avey model with 500 million parameters (referred to as medium Avey), using three different embedding dimensions, 768, 1024, and 1536. These configurations resulted in 90, 54, and 24 layers, respectively, while keeping the total parameter count constant. As shown in Table 7, the setup with an embedding dimension of 768 and 90 layers yielded the best performance among the three tested ones.
Finally, we examined an even larger Avey model with 1.5 billion parameters (referred to as large Avey), using three different embedding dimensions, 1536, 2048, and 2560. To maintain a constant parameter count across configurations, these dimensions corresponded to 80, 48, and 30 layers, respectively. As portrayed in Table 7, the configuration with an embedding dimension of 2048 and 48 layers delivered the best performance among the three considered configurations.
The above results suggest a trend, whereby wider embedding dimensions (e.g., 1024 in small Avey; 1563 in medium Avey; and 2560 in large Avey) paired with shallower architectures (e.g., 11 layers in small Avey; 24 layers in medium Avey; and 30 layers in large Avey) tend to underperform deeper models (e.g., 40 and 20 layers in small Avey; 90 and 54 layers in medium Avey; and 80 and 48 layers in large Avey) with narrower embeddings (e.g., 512 and 768 in small Avey; 768 and 1024 in medium Avey; and 1536 and 2048 in large Avey). As such, in all our experiments discussed in Sections 3.3 and 3.4, we employed deeper models with narrower embedding dimensions, namely, the best performing small Avey, medium Avey, and large Avey in Table 7.
Interestingly, Table 7 also highlights that certain benchmarks benefit more from increased model capacity than others. For instance, a commonsense reasoning benchmark like HellaSwag demonstrates performance improvements of 20.36% and 28.5% under 0.5B-parameter and 1.5B-parameter models, respectively, compared to a 140M-parameter baseline. In contrast, a question-answering benchmark such as SIQA exhibits only a modest gain of 2.4% under both 0.5B-parameter and 1.5B-parameter models relative to the 140M-parameter baseline, suggesting less sensitivity to model size.
Appendix G What are the Best Sequence Length, Split Size, and Top-$k$ Values?
We now analyze how Avey’s perplexity and overall performance are affected by variations in three key parameters, the ranker’s top-$k$ selected splits, the split size $S$, and the sequence length $N$. We consider three values for $N$, $256$, $512$, and $1024$. For each $N$, the split size $S$ is grown geometrically, starting from $16$ and doubling at each step, up to the maximum permissible value $N/2$. Subsequently, for any $N$ and $S$, the number of top-ranked splits $k$ can range from $1$ (i.e., contextualizing the current split with one additional relevant split) up to $N/S-1$. To tame the quantity of experiments, we increase $k$ arithmetically from $1$ to a maximum of $15$, whenever possible, using an increment of $2$. Finally, we use the experimental configurations recommended in Appendices BCDE, and F.
As shown in Table 8, Avey’s perplexity is highest when both $S$ and $k$ are very small (e.g., $S=16$ and $k=1$). While a small $S$ can help filter out irrelevant embeddings and denoise contextualization, pairing it with a very small $k$ can deprive the contextualizer of sufficient context to build expressive representations [^9] Recall from Section 2.2.2 that the context width $C$ is defined as $C=S(k+1)$.. To expand the context (i.e., increase its width $S(k+1)$) and enrich the resulting embeddings, either $S$ or $k$ can be increased. For example, as we increased $k$ under $N=256$ and $S=16$, perplexity decreased and benchmark performance improved. However, a larger context does not always yield better outcomes, especially when involving a high proportion of irrelevant embeddings. This behavior was observed when $k$ was increased beyond $3$ under $N=1024$ and $S=128$, resulting in higher perplexity and diminished downstream accuracy.
In contrast to $S$ and $k$, the sequence length $N$ determines the size of the candidate pool from which the ranker selects the top-$k$ splits for any current split during training. A larger $N$ allows the ranker to reach farther back in the sequence history, potentially retrieving more relevant splits and lowering perplexity. For example, increasing $N$ from $512$ to $1024$, while holding $S=16$ and $k=15$ constant, reduced perplexity by $5.4\%$. Nevertheless, lower perplexity does not always translate into improved downstream performance. For instance, with $N=512$ and $S=16$, $k=15$ yielded the lowest perplexity, yet $k=5$ achieved higher benchmark accuracy. As noted in Appendix B, the loss remains only a proxy for overall modeling capability and does not always exactly predict downstream task performance. As demonstrated in Table 8, the best empirical performance was obtained with $N=512$, $S=64$, and $k=7$, which was, accordingly, adopted as Avey’s default configuration.
Appendix H RMSNorm or LayerNorm
Table 9: Avey’s performance with RMSNorm and LayerNorm. A model with 153 million parameters was trained on 10B tokens using both the neural processor and ranker with the best configuration from Table 8.
Normalization Method Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
RMSNorm 28.02 24.49 39.98 33.77 29.8 65.13 38.08 51.30 40.36
LayerNorm 30.93 23.55 39.65 33.24 28.8 65.07 38.28 49.57 39.74
In all previous runs across the appendices, we used RMSNorm [108] as a normalization technique. We now test Avey with another standard normalization method, namely, LayerNorm [1]. To begin with, we note that Avey normalizes input embeddings before each layer (as illustrated in Fig. 3) and output embeddings once after the final layer and prior to token prediction.
We evaluate each normalization technique using Avey’s complete architecture, including its neural processor and ranker. The ranker is configured according to the best-performing setting identified in Appendix G (i.e., sequence length $N=512$, split size $S=64$, and number of top-ranked splits $k=7$). The neural processor employs the optimal configurations reported in Appendices BCDE, and F. As before, we trained the resulting model with 153 million parameters on 10 billion tokens from the Fineweb dataset. Table 9 summarizes the results. As shown, RMSNorm slightly outperforms LayerNorm on average, and is therefore adopted in Avey and used consistently across all the experiments presented in Sections 3.3 and 3.4.
Appendix I Learning Rate: To Decay or Not to Decay?
Table 10: Avey’s performance with two types of schedules, constant learning rate (LR) and cosine decay, starting from different peak learning rates. All models involved the neural processor and ranker, and were trained with 153 million parameters on 10B tokens, using both the neural processor and ranker with the best configuration from Table 8.
Schedule LR Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
Constant 8e-04 28.21 23.12 39.81 33.52 28.4 64.96 38.13 52.41 40.05
1e-03 27.35 24.06 40.32 33.88 29.6 65.13 38.54 51.46 40.43
3e-03 30.38 23.81 38.34 32.90 28.2 63.38 37.77 52.33 39.53
Cosine Decay 6e-04 26.24 22.87 40.95 33.76 29.2 65.02 37.72 49.01 39.79
8e-04 25.64 24.06 39.31 34.43 29.6 65.83 37.87 49.80 40.13
1e-03 25.00 23.21 41.12 34.76 27.0 65.67 38.38 50.75 40.13
In all previous runs across the appendices, we used a constant learning rate of $1e{-3}$. In this study, we evaluate the performance of Avey under varying maximum learning rates and learning rate schedules. Specifically, we compare two schedules, cosine decay and constant learning rate (which can be viewed as cosine decay with an infinite cycle length, effectively eliminating any decay). The study adopts the experimental configurations suggested in Appendices BCDE, and F. In addition, it employs Avey’s complete architecture, including its neural processor and ranker, with the best configuration from Table 8.
As shown in Table 10, cosine decay consistently achieves lower losses than the constant schedule across the tested learning rates. This observation aligns with findings from the Chinchilla paper [35], which indicates that when the cosine cycle length significantly exceeds the total number of training steps (by at least 25%), model performance tends to deteriorate. Notably, the longest cycle length arises when the schedule is constant. In contrast, setting it to approximately match the training duration yields the best final loss [35].
To this end, we adopt a cosine decay schedule with a peak learning rate of $1e{-3}$ for Avey, especially that it provides the lowest loss across all the runs. We note, however, that Table 10 also shows that the lowest loss does not correspond to the best downstream task performance. For instance, the loss of 3.308 under the constant learning rate schedule resulted in a slightly better benchmark performance than the lower loss of 3.218 under the cosine decay schedule, both using the same peak learning rate. While constant learning rates can be effective for short or exploratory runs (this study uses only 10 billion tokens), it is generally the case that, as the number of training tokens increases, the learning rate must decrease to allow the optimizer to settle into a lower-loss region [103]. Hence, schedules with decay are typically favored for longer or large-scale training runs [35, 4].
Appendix J Scaling Laws
Table 11: Model configurations used in the scaling law experiments. Each model is trained at three different sizes and numbers of training tokens increased proportionally, following the Chinchilla scaling laws.
Model # Layers (# Heads) Embedding Dim. Learning Rate # Tokens
Avey-153M 26 768 1.00e-03 2B
Avey-496M 104 768 1.00e-03 7B
Avey-1.5B 48 2048 1.00e-03 20B
Transformer++-152M 12 (12) 768 3.00e-03 2B
Transformer++-488M 26 (16) 1024 1.50e-03 7B
Transformer++-1.5B 32 (16) 1664 1.25e-03 20B
Mamba-153M 28 768 1.00e-03 2B
Mamba-496M 42 1280 1.00e-03 7B
Mamba-1.5B 52 2048 1.00e-03 20B
RWKV-7-152M 12 768 3.00e-03 2B
RWKV-7-488M 30 1024 1.50e-03 7B
RWKV-7-1.5B 24 2048 1.25e-03 20B
Figure 6: Scaling law results, comparing how perplexity decreases as compute increases, assuming three model sizes of 150M, 500M, and 1.5B parameters and a proportional increase in the number of training tokens with model size, following the Chinchilla scaling laws.
In this section, we present a scaling law study comparing how well Avey, Transformer++, Mamba, and RWKV-7 scale with increasing compute. For Avey, we use the full architecture, including both the ranker and neural processor. All models are trained at three different sizes, as defined in Appendix A and summarized in Table 11. To ensure compute-optimal scaling, we proportionally increase the number of training tokens with model size, following the Chinchilla scaling laws [35]. Specifically, and in line with the methodology from [25], we use 2B, 7B, and 20B tokens to train models with approximately 150M, 500M, and 1.5B parameters, respectively. Lastly, we employ the same batch size across all the models to control for variability in the number of gradient update steps, especially because of training with a limited number of tokens.
Fig. 6 presents the scaling results. The x-axis represents the total training compute budget, calculated as the product of the number of training tokens and model parameters, which serves as a proxy for the total FLOPs required to train each model. As shown, Avey exhibits the steepest decline in perplexity as compute increases. Although it begins with a relatively high perplexity [^10] Avey can achieve substantially lower perplexity under alternative configurations. For example, the small model (150M parameters) with sequence length $N=1024$, split size $S=16$, and top-$k=15$ achieves much lower perplexity as shown in Table 8. The configuration used in this experiment– and in Sections 3.3 and 3.4, was selected based on its strong downstream benchmark performance, rather than optimal perplexity. As discussed in Appendix B, perplexity serves as a useful proxy for modeling capability, but does not always align perfectly with downstream task accuracy., it improves more rapidly than the other models, indicating strong scaling behavior and greater benefit from additional compute. Following Avey, Transformer++ demonstrates the next-best scaling trend, outpacing Mamba and RWKV-7. While Mamba achieves relatively low perplexity at smaller compute budgets, it does not scale as effectively as Avey or Transformer++. Finally, RWKV-7 performs well at low compute but shows the flattest scaling curve, suggesting it gains the least from additional training compute.
Appendix K Ablation Study
Table 12: Ablation results comparing Avey variants, with individual components removed or replaced.
Model Variant ppl ARC-C ARC-E Hella OBQA PIQA SIQA Wino Average
Avey full (all features) 30.00 25.17 39.90 33.59 28.8 65.56 37.62 51.62 40.32
Avey without dynamic parameterization 34.31 25.00 40.66 32.99 28.8 65.34 36.64 50.51 39.99
Avey without bypassing 32.55 22.61 38.38 32.31 28.0 64.20 38.28 52.09 39.41
Avey without embedding expansion 39.94 22.44 37.92 28.75 25.4 62.40 38.64 52.01 38.22
Avey without weighting selected splits 31.17 22.78 38.55 33.25 28.0 65.89 37.82 52.09 39.77
Avey without the ranker 29.48 23.72 38.59 32.52 28.0 63.66 37.67 53.20 39.62
Avey with self-attention in place of neural proc. 31.39 22.61 39.27 31.99 28.0 64.58 38.33 51.38 39.45
In this study, we perform a series of ablation experiments on the core components of Avey, leveraging the best configurations identified in Appendices BCDEFH, and I. All experiments are conducted using Avey’s complete architecture, comprising both the ranker and neural processor, with the small model variant (153 million parameters) as the baseline. For this study, we trained this model on 10 billion tokens from the FineWeb dataset, using the training methodology and hyperparameters detailed in Appendix A.
We begin by examining the effect of dynamic parameterization on both perplexity and downstream benchmark performance. As described in Section 2.2.2, dynamic parameterization allows each neuron in the contextualizer network to compute a cosine similarity between its input embedding and the embeddings of all other neurons, weight those embeddings accordingly, and aggregate them via a learned weighted sum. This mechanism induces selectivity, as defined in [25], into the neural processor, thereby making its parameterization dynamic (or input-dependent). Table 12 summarizes the results. Disabling this component results in a 14.3% increase in perplexity and a 0.8% drop in average performance, highlighting its importance.
Second, we evaluate the impact of the partial-embedding bypassing technique introduced in Section 2.2.1. This method involves forwarding a portion of each expanded embedding directly to the fuser, allowing raw, distinctive features to be preserved and potentially serve in promoting more diverse representations. As shown in Table 12, removing this mechanism results in an 8.5% increase in perplexity and a 2.2% drop in average performance, underscoring its significance.
Third, we set the expansion factor in the enricher to 1, effectively disabling the expansion of input embeddings. As illustrated in Table 12, this modification results in a 33.1% increase in perplexity and a 5.2% drop in average performance, corroborating the critical role of embedding expansion in the model’s effectiveness.
Fourth, we remove the weighting of each selected split by its corresponding normalized MaxSim score, thereby preventing the contextualizer from scaling each split’s contribution during contextualization. As depicted in Table 12, this adjustment leads to a 3.8% increase in perplexity and a 1.37% drop in average performance, indicating the importance of this technique.
Fifth, we evaluate Avey without the ranker to assess its impact on downstream task performance, beyond its primary role of enabling effective extrapolation past the trained context window. As shown in Table 12, the ranker does indeed enhance the neural processor’s performance, primarily by improving the quality of contextualization through more meaningful cross-token interactions. We note, however, the slight increase in loss (by 0.5%), which again highlights (as in Appendix B) the discrepancy between the objectives of pertaining and downstream tasks.
Finally, we replace Avey’s neural processor with self-attention to assess the relative contribution of each component to Avey’s overall performance, given that both are designed to pursue cross-token interactions. As illustrated in Table 12, this substitution leads to a 4.6% increase in perplexity and a 2.1% decline in average performance, underscoring the significance of the neural processor and suggesting that self-attention is less effective within Avey’s architectural framework.
Appendix L Additional Short-Range Benchmark Results
Table 13: Performance of all models across short-range benchmarks at 90B, 95B, and 100B training tokens.
Model (# of Tokens) ARC-C ARC-E HellaSwag PIQA OBQA SIQA Winogrande Avg.
Avey-153M (100BT) 23.98 42.30 39.57 29.8 68.61 39.05 51.85 42.02
Avey-153M (95BT) 24.23 42.09 39.21 31.2 68.23 39.15 50.28 42.06
Avey-153M (90BT) 24.91 42.59 39.31 33.2 68.28 39.20 51.70 42.74
Transformer++-152M (100BT) 23.29 43.43 39.47 29.4 67.03 39.10 50.51 41.90
Transformer++-152M (95BT) 23.55 43.14 39.51 30.2 67.14 38.69 50.99 41.89
Transformer++-152M (90BT) 24.06 42.93 38.97 29.8 66.87 38.89 49.17 41.24
Mamba-144M (100BT) 24.32 43.73 40.82 29.8 68.28 39.00 52.41 42.62
Mamba-144M (95BT) 23.63 43.69 40.51 32.2 68.06 39.82 53.35 43.61
Mamba-144M (90BT) 24.57 43.18 40.33 29.2 68.61 39.41 52.41 42.53
RWKV-7-168M (100BT) 23.89 43.14 41.50 29.8 68.72 39.41 50.99 42.35
RWKV-7-168M (95BT) 24.23 42.89 41.77 29.2 68.99 39.10 51.14 42.48
RWKV-7-168M (90BT) 24.40 43.01 41.38 30.0 68.44 39.00 51.14 42.48
Avey-496M (100BT) 27.13 48.99 52.17 32.0 72.47 40.53 54.54 46.55
Avey-496M (95BT) 27.90 49.20 51.74 33.0 73.07 40.63 53.51 46.72
Avey-496M (90BT) 27.47 48.65 51.56 32.4 71.93 39.30 55.09 46.63
Transformer++-488M (100BT) 25.68 48.02 52.92 31.6 72.69 39.56 55.96 46.06
Transformer++-488M (95BT) 27.39 47.90 52.69 31.6 72.36 40.07 54.22 46.12
Transformer++-488M (90BT) 27.13 48.36 52.37 32.0 71.33 40.17 55.56 46.16
Mamba-500M (100BT) 29.27 51.26 54.45 34.0 73.88 40.38 54.70 48.28
Mamba-500M (95BT) 28.67 51.39 54.25 34.8 72.69 40.89 55.33 48.29
Mamba-500M (90BT) 27.99 50.42 53.76 34.6 72.52 41.25 56.43 48.14
RWKV-7-501M (100BT) 26.96 49.83 54.49 36.0 73.23 39.30 55.17 47.71
RWKV-7-501M (95BT) 27.39 49.24 54.66 35.6 73.78 39.15 55.80 47.95
RWKV-7-501M (90BT) 27.05 49.03 54.46 37.2 73.72 39.76 56.20 48.20
Avey-1.52B (100BT) 30.89 56.36 61.49 34.4 75.84 42.07 56.59 51.09
Avey-1.52B (95BT) 32.34 56.94 61.63 37.6 75.57 41.76 58.09 52.42
Avey-1.52B (90BT) 30.55 56.36 61.15 38.4 75.41 42.17 56.51 51.51
Transformer++-1.5B (100BT) 30.29 56.19 64.28 38.8 76.12 42.27 61.33 52.75
Transformer++-1.5B (95BT) 30.97 57.07 63.87 37.0 76.17 42.07 61.72 52.70
Transformer++-1.5B (90BT) 28.75 55.60 63.45 38.2 75.73 42.37 61.09 52.19
Mamba-1.4B (100BT) 32.42 57.87 64.78 38.4 76.61 42.48 62.27 53.55
Mamba-1.4B (95BT) 32.85 57.91 64.37 35.4 76.33 42.02 60.93 52.69
Mamba-1.4B (90BT) 32.00 58.63 64.38 36.8 76.22 41.50 61.33 52.69
RWKV-7-1.5B (100BT) 32.42 59.55 64.59 37.4 76.82 41.86 59.67 53.19
RWKV-7-1.5B (95BT) 33.11 58.88 64.49 37.0 76.88 41.71 60.38 53.21
RWKV-7-1.5B (90BT) 33.28 58.71 64.21 37.0 76.82 41.56 60.14 53.10
Table 14: Summary statistics for each model with different sizes computed over the last three checkpoints (i.e., at 90B, 95B, and 100B training tokens).
Model Mean Standard Deviation Standard Error 95% Confidence Interval
Avey-153M 42.32 0.3683 0.2126 (41.41, 43.24)
Transformer++-152M 41.72 0.1821 0.1052 (41.27, 42.17)
Mamba-144M 42.73 0.2700 0.1559 (42.06, 43.40)
RWKV-7-168M 42.48 0.0094 0.0054 (42.46, 42.51)
Avey-496M 46.82 0.1895 0.1094 (46.35, 47.29)
Transformer++-488M 46.65 0.0507 0.0293 (46.52, 46.77)
Mamba-500M 48.23 0.0835 0.0482 (48.03, 48.44)
RWKV-7-501M 48.00 0.1807 0.1043 (47.55, 48.45)
Avey-1.52B 51.53 0.4497 0.2596 (50.41, 52.65)
Transformer++-1.5B 52.54 0.3218 0.1858 (51.74, 53.34)
Mamba-1.4B 53.12 0.3783 0.2184 (52.18, 54.06)
RWKV-7-1.5B 53.17 0.0553 0.0320 (53.03, 53.30)
To mitigate the effects of fluctuations in pre-training loss and downstream benchmark scores, we reported in Section 3.3 average results across the final three checkpoints—taken at 5 billion token intervals (i.e., at 90B, 95B, and 100B tokens)—for all models evaluated, namely, Avey, Transformer++, Mamba, and RWKV-7. In this section, we provide the detailed performance scores for each model at each checkpoint in Table 13. In addition, we summarize the mean, standard deviation, standard error, and 95% confidence interval for each model, computed across the three checkpoints, in Table 14. The illustrated statistical results reveal meaningful variance between models and across runs of the same model. For instance, while Mamba achieves the highest mean score of 42.73 among all the models in the small parameter regime ($\sim$150M parameters), it also exhibits a relatively wide confidence interval (42.06, 43.40) and a moderate standard deviation, highlighting nontrivial variability in performance across checkpoints.
Regarding variability across models, Table 14 shows overlapping confidence intervals, indicating that model rankings—particularly which model achieves the highest mean performance—could shift under minor experimental changes (e.g., random initialization, stochastic optimization, etc.). For example, in the small model regime, while Avey does not surpass Mamba in mean performance, their confidence intervals substantially overlap in the range (42.06, 43.24), suggesting that the two models are statistically comparable and that Avey could outperform Mamba in some runs. Similarly, a narrow but meaningful overlap exists between Avey and RWKV-7 in the range (42.46, 42.51), implying that Avey may occasionally match or slightly exceed RWKV-7 in certain cases. Lastly, although Mamba has the highest mean in this setting, its confidence interval also overlaps with RWKV-7, indicating that the difference in performance between the two models is not statistically significant and that RWKV-7 could match or slightly outperform Mamba in some runs.
In the medium model regime ($\sim$500M parameters), Avey outperforms Transformer++ in mean performance, but they are statistically comparable. In contrast, the performance gap between Avey and both Mamba and RWKV-7 is statistically significant at the 95% confidence level, indicating that both models clearly outperform Avey in this setting. In the large model regime ($\sim$1.5B parameters), while Avey does not outpace Transformer++ in average performance, their confidence intervals overlap substantially, suggesting that Avey could potentially surpass Transformer++ in some runs. There is also a limited overlap between Avey and Mamba, indicating that while Mamba generally performs better, Avey might outperform it in certain cases. In contrast, the difference between Avey and RWKV-7 is statistically significant at the 95% level, confirming that RWKV-7 consistently outperforms Avey in this setting. Finally, although RWKV-7 has a slightly higher mean than Mamba (53.17 vs. 53.12), the meaningful overlap in their confidence intervals implies that the difference between them is not statistically significant, and either model could outperform the other depending on minor experimental factors.
Appendix M Additional Long-Range Benchmark Results
Figure 7: Performance comparison between Transformer++, Mamba, RWKV-7, and Avey on S-NIAH-1 and S-NIAH-2. The x-axis denotes the lengths of haystacks (i.e., documents with distractor texts, varying from 2k to 64k tokens). All models use $\sim$150M parameters.
In Section 3.4, we presented results for Avey, Transformer++, Mamba, and RWKV-7 under the medium parameter regime ($\sim$500M parameters) on the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER [38], which is designed to evaluate models’ abilities to handle long-range dependencies. The S-NIAH benchmark, along with two of its common variants—S-NIAH-1 and S-NIAH-2—was described in detail in Section 3.4. In this section, we extend our analysis by reporting results under two additional model regimes, small ($\sim$150M parameters) in Fig. 7 and large ($\sim$1.5B parameters) in Fig. 8. Akin to the experiment in Section 3.4, Transformer++, Mamba, and RWKV-7 were trained with a context window of 2048 tokens, while Avey was trained with a shorter window of only 512 tokens.
Figure 8: Performance comparison between Transformer++, Mamba, RWKV-7, and Avey on S-NIAH-1 and S-NIAH-2. The x-axis denotes the lengths of haystacks (i.e., documents with distractor texts, varying from 2k to 64k tokens). All models use $\sim$1.5B parameters.
In both the small and large model regimes, under S-NIAH-1 and S-NIAH-2, Transformer++, Mamba, and RWKV-7 perform well when the haystack length is 2k, fitting within their trained context windows. Yet, Mamba consistently underperforms Transformer++ and RWKV-7, likely due to solely relying on recurrence, which somehow treats the entire input uniformly, making the model more susceptible to distractions from irrelevant tokens. In contrast, RWKV-7, which combines recurrence with attention, performs better than Mamba but remains inferior to Transformer++, potentially because the attention mechanism allows it to prioritize tokens relevant to the needle, while the recurrent component may still contribute to signal dilution. Transformer++, relying exclusively on full attention, achieves the best performance within the context window by effectively focusing on relevant tokens without interference from recurrence-based mechanisms. Nonetheless, once the haystack length exceeds the models’ context windows, all the three models exhibit a substantial drop in performance. Mamba and RWKV-7, however, show minimal generalization beyond their training limits compared to Transformer++, as previously discussed in Section 3.4.
Compared to Transformer++, Mamba, and RWKV-7, Avey generalizes far beyond its trained context window on both S-NIAH-1 and S-NIAH-2 across all parameter regimes, underscoring its strong extrapolative capabilities (as also shown in Section 3.4). Notably, this holds despite Avey being trained with a context window of only 512 tokens. For example, in the small parameter regime, Avey achieves an accuracy of 97.8% on S-NIAH-1 with a 64k-token haystack, while Transformer++, Mamba, and RWKV-7 drop to 0%, 0%, and 0.6%, respectively. Similarly, on S-NIAH-2 at the same haystack length, Avey attains 68.8% accuracy, whereas Transformer++, Mamba, and RWKV-7 fall to 0%, 2.8%, and 2%, respectively. Comparable trends are observed in the large parameter regime as well, as illustrated in Fig. 8.
An interesting observation arises in the small parameter regime, where Avey outperforms all other models on S-NIAH-1 with a haystack length of 2k, knowing that this length exceeds its trained context window width and enables it to demonstrate its strong extrapolative capability. However, this pattern does not persist in the medium (see Fig. 5 in Section 3.4) and large (see Fig. 8) parameter regimes, where Transformer++ and RWKV-7 outperform Avey on the same benchmark at 2k length, despite this length still surpassing Avey’s trained context window. This suggests that these models, with their increased capacity, are able to compensate for the challenge posed by S-NIAH-1, and entails that Avey might benefit from a longer training context window.
In this paper, we kept Avey’s context window fixed at 512 tokens across all parameter regimes. All tuning experiments related to sequence length, split size, and top $k$ splits (see Section G) were conducted exclusively using the small model size. It is plausible that with larger capacity, Avey could more effectively leverage longer sequences by retrieving and contextualizing a larger set of relevant tokens while filtering out less informative ones, thereby enhancing contextual representations and further boosting performance. Investigating the relationship between sequence length and model size in Avey is an interesting direction for future work.
Appendix N Complexity Analysis
As indicated in Sections 2.1, 2.2.1, 2.2.2, and 2.2.3, the training time complexities of the ranker, enricher, contextualizer, and fuser are $\mathcal{O}(N^{2}d)$, $\mathcal{O}(Nmd)$, $\mathcal{O}(NkCm_{t})$, and $\mathcal{O}(Nmd)$, respectively, where $N$ is the sequence length, $S$ is the split size, $d$ is the original embedding dimension, $m$ is the projected embedding dimension (with $m>d$), $m_{t}$ is the tail part of $m$ forwarded to the contextualizer, $C$ is the context width (with $C\leq N$), and $k$ is the number of splits contextualized with each current split. This yielded an overall training time complexity of $\mathcal{O}(N^{2}d)$, assuming that scalar multiply-add operations (e.g., those used in computing cosine similarity for MaxSim) and comparisons (e.g., those used to determine maximum scores) are constant-time.
During inference, the time complexities of the enricher, contextualizer, and fuser remain unchanged. However, the ranker’s analysis slightly changes, as at each time step $t$ (i.e., upon predicting a new token), the current split is compared against all previous splits. More precisely, at each time step $t$, the current split—denoted as split $i$ and incrementally filled as tokens are generated—is compared against all $i-1$ preceding splits, each consisting of $S$ tokens. Consequently, the cost of comparing $t$ tokens in split $i$ (with $t\leq S$) against $S$ tokens in a previous split is $\mathcal{O}(t\cdot S\cdot d)$.
Now, if we let $M=\lceil N/S\rceil\approx N/S$ for large $N$ be the number of splits, the total inference cost can be defined as:
$\sum_{i=1}^{M}\sum_{t=1}^{S}(i-1)\cdot\mathcal{O}(t\cdot S\cdot d)$
Simplifying the inner summation yields:
$\sum_{t=1}^{S}(i-1)\cdot\mathcal{O}(t\cdot S\cdot d)=(i-1)\cdot\mathcal{O}(S\cdot d)\cdot\sum_{t=1}^{S}t=(i-1)\cdot\mathcal{O}(S\cdot d)\cdot\frac{S(S+1)}{2}=(i-1)\cdot\mathcal{O}(S^{2}\cdot d)$
Substituting this back into the outer summation gives:
$\sum_{i=1}^{M}(i-1)\cdot\mathcal{O}(S^{2}\cdot d)=\mathcal{O}(S^{2}\cdot d)\cdot\sum_{i=1}^{M}(i-1)=\mathcal{O}(S^{2}\cdot d)\cdot\frac{M(M-1)}{2}$
Substituting $M$ with $N/S$ for large $N$ results in:
$\mathcal{O}(S^{2}\cdot d)\cdot\frac{(N/S)^{2}}{2}=\mathcal{O}(S^{2}\cdot d)\cdot\mathcal{O}(N^{2}/S^{2})=\mathcal{O}(N^{2}\cdot d)$
Therefore, the cost per token becomes:
$\frac{\mathcal{O}(N^{2}d)}{N}=\mathcal{O}(Nd)$
Appendix O Related Work
Recurrent Neural Networks (RNNs) [19, 79] are designed to process sequential data by capturing temporal dependencies, making them well-suited for tasks where input order is essential. However, their cyclical nature limits their potential for parallel computation and exposes them to vanishing and exploding gradient problems. As a result, they typically struggle to effectively learn long-range dependencies. While architectures like Long Short-Term Memory (LSTM) [34] and Gated Recurrent Units (GRU) [8] mitigate these gradient-related issues, they remain slow to optimize and challenging to scale due to retaining RNN’s core recurrent and sequential structure.
In contrast, the Transformer [98] employs a self-attention mechanism to process each sequence of tokens simultaneously. More precisely, it promotes two key design principles: (1) a recurrent-free architecture, which enables parallel computation of token embeddings, while still capturing their order through positional encodings, and (2) a multi-head self-attention approach, which facilitates cross-token interactions to further enrich the expressiveness of embeddings. These innovations make the Transformer highly effective, as well as parallelizable and efficient to train. However, they also cause its computational and memory requirements to scale quadratically with sequence length, making it expensive and less efficient for very long sequences.
To address the Transformer’s quadratic computation and memory costs, a wide array of approaches have been proposed, including linear attention [50, 48, 9, 71, 106], sparse or local attention [104, 7, 67], context compression [74, 99, 90, 78], and modified attention computations [93, 100, 92], to mention just a few. Notably, the Attention-Free Transformer (AFT) [106] offers a linear drop-in replacement for the quadratic self-attention mechanism. In particular, it weights key and value vectors using learned positional biases and integrates them with query vectors via element-wise multiplication. As such, it eliminates the need to compute and store the costly attention matrix while still preserving global query-value interactions—without requiring architectural modifications or additional tuning. Furthermore, AFT introduces variants such as AFT-local and AFT-conv, which leverage local attention patterns to reduce parameter count and further improve computational and memory efficiency.
RWKV-4 [68] (the first 3 versions were experimental [52]) capitalizes on AFT and suggests combining the strengths of both Transformers and RNNs. To elaborate, unlike Transformers and akin to RNNs, it does not process each input token solely based on its own embedding, but rather as a weighted sum of its embedding and that of the preceding one. To the contrary of traditional RNNs and similar to Transformers, it adopts self-attention, but an extended version of it, namely, that of AFT. This hybrid approach allows RWKV-4 to maintain some of the modeling capabilities of RNNs, while benefiting from the parallelization and scalability features of Transformers.
More precisely, RWKV-4 extends AFT in two distinct ways: (1) it introduces an additional parameter to handle each current token independently, and (2) it implements a per-time-step decay mechanism that selectively removes older content from the recurrent hidden state in a data-dependent manner. This decay mechanism addresses a central limitation of linear attention, which pertains to the lack of a systematic way to discard outdated information [83, 102].
Architecturally, RWKV-4 consists of homogeneous stacked residual blocks, each encompassing two units, a time-mixing and a channel-mixing ones. The time-mixing unit applies linear attention across tokens, while the channel-mixing unit integrates each element of the current token’s embedding with its corresponding element from the preceding token embedding, leveraging the output of the time-mixing unit.
RWKV-5 [69] enhances RWKV-4’s architecture and learning decay mechanism by replacing traditional vector-valued states with more expressive multi-head, matrix-valued ones. Moreover, it reconfigures receptive states, incorporates supplementary gating mechanisms, and dynamically learns the linear interpolation between the current and preceding token embeddings instead of relying on pre-defined hyperparameters.
RWKV-6 [69] promotes a new application of low-rank adaptation functions [39, 52]. Specifically, it makes the linear interpolation between the current and preceding tokens data-dependent to improve the selectivity of the model in retaining and discarding information. Additionally, it replaces the static decay mechanism with a dynamic one, allowing each element in the decay vector to fluctuate independently over time in response to the input.
The decay strategies of RWKV-4, RWKV-5, and RWKV-6 still cannot remove values stored at specific keys. DeltaNet [83] overcomes this drawback by partially replacing the values stored at current keys with equivalent new values, enabling models to erase outdated memories and include up-to-date ones on a per-key basis. However, it only allows a fixed scalar fraction of a value to be replaced from a state via an in-context learning rate parameter, thus demonstrating rigidity in adapting to varying data contexts.
RWKV-7 [70] builds upon the principles of DeltaNet and introduces a vector-valued in-context learning rate instead of a scalar-valued one. This allows selective replacement of state data on a channel-wise basis. Furthermore, RWKV-7 employs a vector-valued decay mechanism and uses additional low-rank projections to optimize the trade-off between the number of parameters, computational efficiency, and downstream performance. Lastly, it incorporates Value Residual Learning [110], which improves the propagation of initial local information via utilizing a residual connection between the value vectors of the current layer and those of the first layer prior to the attention operation, resulting in enhanced language modeling performance.
Most recently, RWKV-X [37] proposed combining the strengths of RWKV and sparse attention, drawing inspiration from Mixture of Block Attention (MoBA) [57]. In particular, RWKV-X restricts each query to attend only to a small, relevant subset of the input, thus reducing computational cost and facilitating the modeling of longer-range dependencies. More precisely, rather than allowing each token to attend to every other token in the sequence (as in traditional self-attention), it constrains each token’s attention to a limited subset (hence, making it sparse), while maintaining the coupling between the input sequence and context window.
Similar to RWKV, RetNet [91] adopts linear attention and promotes a hybrid approach that blends Transformer- and RNN-like representations, yet with a decay-based memory unit. Specifically, it divides the input sequence into chunks, wherein the Transformer-like parallel representation is applied. Additionally, it enables propagating information sequentially across chunks using the RNN-like representation. Lastly, it uses multiple attention heads, each governed by a distinct decay rate, and replaces LayerNorm [1] with GroupNorm [101].
Although linear attention has been proposed as a promising alternative to quadratic softmax attention [48, 9, 47, 71], existing implementations of it are in practice slower than optimized versions of softmax attention [14, 12, 102]. From an accuracy standpoint, linear attention generally underperforms conventional softmax attention, sometimes by a significant margin in language modeling [47, 102].
To this end, and in light of the exponentially growing complexity associated with overcoming the limitations of Transformer-based architectures, there has been a renewed interest in RNN-based alternatives in recent years. Notably, Structured State Space Sequence (S4) models [27, 28], inspired by the classical state space models (SSMs) [44], have emerged as a promising paradigm for sequence modeling. These models describe the temporal evolution of a system using differential equations, offering a continuous-time formulation of dynamics, and can be viewed as generalized versions of RNNs.
An SSM as a concept has a broad meaning, which simply refers to the notion of any recurrent process with a latent state [25]. From this perspective, the RNN-like linear attention model proposed and formulated by [48] can be interpreted as a degenerate linear SSM. Interestingly, this justifies the usage of a decay factor in RetNet and RWKV, especially that a decay term (or a forget gate) has been shown to be crucial in RNNs [34, 97, 8].
Numerous variants of SSMs [27, 29, 30, 51, 58, 65, 87] have demonstrated strong performance across a range of domains, including audio and vision [24, 63, 81]. Nonetheless, these variants have struggled with language modeling, often lagging behind Transformers by several points in perplexity [27].
From an efficiency standpoint, however, SSMs have shown encouraging results in language modeling. For instance, S4 [27, 28], a prominent SSM, converts the continuous-time state update equation of SSMs into a discrete form, hence, enabling parallel sequence modeling. Moreover, it utilizes the HiPPO (High-Order Polynomial Projection Operator) initialization [26], which alleviates the vanishing gradient problem and facilitates processing longer sequences.
Another example of SSMs is H3 [20], which improves language modeling by allowing both, the recall of earlier tokens and token-wise comparisons within a sequence. H3 extends S4 by suggesting a state-passing algorithm that enhances computational efficiency on modern hardware. This advancement reduces the hardware-related barriers that have traditionally limited the scalability of SSM-based architectures.
Hyena [72] capitalizes on H3 by replacing its S4 layer with an MLP-parameterized global convolution [77]. S5 [87] proposes using parallel scan [59] to parallelize S4. Liquid S4 [31] augments S4 with an input-dependent state transition matrix, computed convolutionally in the frequency domain (which is computationally efficient) and mapped back to the time domain using an inverse Fourier transformation. SGConv [51], LongConv [21], MultiresConv [86], and Toeplitz Neural Network [73] all focus on the convolutional representation of S4 as well, aiming to enhance its efficiency [25].
Most recently, Mamba [25] introduced a new class of SSMs known as selective SSMs, specifically designed to improve the performance of language modeling tasks. Mamba addresses a key limitation in SSMs, namely, their inability to selectively process inputs in an input-dependent manner (i.e., focus on or ignore specific parts of the input sequence). Consequently, it makes the SSM parameters input-dependent, but introduces a technical challenge since traditional SSMs are inherently designed to be time- and input-invariant to ensure computational efficiency. To overcome this challenge, Mamba proposes a hardware-efficient parallel scan (or prefix sum) algorithm [6], which enables recurrent-style computation without explicitly materializing the expanded state. This design precludes costly I/O operations across GPU memory hierarchies and accelerates both, training and inference.
Tri Dao and Albert Gu [13] argue that various approaches to operating SSMs can be reframed as matrix multiplication algorithms involving a specific class of structured matrices known as semiseparable matrices. They further leverage the language of tensor contractions to prove the recurrent formulation of linear attention as proposed by [48], before generalizing it to a new family of structured masked attention (SMA).
Subsequently, Tri Dao and Albert Gu demonstrated that SSMs and Transformers are fundamentally connected, governed by the mathematical framework of semiseparable matrices and SMA. Additionally, they developed a rich state space duality (SSD) framework of theoretical connections between SSMs and various forms of attention. This framework facilitated the design of Mamba-2, an extended version of Mamba, which aims to improve its efficiency (not performance). Mamba-2 utilizes a scalar data-dependent gating mechanism (like the ones proposed by [71, 91, 3]), which enables transforming its recurrent structure into a matrix-multiply form, thus allowing for efficient execution on tensor cores and better support for larger hidden state sizes.
The strategy of the SSD framework mirrors that of linear attention [48], which established a connection between autoregressive attention mechanisms and linear RNNs via showing an equivalence between "dual forms" of quadratic kernelized attention and a specific type of linear recurrence. Conceptually, the SSD framework seeks to transfer algorithmic and systems-level optimizations originally developed for Transformers to the realm of SSMs. Its overarching goal is to enable the development of architectures that outperform Transformers, while scaling more efficiently with sequence length.
Finally, several works, including Tolstikhin et al. [95], Melas-Kyriazi [60], Touvron et al. [96], and Ding et al. [18], among others, have questioned the necessity of self-attention, particularly in the context of Vision Transformers. In contrast, Liu et al. [53] introduced gMLP, an MLP-based alternative to BERT-style Transformers [16] that (partially) eliminates self-attention but ultimately underperforms average performance on downstream NLP tasks. gMLP encompasses channel (hidden) and spatial (cross-token) projections with multiplicative gating and static parameterization. Its gating mechanism is reminiscent of Gated Linear Units (GLUs) [15, 84, 100], as well as earlier architectures such as Highway Networks [88] and LSTM-RNNs [34]. A key distinction, however, is that gMLP applies gating on the spatially projected dimension and not the hidden one. The gated embedding-wise neural network in Avey’s contextualizer draws inspiration from gMLP.
Unlike all previously mentioned models, Avey entirely abandons self-attention and recurrence, introducing a new architecture composed of a ranker and a dynamically parameterized neural processor. The ranker identifies the most relevant tokens for contextualization, while the neural processor contextualizes them data-dependently. This design decouples sequence length from context width, enabling efficient processing of arbitrarily long sequences without diminishing the influence of distant yet important tokens.
At the core of Avey’s architecture is a weighted-selective-split interaction mechanism, which filters out irrelevant tokens beyond the context window and enables direct interactions only with relevant ones, thus preserving their influence irrespective of sequence length. In addition, Avey employs a partial-embedding bypassing technique that retains a portion of each token’s raw, distinctive features before fusing them with its contextualized ones through a neural network. This technique boosts the performance of Avey (as shown in Appendix K) and might help mitigate issues such as entropy collapse [107] and over-smoothing [109, 85], especially at large-scale, when the depth of the model is increased significantly.
Appendix P Limitations
The scope of our work is limited to textual data and does not involve other modalities such as images, audio, or genomics. Additionally, our evaluation of Avey is restricted to standard autoregressive language modeling, benchmarking it against popular open-source architectures using both pretraining metrics (perplexity) and zero-shot evaluations on established NLP benchmarks. As a result, we do not investigate Avey’s ability to construct bidirectional contextualized word representations, as done in BERT [16]. We leave this for future work. Finally, the paper focuses solely on effectiveness rather than efficiency. While we provide a complexity analysis showing that Avey exhibits quadratic training time like Transformers, our current implementation is slower. As such, further engineering efforts are required to optimize it.

### v2

@@
 Don’t Pay Attention
-Abstract
-The Transformer has become the de facto standard for large language models and a wide range of downstream tasks across various domains.
-Despite its numerous advantages like inherent training parallelism, the Transformer still faces key challenges due to its inability to effectively process sequences beyond a fixed context window and the quadratic complexity of its attention mechanism.
-These challenges have renewed interest in RNN-like architectures, which offer linear scaling with sequence length and improved handling of long-range dependencies, albeit with limited parallelism due to their inherently recurrent nature.
-In this paper, we propose Avey, a new neural foundational architecture that breaks away from both attention and recurrence.
-Avey comprises a ranker and an autoregressive neural processor, which collaboratively identify and contextualize only the most relevant tokens for any given token, regardless of their positions in the sequence.
-Specifically, Avey decouples sequence length from context width, thus enabling effective processing of arbitrarily long sequences.
-Experimental results show that Avey compares favorably to the Transformer across a variety of standard short-range NLP benchmarks, while notably excelling at capturing long-range dependencies.
+Don’t Pay Attention
+[…]
+svg-inkscape/.
+Don’t Pay Attention Mohammad Hammoud,  Avey AI,  mhh@avey.ai Devang Acharya,  Avey AI,  dacharya@avey.ai The Transformer has become the de facto standard for modern language models owing to its parallelizable training and effective autoregressive decoding.
+However, its fixed context window and the quadratic time and memory costs of its self-attention mechanism remain central bottlenecks.
+These constraints have revived interest in recurrent architectures that scale linearly with sequence length, but at the cost of reduced parallelism.
+In this paper, we introduce Avey, a new foundational architecture that breaks away from both attention and recurrence.
+Avey pairs a ranker with an autoregressive neural processor to select and contextualize only the most relevant tokens for any given token.
+Specifically, it decouples sequence length from context width, thus enabling effective and efficient processing of arbitrarily long sequences.
+Results show that Avey compares favorably to the Transformer across a variety of standard short-range NLP benchmarks, while significantly outperforming it on tasks requiring long-range dependency modeling.
+Date: November 14, 2025 Website: https://avey.ai/research Equal contribution.
 1 Introduction
-The Transformer [98] has emerged as one of the most influential AI innovations in recent years, profoundly impacting various aspects of modern life, including work, science, and art, to mention just a few.
-Notably, Large Language Models (LLMs) are almost universally based on the Transformer [25], which has demonstrated remarkable performance in natural language processing (NLP) [66, 55, 75] and various other fields [32, 56, 2].
+The Transformer (vaswani2017attention) has emerged as one of the most influential AI innovations in recent years, profoundly impacting various aspects of modern life, including work, science, and art, to mention just a few.
+Notably, Large Language Models (LLMs) are almost universally based on the Transformer (gu2023mamba), which has demonstrated remarkable performance in natural language processing (NLP) (ouyang2022training; liu2019roberta; raffel2020exploring) and various other fields (he2022masked; liu2021swin; baevski2020wav2vec).
 The Transformer’s state-of-the-art performance is largely driven by its recurrence-free self-attention mechanism, which enables parallel processing of entire token sequences.
 Nevertheless, the computational and memory costs of self-attention grow quadratically with sequence length, making it inefficient for handling arbitrarily long sequences.
-Extensive research has been conducted over the years to address this limitation [94], with a noticeable emphasis on linearizing attention [48, 9, 106, 99].
+Extensive research has been conducted over the years to address this limitation (tay2022efficient), with a noticeable emphasis on linearizing attention (katharopoulos2020transformers; choromanski2020rethinking; zhai2021attention; wang2020linformer).
 These linear approaches aim at approximating self-attention in a more computationally efficient manner, without considerably compromising performance.
-Nonetheless, linear attention mechanisms have generally underperformed the original self-attention mechanism, often by a significant margin in language modeling tasks [102, 47].
-While recent linear models such as RWKV [70] and RetNet [91] have shown promising results, substantial progress is still needed before they can reliably and consistently surpass the Transformer [52].
-In addition, these models have yet to demonstrate definitive empirical effectiveness at scale [25].
+Nonetheless, linear attention mechanisms have generally underperformed the original self-attention mechanism, often by a significant margin in language modeling tasks (yang2023gated; kasai2021finetuning).
+While recent linear models such as RWKV (peng2025rwkv) and RetNet (sun2023retentive) have shown promising results, substantial progress is still needed before they can reliably and consistently surpass the Transformer (li2024survey).
+In addition, these models have yet to demonstrate definitive empirical effectiveness at scale (gu2023mamba).
 This persistent performance gap between quadratic and linear approaches has spurred renewed interest in RNN-based architectures, which offer linear scalability with sequence length but limit parallelism due to their inherently cyclical nature.
-To exemplify, state space models (SSMs) [44, 28], which are viewed as extensions of RNNs, have recently emerged as a compelling class of architectures.
+To exemplify, state space models (SSMs) (kalman1960new; gu2021combining), which are viewed as extensions of RNNs, have recently emerged as a compelling class of architectures.
 Unlike traditional RNNs, SSMs can parameterize their state transition matrices in a structured manner (e.g., via using a diagonal plus low-rank decomposition) to improve computational efficiency and enhance gradient flow.
-A specialized subclass of these models, known as structured state space sequence (S4) models [27, 28], has garnered growing attention.
-Yet, despite their theoretical appeal, S4 models struggled with language modeling tasks, typically trailing Transformers by several points [20, 27].
-Most recently, Mamba [25] advanced S4 models by enhancing their selectivity and effectiveness while enabling high training concurrency.
+A specialized subclass of these models, known as structured state space sequence (S4) models (gu2021efficiently; gu2021combining), has garnered growing attention.
+Yet, despite their theoretical appeal, S4 models struggled with language modeling tasks, typically trailing Transformers by several points (fu2022hungry; gu2021efficiently).
+Most recently, Mamba (gu2023mamba) advanced S4 models by enhancing their selectivity and effectiveness while enabling high training concurrency.
 It demonstrated strong performance on tasks involving long-range dependencies and compared favorably to Transformers in language modeling.
-However, training, scaling, and interpreting Mamba—and SSMs more broadly [87, 72, 31]—remain challenging, while continue to be promising [13].
+However, training, scaling, and interpreting Mamba—and SSMs more broadly (smith2022simplified; poli2023hyena; hasani2022liquid)—remain challenging, while continue to be promising (dao2024transformers).
 We posit that the primary limitation of the Transformer lies in its inability to effectively model dependencies beyond its fixed context window.
@@
 Under this paradigm shift, the quadratic training complexity becomes less of a concern when small context windows are used, especially if the models maintain high parallelizability.
-This paper introduces Avey [^1] Avey is not an acronym, but a name that the authors like., a new architecture for language modeling that entirely departs from Transformer-based and RNN-like designs.
+This paper introduces Avey [^1] Avey is not an acronym, but a name that the authors like., a new architecture for language modeling that departs from Transformer-based and RNN-like designs.
 Avey is a flexible, sequence-length-invariant model that decouples sequence length from context width, thus enabling effective processing of long-range sequences.
@@
 Figure 1: Needle-in-a-Haystack test performance comparison between Transformer++, Mamba, RWKV-7, and Avey, all using 1.5B parameters.
-The x-axis denotes the lengths of haystacks (i.e., documents with distractor texts, varying from 2k to 64k tokens) and the y-axis refers to the position of the needle (i.e., a short sentence) within any of the haystacks.
-A green cell means that the model was able to recall the needle and a red one entails that it could not.
-Transformer++, Mamba, and RWKV-7 were all trained with 2k-token contexts, while Avey was trained with only a 512-token context and was able to extrapolate to the maximum needed modeling capacity.
+The x-axis denotes the lengths of haystacks (i.e., documents with distractor texts, varying from 2k to 64k tokens) and the y-axis refers to the position of the needle (i.e., a short sentence) within each of the haystacks.
+A green cell indicates successful needle recall, while a red cell indicates failure.
+Transformer++, Mamba, and RWKV-7 were trained with 2k-token context windows, whereas Avey was trained with only a 512-token window yet was able to extrapolate to the longest sequences evaluated.
 Fig.
 1 illustrates Avey’s ability to generalize beyond its training context.
-A popular benchmark for evaluating this capability is Needle-in-a-Haystack (NiaH) [45].
+A popular benchmark for evaluating this capability is Needle-in-a-Haystack (NiaH) (kamradt2023needle).
 This benchmark measures a model’s capacity to recite a specific sentence (i.e., the needle) placed at an arbitrary position within a large body of distractor text (i.e., the haystack).
-Since its introduction, NiaH has become a widely used sandbox for probing the limits of long-context language models in capturing distant dependencies, and smaller models in generalizing beyond their trained context windows [22].
-As shown in the figure, Transformer++ (i.e., the Transformer with an enhanced architecture and training recipe– see Section 3.1), which was trained with a 2k-token context window, could not generalize beyond that limit.
-In contrast, Mamba and RWKV-7 [70], also trained with 2k-token context windows, managed to generalize to nearly 8k and 16k tokens, respectively.
+Since its introduction, NiaH has become a widely used sandbox for probing the limits of long-context language models in capturing distant dependencies, and smaller models in generalizing beyond their trained context windows (fu2024data).
+As shown in the figure, Transformer++ (i.e., the Transformer with an enhanced architecture and training recipe– see Section 3.1 for details), which was trained with a 2k-token context window, could not generalize beyond that limit.
+In contrast, Mamba and RWKV-7 (peng2025rwkv), also trained with 2k-token windows, managed to generalize to nearly 8k and 16k tokens, respectively.
 Most notably, Avey, despite being trained on a context window of only 512 tokens, successfully generalized to the maximum tested sequence length of 64k tokens, demonstrating strong extrapolative capability far beyond its original training regime.
@@
   • We show that Avey performs comparably to the Transformer—outperforming it at two model sizes and underperforming it at one—across a range of popular zero-shot NLP benchmarks, thereby establishing an initial foundational architecture with potential for more scalable and effective language modeling.
-  • In contrast to the Transformer, we demonstrate that Avey can scale far beyond its context window using the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER [38].
+  • In contrast to the Transformer, we demonstrate that Avey can scale far beyond its context window using the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER (hsieh2024ruler).
   • We show that Mamba (representing SSMs) and RWKV-7 (representing linear attention models) exhibit some ability to generalize beyond their training context windows, but their performance decline significantly as the sequence length increases far beyond them.
@@
   • We conduct extensive ablation studies to assess the impact of each design choice in Avey.
-  • We provide a comprehensive survey of related work in Appendix O.
   • We open-source the code and pretrained checkpoints of Avey to facilitate reproducibility and foster future research [^2] https://github.avey.ai/avey-dpa..
 2 Avey
-As indicated earlier, Avey comprises two components, a ranker and an autoregressive neural processor. Below, we provide a detailed explanation of each component.
+As indicated earlier, Avey comprises two components, a ranker and a neural processor. We next describe each component in detail (see Appendix G for more design intuitions behind them).
 2.1 Ranker
@@
 The current split is defined as the one that either contains the token to be predicted (e.g., split 4 may contain only embedding 7, and Avey will aim to predict token 8) or contributes to predicting the first token in its subsequent split (e.g., split 4 serves in predicting token 9, which will belong to split 5 once predicted).
-To determine relevance, the ranker computes a similarity score between the current split, say $S_{c}$, and each preceding split, say $S_{p}$, using the MaxSim operator [49], originally proposed and utilized in Information Retrieval.
+To determine relevance, the ranker computes a similarity score between the current split, say $S_{c}$, and each preceding split, say $S_{p}$, using the MaxSim operator (khattab2020colbert), originally proposed and utilized in Information Retrieval.
 Specifically, pairwise similarities are calculated (e.g., using a cosine function) between each embedding in $S_{c}$ and all embeddings in $S_{p}$.
@@
 This empowers the neural processor to judiciously leverage global information (i.e., splits beyond the context width) by focusing selectively on only relevant features, emphasizing informative ones and deemphasizing less useful ones, thus enhancing performance.
-We analyze the impact of weighting the top k splits in Appendix K.
-Note that the ranker is invoked only once per full forward and backward pass.
+We analyze the impact of weighting the top k splits in Section 3.3.
+Note that the ranker is invoked only once per full forward and backward passes [^3] It is worth emphasizing that the ranker is an internal module that operates solely over in-sequence splits already present in the input.
+In other words, it does not access external corpora or indexes, and therefore introduces neither retrieval latency nor corpus-freshness dependencies.
+Consequently, it is not a RAG component, which retrieves external evidence at inference (and/or training) time from a separate knowledge base.
+The two approaches are orthogonal in fact (the ranker allocates internal context, while RAG modifies the available evidence set).
+See Appendix E for a detailed discussion..
 To elucidate, Avey’s depth can be increased by stacking multiple layers within the neural processor (see Fig.
@@
 This complexity assumes that scalar multiply-add operations (e.g., those used in computing cosine similarity for MaxSim) and comparisons (e.g., those utilized to determine maximum scores) are constant-time.
-We next discuss the neural processor.
 2.2 Neural Processor
@@
 The enricher can expand each input embedding by an arbitrary factor.
-We study the effect of varying the expansion factor on Avey’s performance in Appendix D, and ablate the enricher’s contribution in Appendix K.
+We study the effect of varying the expansion factor on Avey’s performance in Section 3.2.3, and ablate the enricher’s contribution in Section 3.3.
 Equation 1 formalizes the enricher, where $\mathbf{X}\in\mathbb{R}^{C\times d}$ is a matrix of $C$ input embeddings ($C\leq N$, where $N$ is the sequence length), each of dimension $d$; $\sigma$ is an activation function; $\mathbf{U}\in\mathbb{R}^{d\times m}$ is a learnable weight matrix defining a linear projection from dimension $d$ to $m$, where $m>d$; and $\mathbf{b}\in\mathbb{R}^{C\times m}$ denotes biases.
@@
 In particular, it bypasses a portion of each expanded embedding directly to the fuser in a technique that we refer to as partial-embedding bypassing.
-More precisely, the output of the enricher, $\mathbf{Z}\in\mathbb{R}^{C\times m}$, is split into two parts: (1) the head $\mathbf{Z}_{h}\in\mathbb{R}^{C\times m_{h}}$, which is bypassed directly to the fuser, and (2) the tail $\mathbf{Z}_{t}\in\mathbb{R}^{C\times m_{t}}$, which is forwarded to the contextualizer, where $m=m_{h}+m_{t}$.
+More formally, the output of the enricher, $\mathbf{Z}\in\mathbb{R}^{C\times m}$, is split into two parts: (1) the head $\mathbf{Z}_{h}\in\mathbb{R}^{C\times m_{h}}$, which is bypassed directly to the fuser, and (2) the tail $\mathbf{Z}_{t}\in\mathbb{R}^{C\times m_{t}}$, which is forwarded to the contextualizer, where $m=m_{h}+m_{t}$.
 Consequently, varying the tail size alters the head size, which can influence Avey’s performance.
-We investigate the impact of different tail sizes on Avey’s performance in Appendix E.
+We investigate the impact of different tail sizes on Avey’s performance in Section 3.2.4.
 The partial-embedding bypassing technique allows preserving raw distinctive features of each embedding, thus inducing representations with inherent diversity.
-This diversity may serve in alleviating issues like entropy collapse [107], where the contextualizer increasingly focuses on a few tokens, and over-smoothing [109, 85, 110], where embeddings become increasingly similar, as Avey’s depth is increased.
-We analyze the significance of partial-embedding bypassing on Avey’s effectiveness in Appendix K.
-Lastly, Equation 1 implies that each neuron performs a weighted sum of $d$ input features (i.e., the elements of an embedding), incurring $d-1$ multiply-add operations.
-Since $d$ is projected to a higher dimension $m$ [^3] In our case, we experiment with $m$ being a multiple of $d$, entailing that $m\geq 2d$ (see Appendix D)., the total computational cost is $m(d-1)$ per token.
-For a sequence of $N$ tokens, the cost becomes $Nm(d-1)$, or asymptotically $\mathcal{O}(Nmd)$.
+This diversity may serve in alleviating issues like entropy collapse (zhai2023stabilizing), where the contextualizer increasingly focuses on a few tokens, and over-smoothing (zhou2021deepvit; shi2022revisiting; zhou2024value), where embeddings become increasingly similar, as Avey’s depth is increased.
+We analyze the significance of partial-embedding bypassing on Avey’s effectiveness in Section 3.3.
+Lastly, Equation 1 implies that each neuron performs a weighted sum of $d$ input features (i.e., the elements of an embedding), incurring d1d-1 multiply-add operations.
+Since $d$ is projected to a higher dimension mm [^4] In our case, we experiment with $m$ being a multiple of $d$, entailing that $m\geq 2d$ (see Section 3.2.3)., the total computational cost is m(d1)m(d-1) per token.
+For a sequence of $N$ tokens, the cost becomes Nm(d1)Nm(d-1), or asymptotically $\mathcal{O}(Nmd)$.
 2.2.2 The Contextualizer
@@
 Both $m_{tl}$ and $m_{tr}$ are learnable by the model, hence, allowing $m_{tl}$ to dynamically capture the significance of each $m_{tr}$’s feature, and emphasize or deemphasize its influence accordingly.
-This gating mechanism was inspired from gMLP [53] and resembles that of Gated Linear Units [15, 84, 100].
+This gating mechanism was inspired from gMLP (liu2021pay) and resembles that of Gated Linear Units (dauphin2017language; shazeer2020glu; wu2019pay).
 More formally, $\mathbf{Z}_{t}\in\mathbb{R}^{C\times m_{t}}$ is partitioned into two equal parts, $\mathbf{Z}_{tl}\in\mathbb{R}^{C\times\left(m_{t}/2\right)}$, which is bypassed to a multiplicative element-wise operation as part of a gating mechanism, and $\mathbf{Z}_{tr}\in\mathbb{R}^{C\times\left(m_{t}/2\right)}$, which is contextualized via a neural network, where each neuron takes as input an embedding of dimension $m_{t}/2$.
@@
 Equation 2 suggests that each neuron in the contextualizer’s network performs a weighted sum of the cosine similarities between embeddings (denoted by $\mathcal{N}(\mathbf{Z}_{tr})\mathcal{N}(\mathbf{Z}_{tr}^{\top})$) and the embeddings themselves (denoted by $\mathbf{Z}_{tr}$).
-This introduces a level of selectivity into the neural processor, as described and advocated by [25].
+This introduces a level of selectivity into the neural processor, as advocated by (gu2023mamba).
 Specifically, it makes the parametrization of the neural processor dynamic, enabling it to disregard or focus on information during inference based on the input.
-We examine the influence of dynamic parametrization on Avey’s performance in Appendix K.
+We examine the influence of dynamic parametrization on Avey’s performance in Section 3.3555See also a discussion on neural contextualization versus attention in Appendix F..
 Finally, we note that the contextualizer inherently models the relationships between tokens, making the neural processor naturally aware of their positions in the sequence (i.e., positional encodings are not needed).
@@
 2.2.3 The Fuser
-The fuser is designed to learn an optimal function, referred to as fusion [^4] The name is inspired from the CNN literature [40]., that integrates uncontextualized features (i.e., those of dimension $m_{h}$, bypassed by the partial-embedding bypassing technique presented in Section 2.2.1) with contextualized features (i.e., those of dimension $m_{t}/2$, output by the contextualizer).
+The fuser is designed to learn an optimal function, referred to as fusion [^6] The name is inspired from the CNN literature (hu2018squeeze)., that integrates uncontextualized features (i.e., those of dimension $m_{h}$, bypassed by the partial-embedding bypassing technique) with contextualized features (i.e., those of dimension $m_{t}/2$, output by the contextualizer).
 Subsequently, it produces, for each input token, a contracted representation that matches the token’s original embedding dimension $d$ (see Fig.
@@
 Akin to the enricher, it is a one-layer, position-wise neural network, which operates on each embedding of dimension $m_{h}+m_{t}/2$ independently.
-[…]
+Equation 3 provides a mathematical definition of the fuser, where $\mathbf{Z}_{h}\in\mathbb{R}^{C\times m_{h}}$ (as described in Section 2.2.1) and $\mathbf{c}(\mathbf{Z}_{t})\in\mathbb{R}^{C\times(m_{t}/2)}$ (as suggested by Equation 2) are concatenated, and $\mathbf{O}\in\mathbb{R}^{(m_{h}+m_{t}/2)\times d}$ is a learnable weight matrix representing a linear projection from dimension $m_{h}+m_{t}/2$ back to dimension $d$, where $d<m_{h}+m_{t}/2$ [^7] This inequality will always hold if $m_{h}+m_{t}\geq 2d$, as is the case in our experiments (see Section 3)..
 $f(\mathbf{Z})=[\mathbf{Z}_{h}\,\|\,\mathbf{c}(\mathbf{Z}_{t})]\mathbf{O}$ (3)
@@
 During inference, the complexity reduces to $\mathcal{O}(N)$, or linear per token.
-We elaborate on Avey’s inference time complexity in Appendix N.
-Figure 4: The Time to First Token (TTFT) for Avey, Transformer++, Mamba, and RWKV-7 using $\sim$1.5B parameters, across varying sequence lengths.
-See Section 3.1 for details on the experimental methodology.
-To characterize Avey’s inference-time efficiency relative to other models, we benchmarked Time to First Token (TTFT)—a key latency metric for real-time applications [36, 54, 17]—on a single NVIDIA H200 GPU.
+We provide a detailed analysis of Avey’s time complexity in Appendix D.
+Figure 4: The Time to First Token (TTFT) for Avey, Transformer++, Mamba, and RWKV-7 across varying sequence lengths.
+Lastly, to characterize Avey’s inference-time efficiency relative to other models, we benchmarked Time to First Token (TTFT)—a key latency metric for real-time applications (horton2024kv; liu2025speculative; dexter2025llm)—on a single NVIDIA H200 GPU.
 The evaluation included Avey, Transformer++, Mamba, and RWKV-7 across a range of sequence lengths.
-As shown in Figure 4, Transformer++ demonstrates an approximately quadratic increase in TTFT as the sequence length $N$ grows, owing to its full self-attention mechanism, which must operate over the entire prompt before generating the first token.
+As shown in Fig.
+4, Transformer++ demonstrates an approximately quadratic increase in TTFT as the sequence length $N$ grows, owing to its full self-attention mechanism, which must operate over the entire prompt before generating the first token.
 In contrast, Mamba and RWKV-7 scale linearly with $N$, as they require a full forward pass to construct their RNN-style hidden states prior to emitting the first token.
 While Avey is also theoretically expected to scale linearly, its empirical TTFT is significantly lower than that of Transformer++, Mamba, and RWKV-7.
-This efficiency gain stems from Avey’s architectural design, which invokes its dominant contributor to inference complexity (see Appendix N), namely, the ranker, only once per forward pass.
+This stems from Avey’s architectural design, which invokes its dominant contributor to inference complexity (see Appendix D), namely, the ranker, only once per forward pass.
 As a result, the ranker imposes minimal computational overhead in practice, enabling Avey to achieve substantially lower TTFT and making it particularly well-suited for latency-sensitive, real-world applications (e.g., conversational AI and edge deployments).
 3 Experiments
-3.1 Experimental Setup
-We compare Avey against three leading open-source models, namely, Mamba [42], RWKV-7 [43], and Transformer++, extended to the strongest known architectural recipe of the standard Transformer [46] (see Appendix A for details).
-All models were trained using their best-known hyperparameters under a fixed budget of 100 billion tokens drawn from the FineWeb dataset [41].
-Complete training and model hyperparameters for all the baselines are provided in Appendix A.
-To assess each model’s accuracy, we employed a suite of widely used NLP benchmarks, including ARC-E and ARC-C [11], HellaSwag [105], PIQA [5], OBQA [61], SIQA [82], and Winogrande [80].
-Additionally, we evaluated the long-context retrieval capabilities of all the models using the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER [38].
-Full details of all the benchmarks and additional experimental setups are included in Appendix A.
-3.2 Design Choices and Ablations
+3.1 Methodology
+We now describe our experimental methodology in detail.
+To begin with, we adopt a cascaded search to identify the best configuration of Avey.
+Starting from a baseline neural processor (excluding the ranker), we sequentially evaluate individual architectural choices.
+After each empirical finding, we integrate the best-performing architectural element into the processor (one at a time) and resume the search process from the updated configuration.
+This cascaded procedure unfolds chronologically in Sections 3.2.1, 3.2.2, 3.2.3, 3.2.4, and 3.2.5.
+More precisely, we initiate the cascaded search with an expansion factor of $4\times$ in the enricher, a tail size of $50\%$ (i.e., half of each expanded embedding is forwarded to the contextualizer), RMSNorm (zhang2019rmsnorm) for normalization, no activations in either the enricher or contextualizer, a global batch size of 0.5M, a context width of 1024, and a constant learning rate of 1e31e{-3}.
+As each architectural element is explored and empirically selected, we incorporate it into the processor and resume the search from the updated configuration.
+For instance, after identifying $\mathrm{ReLU}^{2}$ as the most effective activation for the enricher, we included it in the architecture and proceeded with the remaining explorations.
+After completing the exploratory experiments described above, we integrated the ranker into the neural processor and conducted an extensive study—spanning over 138 training and inference runs—to identify the optimal sequence length ($N$), split size ($S$), and number of top-$k$ splits ($k$).
+The results of this sensitivity analysis are summarized in Section 3.2.6.
+Subsequently, we evaluated normalization strategies in Section 3.2.7 and tuned the peak learning rate and learning-rate schedule for the full architecture in Section 3.2.8.
+Table 1: Training and model hyperparameters for Avey, Transformer++, Mamba, and RWKV-7 at three scales.
+Avey Transformer++ Mamba RWKV-7 Scale Small Medium Large Small Medium Large Small Medium Large Small Medium Large # of Parameters 153M 496M 1.52B 152M 488M 1.5B 144M 500M 1.4B 168M 501M 1.5B Optimizer AdamW AdamW AdamW AdamW Betas (0.9, 0.95) (0.9, 0.95) (0.9, 0.95) (0.9, 0.95) Epsilon $1\mathrm{e}{-12}$ $1\mathrm{e}{-12}$ $1\mathrm{e}{-12}$ $1\mathrm{e}{-12}$ Peak learning rate $1\mathrm{e}{-3}$ $3\mathrm{e}{-3}$ $1.5\mathrm{e}{-3}$ $1.25\mathrm{e}{-3}$ $3\mathrm{e}{-3}$ $1.5\mathrm{e}{-3}$ $1.0\mathrm{e}{-3}$ $6\mathrm{e}{-4}$ $4\mathrm{e}{-4}$ $6\text{e}{-4}$ $4\text{e}{-4}$ $4\text{e}{-4}$ Schedule Cosine decay to 10% of peak; no warmup Linear warmup 10% of steps; cosine decay to 10% Linear warmup 10% of steps; cosine decay to 10% Cosine decay to 10% of peak; no warmup Batch size 0.5M 0.5M 1M 0.5M 0.5M 1M 0.5M 0.5M 1M 1M 1M 2M Gradient norm clip 1.0 1.0 1.0 1.0 Weight decay 0.1 (matrices only) 0.1 (matrices only) 0.1 (matrices only) 0.1 (matrices only) Context width 512 2048 2048 2048 Split size ($S$) 64 N/A N/A N/A Top-$k$ splits 7 N/A N/A N/A Vocabulary size 50,304 50,304 50,304 50,304 Expansion factor 4 N/A N/A N/A Tail size 0.5 N/A N/A N/A Embedding dimension 768 768 2048 768 1024 1664 768 1280 2048 768 1024 2048 Number of layers 26 104 48 12 26 32 28 42 52 12 30 24 Number of heads N/A 12 N/A 16
+All the aforementioned experiments were conducted with a 145M-parameter model trained on 10B tokens from the FineWeb dataset [^8] This dataset is released under the Open Data Commons Attribution License (ODC-By) v1.0.
+(fineweb2023) (specifically, the sample-100BT subset).
+These results informed the final choices of training and model hyperparameters for Avey across three parameter scales, 153M (small), 496M (medium), and 1.52B (large), as reported in Table 1.
+For baselines, we compare Avey against three leading open-source models, namely, Transformer++ (nanogpt2023), Mamba (mambai2023), and RWKV-7 (rwkvi2023).
+For Transformer++, we implement the strongest recipe known to us, incorporating rotary positional encodings (RoPE) (su2024roformer), SwiGLU MLPs (shazeer2020glu), and RMSNorm in place of LayerNorm (zhang2019rmsnorm).
+All models are trained with their best-known hyperparameters (see Table 1) at three parameter scales (i.e., small, medium, and large) under a fixed budget of 100B tokens drawn from the FineWeb dataset [^9] More precisely, all models—Avey, Transformer++, Mamba, and RWKV-7—are trained for one epoch over the sample-100BT subset of FineWeb..
+For consistency and comparability, we use the p50k_base tokenizer (tiktoken2022) across all models, matching the GPT-2–derived token counts reported for this dataset.
+To compare all models, we employ a suite of widely used NLP benchmarks, including ARC-E and ARC-C (for scientific reasoning and reading comprehension) (clark2018think), HellaSwag (for commonsense inference) (zellers2019hellaswag), PIQA (for physical reasoning) (bisk2020piqa), OBQA (for open-book science reasoning) (mihaylov2018can), SIQA (for social interaction understanding) (sap2019socialiqa), and Winogrande (for coreference and commonsense reasoning) (sakaguchi2021winogrande).
+In addition, we evaluate long-context retrieval capabilities using the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER (hsieh2024ruler), which measures a model’s ability to recover passkeys from large distractor contexts at sequence lengths ranging from 2k to 64k tokens.
+All evaluations are conducted with the widely adopted LM Evaluation Harness from EleutherAI (gao2021framework), consistent with prior work (gu2023mamba).
+For all models, we report performance in terms of benchmark accuracy [^10] Additionally, all reported perplexity values in this paper refer to training perplexity..
+When available, we use the normalized-accuracy metric (acc_norm) from the LM Evaluation Harness.
+To mitigate variance due to training stochasticity, each model’s score on a given NLP benchmark is computed as the mean accuracy over its final three checkpoints (at 90B, 95B, and 100B tokens).
+Complete results across these checkpoints, along with summary statistics and discussion, are provided in Appendix B.
 Table 2: Summary of studies for key design choices and corresponding experimental references.
 Question Answer Experiments
-RMSNorm or LayerNorm? RMSNorm Appendix H
-LR: to decay or not to decay? Yes, cosine decay with peak LR of $1\text{e}{-3}$ Appendix I
-Best values for sequence length $N$, split size $S$, and top-$k$ splits? $N=512$, $S=64$, $k=7$ Appendix G
-Activation in the enricher? Yes, ReLU2 Appendix B
-Activation in the contextualizer? No Appendix C
-Deeper model and narrower embeddings, or shallower model and wider embeddings? Deeper model and narrower embeddings Appendix F
-Weight ranked splits? Yes, using normalized scores Appendix K
-Enrich embeddings before contextualization? Yes, by 4x Appendix K, D
-Bypass uncontextualized features to the fuser? Yes, 50% of each enriched embedding Appendix K, E
-Static or dynamic parametrization for the contextualizer? Dynamic parametrization Appendix K
-Replace the contextualizer with self-attention? No Appendix K
-Besides enabling extrapolation past the context window, does the ranker improve performance? Yes Appendix K
-We conducted over 200 experiments to investigate and ablate several key design choices in our architecture.
-These included, whether to: (1) utilize activation functions in the enricher and contextualizer; (2) trade off additional layers for reduced embedding dimensionality (or vice versa) under a fixed parameter budget; (3) apply weights to ranked splits; (4) enrich embeddings; (5) bypass uncontextualized features to the fuser; (6) use dynamic parameterization in the neural processor; and (7) replace the neural contextualizer with self-attention.
-Furthermore, we evaluated Avey without the ranker to assess its contribution to downstream task performance, beyond its primary role of enabling effective extrapolation past the trained context window.
-Table 1 summarizes our findings and provides references to the corresponding experiments that support each conclusion.
-3.3 Short-Range Benchmark Results
-Table 2: Zero-shot performance across multiple NLP tasks.
-Model ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
-Avey-153M 24.37 42.33 39.36 31.40 68.37 39.13 51.28 42.32
-Transformer++-152M 23.63 43.17 39.32 29.80 67.01 38.89 50.22 41.72
-Mamba-144M 24.17 43.53 40.55 30.40 68.32 39.41 52.72 42.73
-RWKV-7-168M 24.17 43.01 41.55 29.67 68.72 39.17 51.09 42.48
-Avey-496M 27.50 48.95 51.82 32.47 72.49 40.15 54.38 46.82
-Transformer++-488M 26.73 48.09 52.66 31.73 72.13 39.93 55.25 46.65
-Mamba-500M 28.64 51.02 54.15 34.47 73.03 40.84 55.49 48.23
-RWKV-7-501M 27.13 49.37 54.54 36.27 73.58 39.40 55.72 48.00
-Avey-1.52B 31.26 56.55 61.42 36.80 75.61 42.00 57.06 51.53
-Transformer++-1.5B 30.00 56.29 63.87 38.00 76.01 42.24 61.38 52.54
-Mamba-1.4B 32.13 57.74 63.74 36.85 76.19 42.00 60.40 52.72
-RWKV-7-1.5B 32.94 59.05 64.43 37.13 76.84 41.71 60.06 53.17
-We now evaluate Avey on standard autoregressive language modeling benchmarks, comparing it against Transformer++, Mamba, and RWKV-7 across three model sizes, small, medium, and large, as defined in Section 3.1.
-We utilize a range of widely used zero-shot downstream evaluation tasks, all detailed in Section 3.1.
-Table 2 summarizes the results.
-With small models, Avey, Mamba, and RWKV-7 outperformed Transformer++ by average margins of 1.43%, 2.41%, and 1.82%, respectively.
-Mamba and RWKV-7 slightly exceeded Avey’s performance, with average margins of 0.9% and 0.3%, respectively.
-With medium models, Avey, Mamba, and RWKV-7 again outperformed Transformer++ by averages of 0.3%, 3.4%, and 2.9%, respectively.
-Lastly, with large models, Avey underperformed Transformer++ by an average of 1.9%, while Mamba and RWKV-7 marginally outpaced it by 0.71% and 1.19%, respectively.
-The results above assume a fixed training budget of 100B tokens.
-To better understand how Avey scales with increasing model size, we conducted additional experiments following the Chinchilla scaling laws [35], which recommend increasing the number of training tokens proportionally with model size.
-Consequently, we adjusted the number of training steps and tokens to align with these laws.
-Appendix J outlines the configurations of the trained models, including the numbers of layers, embedding dimensions, training steps, learning rates, and total training tokens.
-The methodology of these experiments closely follows that of [25], with slight modifications (e.g., to accommodate parameter budget constraints).
-As demonstrated in Appendix J, Avey exhibits more favorable scaling behavior than Transformer++, Mamba, and RWKV-7, particularly when both model size and token count are increased proportionally in accordance with the Chinchilla scaling laws.
-Finally, to provide deeper insights into the behaviors of the compared models, we performed a detailed statistical analysis, presented in Appendix L.
-The analysis reports the standard deviation, standard error, and 95% confidence interval for each model, computed over its final three checkpoints—captured at 5-billion-token intervals (i.e., at 90B, 95B, and 100B tokens).
-The results revealed meaningful variance among models.
-For example, in the small model regime, although Avey does not surpass Mamba in mean performance, their confidence intervals overlap substantially within the range (42.06, 43.24), indicating that the two models are statistically comparable and that Avey may outperform Mamba in some runs.
-A similar, though narrower, overlap is observed between Avey and RWKV-7.
-In the large model regime, while Avey does not exceed Transformer++ in average performance, their confidence intervals also exhibit significant overlap, suggesting that Avey could potentially outperform Transformer++ in certain cases.
-A more limited overlap is seen between Avey and Mamba, implying that although Mamba generally performs better, Avey may outperform it in some runs.
-3.4 Long-Range Benchmark Results
-Figure 5: Performance comparison between Transformer++, Mamba, RWKV-7, and Avey on S-NIAH-1 and S-NIAH-2.
-The x-axis denotes the lengths of haystacks (i.e., documents with distractor texts, varying from 2k to 64k tokens).
-All models use 0.5B parameters.
-Similar results are shown in Appendix M for other model sizes.
-In this section, we evaluate Avey, Transformer++, Mamba, and RWKV-7 on benchmarks designed to assess performance on tasks with long-range dependencies.
-Specifically, we use the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER[38], as described in Section3.1.
-The S-NIAH suite includes multiple variants, notably S-NIAH-1 (pass-key retrieval) and S-NIAH-2 (number in haystack).
-S-NIAH-1 involves retrieving the specific value associated with a given key (the pass-key) from a lengthy distractor text (the haystack) containing many key-value pairs.
-The pass-key, serving as the needle, appears only once, and the model must accurately recall its corresponding value regardless of its position in the haystack.
-S-NIAH-2 is similar to S-NIAH-1 but poses a greater challenge, whereby the value to be retrieved is numerical (e.g., a random 9-digit number).
-This task requires exact recall, where even a single-digit error is considered incorrect, thereby testing the model’s precision in extracting structured numerical information from long haystacks.
-Fig.
-5 demonstrates the results of Avey, Transformer++, Mamba, and RWKV-7 on both S-NIAH-1 and S-NIAH-2 benchmarks.
-As described in Section 3.1, Transformer++, Mamba, and RWKV-7 were all trained with a context window of 2,048 tokens.
-As shown, Transformer++ performs strongly on both benchmarks as long as the haystack length remains within its trained context window.
-Once the haystack’s length exceeds its context window width, Transformer++ fails to recall the correct values associated with the keys.
-In contrast, Mamba and RWKV-7 demonstrate some ability to generalize beyond their training windows, but their performance also declines significantly as the haystack length increases far beyond those limits.
-On the flip side, Avey achieves good performance across both benchmarks, despite being trained with a context window of only 512 tokens.
-For instance, on S-NIAH-2 with a 64k-token haystack, Avey outperforms Mamba and RWKV-7 by averages of 85.25% and 23.6%, respectively.
-In addition, on S-NIAH-1 under the same 64k-token setting, Avey achieves an accuracy of 97.8%, while Mamba and RWKV-7 drop to 0% and 0.8%, respectively.
-Interestingly, Avey’s performance tends to improve as the haystack length increases, highlighting its strong extrapolative capability.
-This behavior can be attributed to the fact that as the haystack length (i.e., sequence length $N$) grows, the candidate pool from which the ranker selects the top-$k$ splits for contextualization also expands.
-As discussed in Appendix G, a larger $N$ enables the ranker to identify and retrieve more relevant splits while discarding less informative ones, thereby improving the overall quality of contextualization and potentially enhancing performance.
-This effect is further supported by the results in Appendix K, where the inclusion of the ranker led to measurable performance gains.
-Notably, embeddings containing a needle—whether in S-NIAH-1 or S-NIAH-2—are not processed in isolation but rather contextualized alongside other embeddings.
-As such, an improved quality of contextualization driven by the ranker may contribute to more accurate value recall.
-However, whether this mechanism fully explains Avey’s increasing performance with longer haystacks remains uncertain, and further interpretability studies are needed to better understand the underlying drivers of this behavior.
-4 Related Work
-Extensive research has focused on addressing the limitations of the Transformer and exploring alternative RNN-like architectures.
-Appendix O provides a comprehensive survey of these efforts and offers a comparison with Avey.
-5 Conclusion
-In this paper, we introduced Avey, a new foundational architecture for autoregressive language modeling.
-Unlike traditional models, Avey relies neither on recurrence nor attention.
-Instead, it employs a neural approach to enrich and contextualize embeddings.
-Additionally, it leverages a ranker that enables the model to flexibly and effectively handle sequences of arbitrary lengths, despite being trained with only a small context window.
-We hope this work lays the groundwork for future research and inspires further advances in scalable and effective language modeling.
-Limitations
-We define the scope of our work and discuss limitations in Appendix P.
-Appendix A Experimental Methodology
-In this section, we describe the experimental methodology employed throughout the paper.
-To begin with, we adopted a cascaded search process to identify the best configuration of Avey.
-Specifically, we began with a baseline version of the neural processor—excluding the ranker—and sequentially explored several architectural design choices.
-After each empirical conclusion regarding a specific architectural element, we integrated that element into the processor (one at a time) and resumed the search process from the updated configuration.
-This cascaded process is captured chronologically in Appendices B, C, D, E, and F.
-To elaborate, we started with an expansion factor of 4×4$\times$4 × in the enricher, a tail size of 50% (i.e., half of each expanded embedding is forwarded to the contextualizer), RMSNorm [108] as a normalization technique, no activation functions in the enricher and contextualizer, global batch size of 0.5M, context width of 1024, and a constant learning rate of $1e{-3}$.
-As we empirically verified and decided upon each architectural element, we updated the processor accordingly.
-For example, after determining that ReLU2 was the most effective activation function for the enricher, we integrated it into the model and proceeded with the remaining search.
-After finalizing the above exploratory set of experiments, we incorporated the ranker into the neural processor and conducted an extensive study—comprising over 138 training and inference runs—to identify the optimal sequence length (i.e., $N$), split size (i.e., $S$), and number of top-$k$ splits (i.e., $k$).
-The results of this sensitivity study are summarized in Appendix G.
-Following this, we evaluated the best normalization technique (Appendix H) as well as the optimal peak learning rate and learning rate schedule (Appendix I) for the full architecture.
-All the experiments described above were conducted using a 145-million-parameter model trained on 10 billion tokens from the FineWeb dataset [^6] This dataset is released under the Open Data Commons Attribution License (ODC-By) v1.0.
-[41] (specifically, the sample-100BT subset of FineWeb).
-The results of these experiments informed the following final selection of training and model hyperparameters for Avey across three parameter scales, 153M (small), 496M (medium), and 1.52B (large).
-Training and Model Hyperparameters of Avey:
-  • Training hyperparameters:
-    • Optimizer: AdamW
-    • Betas: (0.9, 0.95)
-    • Epsilon: $1\text{e}{-12}$
-    • Peak learning rate: $1\text{e}{-3}$
-    • Schedule: Cosine decay to 10% of the peak learning rate, with no warmup
-    • Batch size: 0.5M for the small and medium models, and 1M for the large model
-    • Gradient norm clip: 1.0
-    • Weight decay: 0.1 (applied only to matrices)
-  • Model hyperparameters:
-    • All models:
-      • Context width: 512
-      • Split size ($S$): 64
-      • Number of top-$k$ splits: 7
-      • Vocabulary size: 50,304
-      • Expansion factor: 4
-      • Tail size: 0.5
-    • Small model (153M parameters):
-      • Embedding dimension: 768
-      • Number of layers: 26
-    • Medium model (496M parameters):
-      • Embedding dimension: 768
-      • Number of layers: 104
-    • Large model (1.52B parameters):
-      • Embedding dimension: 2048
-      • Number of layers: 48
-For baselines, we compared Avey against three leading open-source models, namely, Mamba [42], RWKV-7 [43], and Transformer++ [46].
-For Transformer++, we implemented the strongest architectural recipe known to us, incorporating rotary positional encodings [89], SwiGLU MLPs [84], and RMSNorm in place of LayerNorm [108].
-All models were trained using their best-known hyperparameters (see details below, assuming three model sizes, small, medium, and large) under a fixed budget of 100 billion tokens drawn from the aforementioned FineWeb dataset  [^7] More precisely, all models, including Avey, Transformer++, Mamba, and RWKV-7 were trained for 1 epoch over the sample-100BT subset of FineWeb..
-For consistency and comparability, we used the p50k_base tokenizer [64] across all the models, as it aligns with the GPT-2-derived token counts reported for this dataset.
-Training and Model Hyperparameters of Transformer++:
-  • Training hyperparameters:
-    • Optimizer: AdamW
-    • Betas: (0.9, 0.95)
-    • Epsilon: $1\text{e}{-12}$
-    • Peak learning rates:
-      • Small model: $3\text{e}{-3}$
-      • Medium model: $1.5\text{e}{-3}$
-      • Large model: $1.25\text{e}{-3}$
-    • Schedule: A linear warmup for 10% of steps, followed by cosine decay to 10% of the peak learning rate
-    • Batch size: 0.5M for the small and medium models, and 1M for the large model
-    • Gradient norm clip: 1.0
-    • Weight decay: 0.1 (only applied to matrices)
-  • Model hyperparameters:
-    • All models:
-      • Context width: 2048
-      • Vocabulary size: 50,304
-      • Intermediate size in FFN: 4$\times$ the embedding dimension
-    • Small model (152M parameters):
-      • Embedding dimension: 768
-      • Number of layers: 12
-      • Number of heads: 12
-    • Medium model (488M parameters):
-      • Embedding dimension: 1024
-      • Number of layers: 26
-      • Number of heads: 16
-    • Large model (1.5B parameters):
-      • Embedding dimension: 1664
-      • Number of layers: 32
-      • Number of heads: 16
-Training and Model Hyperparameters of Mamba:
-  • Training hyperparameters:
-    • Optimizer: AdamW
-    • Betas: (0.9, 0.95)
-    • Epsilon: $1\text{e}{-12}$
-    • Peak learning rates:
-      • Small model: $3\text{e}{-3}$
-      • Medium model: $1.5\text{e}{-3}$
-      • Large model: $1.0\text{e}{-3}$
-    • Schedule: A linear warmup for 10% of steps, followed by cosine decay to 10% of the peak learning rate
-    • Batch size: 0.5M for the small and medium models, and 1M for the large model
-    • Gradient norm clip: 1.0
-    • Weight decay: 0.1 (applied only to matrices)
-  • Model hyperparameters:
-    • All models:
-      • Context width: 2048
-      • Vocabulary size: 50,304
-      • All hyperparameters other than the ones specified are left at their default values according to [42]
-    • Small model (144M parameters):
-      • Embedding dimension: 768
-      • Number of layers: 28
-    • Medium model (500M parameters):
-      • Embedding dimension: 1280
-      • Number of layers: 42
-    • Large model (1.4B parameters):
-      • Embedding dimension: 2048
-      • Number of layers: 52
-Training and Model Hyperparameters of RWKV-7:
-  • Training hyperparameters:
-    • Optimizer: AdamW
-    • Betas: (0.9, 0.95)
-    • Epsilon: $1\text{e}{-12}$
-    • Peak learning rates:
-      • Small model: $6\text{e}{-4}$
-      • Medium model: $4\text{e}{-4}$
-      • Large model: $4\text{e}{-4}$
-    • Schedule: Cosine decay to 10% of the peak learning rate, with no warmup
-    • Batch size: 1M for the small and medium models, and 2M for the large model
-    • Gradient norm clip: 1.0
-    • Weight decay: 0.1 (applied only to matrices)
-  • Model hyperparameters:
-    • All models:
-      • Context width: 2048
-      • Vocabulary size: 50,304
-      • All hyperparameters other than the ones specified are left at their default values according to [43]
-    • Small model (168M parameters):
-      • Embedding dimension: 768
-      • Number of layers: 12
-    • Medium model (501M parameters):
-      • Embedding dimension: 1024
-      • Number of layers: 30
-    • Large model (1.5B parameters):
-      • Embedding dimension: 2048
-      • Number of layers: 24
-To compare all models, we employed a suite of widely used NLP benchmarks, including ARC-E and ARC-C (for scientific reasoning and reading comprehension) [11], HellaSwag (for commonsense inference) [105], PIQA (for physical reasoning) [5], OBQA (for open-book science reasoning) [61], SIQA (for social interaction understanding) [82], and Winogrande (for coreference and commonsense reasoning) [80].
-In addition, we evaluated long-context retrieval capabilities using the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER [38], which measures a model’s ability to extract pass-keys from large distractor corpora, with sequence lengths ranging from 2k to 64k tokens.
-All evaluations were conducted using the widely adopted LM Evaluation Harness from EleutherAI [23], consistent with the prior work in the field.
-For all models, we reported performance in terms of benchmark accuracy [^8] Additionally, throughout the paper, all reported perplexity values specifically refer to training perplexity..
-Specifically, we used normalized accuracy (acc-norm) from the LM Evaluation Harness whenever available.
-For each model, the reported score on each NLP benchmark is the average accuracy across its final three checkpoints (taken at 90B, 95B, and 100B tokens) to account for variability due to training randomness.
-Complete benchmark results across these checkpoints, along with key summary statistics and discussions, are provided in Appendix L.
-Finally, all training and evaluation runs were executed on 208 NVIDIA H200 GPUs, with mixed-precision (bfloat16) enabled for training.
-The total training time for all models—Avey, Transformer++, Mamba, and RWKV-7—across the three presented model sizes and 100B training tokens is estimated to be approximately 80–90 hours, assuming optimal parallelization across the 208 GPUs and using the implementations referenced above.
-To avoid potential sources of randomness and ensure consistency across results, we disabled Torch Compile during all design choice and sensitivity experiments.
-For the ablation studies and final training runs, however, Torch Compile was enabled whenever possible to accelerate training.
-Additionally, we fixed the random seed to 11 (arbitrarily chosen) for all training runs to further reduce variability due to stochastic effects.
-Appendix B Activation or No Activation in the Enricher
-Table 3: Avey’s performance without and with an activation function in the enricher. The study involves only the neural processor within Avey and trains it on 10B tokens.
+Activation in the enricher? Yes, ReLU2 Section 3.2.1
+Activation in the contextualizer? No Section 3.2.2
+Enrich embeddings before contextualization? Yes, by 4x Sections 3.3, 3.2.3
+Bypass uncontextualized features to the fuser? Yes, 50% of each enriched embedding Sections 3.3, 3.2.4
+Deeper model and narrower embeddings, or shallower model and wider embeddings? Deeper model and narrower embeddings Section 3.2.5
+Best values for sequence length $N$, split size $S$, and top-$k$ splits? $N=512$, $S=64$, $k=7$ Section 3.2.6
+RMSNorm or LayerNorm? RMSNorm Section 3.2.7
+LR: to decay or not to decay? Yes, cosine decay with peak LR of $1\text{e}{-3}$ Section 3.2.8
+Weight ranked splits? Yes, using normalized scores Section 3.3
+Static or dynamic parametrization for the contextualizer? Dynamic parametrization Section 3.3
+Replace the contextualizer with self-attention? No Section 3.3
+Finally, all training and evaluation runs were executed on 208 NVIDIA H200 GPUs with mixed precision (bfloat16) enabled for training.
+Across the three model sizes and a 100B-token budget, the total wall-clock training time was approximately 80–90 hours under near-ideal parallelization using the above referenced implementations.
+To minimize sources of randomness and ensure comparability, we disabled torch.compile during all design-choice and sensitivity experiments.
+For ablations and final training runs, we enabled torch.compile when compatible to accelerate training.
+We also fixed the random seed to 11 (arbitrarily chosen) for all runs to further reduce variability due to stochastic effects.
+3.2 Design Choices
+We conducted more than 200 experiments to investigate key design choices.
+Table 2 summarizes the main findings and cites the corresponding experiments that support each conclusion.
+We next elaborate on all these experiments.
+3.2.1 Activation or No Activation in the Enricher
+Table 3: Avey’s performance without and with an activation function in the enricher. The study isolates the neural processor (excluding the ranker) and trains it on 10B tokens.
 Configuration Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
@@
 SiLU 30.81 22.18 38.43 31.30 28.20 62.30 37.05 53.20 38.95
-In this study, we illustrate the performance of Avey with and without an activation function in the enricher.
-To this end, we trained a model with 145 million parameters using 10 billion tokens from the Fineweb dataset [41].
-The model employs an expansion factor of 4×4$\times$4 × in the enricher (i.e., each embedding dimension is expanded fourfold by the enricher), a tail size of 50% (i.e., half of each expanded embedding is forwarded to the contextualizer), RMSNorm [108] as a normalization technique, and no activation function in the contextualizer.
-Additionally, the context width (i.e., the maximum number of tokens that can be input to and processed by the contexutalizer simultaneously) is set to 1024 and a constant learning rate of $1e{-3}$ is maintained throughout training.
-The study excludes the ranker and focuses solely on the neural processor.
-Besides, it evaluates four activation functions, namely, GELU [33], ReLU [62], ReLU2 [10], and SiLU [76].
-All other experimental settings follow the methodology detailed in Appendix A.
-Table 3 summarizes the results.
-As shown, ReLU2 yielded an improvement in performance versus a baseline with no activation, hence, was adopted as the default activation function for the enricher throughout our experiments presented in Sections 3.3 and 3.4.
-It is important to note, however, that the lowest perplexity was provided by GELU and not ReLU2 (although the difference in perplexity was very minimal).
-While perplexity quantifies how effectively the model predicts the next token in the training dataset, it remains a proxy for overall modeling capability and does not always precisely predict downstream task performance.
-Appendix C Activation or No Activation in the Contextualizer
+In this study, we evaluate Avey’s performance with and without an activation function in the enricher.
+We train a 145M-parameter model on 10B tokens from the FineWeb dataset (fineweb2023).
+The configuration uses an enricher expansion factor of $4\times$ (each embedding is expanded fourfold), a tail size of $50\%$ (i.e., half of each expanded embedding is forwarded to the contextualizer), RMSNorm (zhang2019rmsnorm) for normalization, and no activation in the contextualizer.
+The context width (i.e., the maximum number of tokens processed concurrently by the contextualizer) is set to 1024, and we employ a constant learning rate of 1e31e{-3} throughout training.
+This study excludes the ranker and focuses solely on the neural processor.
+Besides, it evaluates four activation functions in the enricher, namely, GELU (hendrycks2016gelu), ReLU (nair2010relu), $\mathrm{ReLU}^{2}$ (chowdhery2022palm), and SiLU (ramachandran2017swish).
+All other experimental settings follow the methodology in Section 3.1.
+Results are summarized in Table 3.
+As shown, $\mathrm{ReLU}^{2}$ yields a clear accuracy improvement over the no-activation baseline and is therefore adopted as the default enricher activation for Avey.
+Notably, GELU attains the lowest perplexity, though the gap relative to $\mathrm{ReLU}^{2}$ is minimal.
+While perplexity measures next-token predictive quality on the training distribution, it is an indirect proxy for modeling capability and does not always precisely predict downstream task performance.
+3.2.2 Activation or No Activation in the Contextualizer
 Table 4: Avey’s performance without and with an activation function in the contextualizer.
-The study involves only the neural processor within Avey, trains it on 10B tokens, and uses ReLU2 within the enricher, capitalizing on the results shown in Table 3.
+The study isolates the neural processor (without the ranker), trains on 10B tokens, and fixes $\mathrm{ReLU}^{2}$ in the enricher, following the findings in Table 3.
 Configuration Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
@@
 SiLU 30.92 23.63 36.74 31.51 27.40 64.20 36.18 50.04 38.53
-We now evaluate Avey with and without an activation function in the contextualizer.
-We use the same experimental settings outlined in Appendix B and add to that ReLU2 as an activation function in the enricher, capitalizing on the findings therein.
-We also experiment with four activation functions, namely, GELU, ReLU, ReLU2, and SiLU.
-Results are summarized in Table 4.
-As illustrated, the best performance was achieved without any activation function in the contextualizer, thus was employed as the default configuration in all our experiments reported in Sections 3.3 and 3.4.
-Appendix D What is the Best Expansion Factor?
+Next, we evaluate Avey both with and without an activation function in the contextualizer.
+We employ the same experimental setup as in Section 3.2.1 and, consistent with the findings reported therein, fix $\mathrm{ReLU}^{2}$ as the activation for the enricher.
+We then assess four candidate activations for the contextualizer, namely, GELU, ReLU, $\mathrm{ReLU}^{2}$, and SiLU.
+The results, summarized in Table 4, indicate that the best performance is achieved without any activation in the contextualizer.
+Accordingly, we adopt this configuration as Avey’s default setting.
+3.2.3 What is the Best Expansion Factor?
 Table 5: The effect of the expansion factor on Avey’s performance.
-The study involves only the neural processor within Avey, trains it on 10B tokens, adopts ReLU2 within the enricher, and does not use an activation function within the contextualizer, building upon the results portrayed in Tables 3 and 4.
+The study isolates the neural processor (without the ranker), trains on 10B tokens, fixes $\mathrm{ReLU}^{2}$ as the enricher activation, and uses no activation in the contextualizer, following the configurations established in Tables 3 and 4.
 Expansion Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
@@
 30.00 23.21 37.79 31.48 26.20 63.82 36.59 50.75 38.55
-In this study, we vary the expansion factor in the enricher, defined as the degree to which each input embedding is expanded.
-Specifically, we evaluate several expansion factors, ranging from $2x$ to $8x$ as shown in Table 5, while keeping the total model parameter count constant (e.g., with a $2x$ expansion factor we use 34 layers, while with a $4x$ one we utilize 20 layers).
-The experimental setup follows the methodology outlined in Appendix B, but incorporates ReLU2 as an activation function in the enricher and omits any activation function in the contextualizer, aligning with the findings reported in Appendices B and C.
-As depicted in the table, an expansion factor of $4x$ yielded the best performance, hence, was set as the default configuration in the enricher throughout all our experiments presented in Sections 3.3 and 3.4.
-Appendix E What is the Best Tail Size?
+In this study, we investigate the effect of the expansion factor in the enricher, defined as the degree to which each input embedding is expanded.
+Specifically, we evaluate expansion factors ranging from $2\times$ to $8\times$ (see Table 5) while maintaining a fixed total parameter budget (e.g., $2\times$ uses 34 layers, whereas $4\times$ uses 20 layers).
+The experimental setup follows that of Section 3.2.1, employing $\mathrm{ReLU}^{2}$ in the enricher (chowdhery2022palm) and no activation in the contextualizer, consistent with the findings from Sections 3.2.1 and 3.2.2.
+As shown in Table 5, an expansion factor of $4\times$ delivers the strongest performance and is therefore adopted as the default configuration for the enricher.
+3.2.4 What is the Best Tail Size?
 Table 6: The effect of the tail size on Avey’s performance.
-The study involves only the neural processor within Avey, trains it on 10B tokens, adopts ReLU2 within the enricher, does not use an activation function within the contextualizer, and utilizes an expansion factor of $4x$, as recommended by the findings demonstrated in Tables 34, and 5.
+The study isolates the neural processor (without the ranker), trains on 10B tokens, fixes $\mathrm{ReLU}^{2}$ as the enricher activation, uses no activation in the contextualizer, and adopts an expansion factor of $4\times$, following the findings in Tables 3, 4, and 5.
 Tail Size Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
@@
 We now examine the impact of forwarding a tail portion of each enriched embedding to the contextualizer.
-More precisely, we vary the size of this tail portion, referred to as the tail size, from 10% to 90% of each enriched embedding, as illustrated in Table 6.
-The study follows the experimental setup described in Appendix B and employs ReLU2 as an activation function in the enricher, no activation function in the contextualizer, and an expansion factor of $4x$, based on the findings presented in Appendices B, C, and D, respectively.
-As depicted in Table 6, the best performance was accomplished using a tail size of 50%, thus was adopted as the default configuration for Avey in all our experiments reported in Sections 3.3 and 3.4.
-Appendix F Deeper Models and Narrower Embeddings, or Shallower Models and Wider Embeddings
-Table 7: Avey’s performance across different model configurations, including wider embedding dimensions (e.g., 1536 under 0.5B-parameter model) with shallower layers (e.g., 24 layers under 0.5B-parameter model), or narrower embedding dimensions (e.g., 768 under 0.5B-parameter model) with deeper layers (e.g., 90 layers under 0.5B-paramter model).
-The models with 140M, 0.5B, and 1.5B parameters are referred to as small Avey, medium Avey, and large Avey in the text.
+Concretely, we vary the tail size (or the fraction of each enriched embedding passed to the contextualizer) from $10\%$ to $90\%$ (see Table 6).
+The study follows the setup outlined in Section 3.2.1, using $\mathrm{ReLU}^{2}$ in the enricher, no activation in the contextualizer, and an expansion factor of $4\times$, consistent with the findings in Sections 3.2.1, 3.2.2, and 3.2.3.
+As depicted in Table 6, a tail size of $50\%$ yields the strongest performance and is accordingly employed as Avey’s default configuration.
+3.2.5 Deeper Models and Narrower Embeddings, or Shallower Models and Wider Embeddings
+Table 7: Avey’s performance across configurations that trade embedding width for depth, including wider embeddings with fewer layers (e.g., 1536-dimensional embeddings with 24 layers for the 0.5B-parameter model) versus narrower embeddings with more layers (e.g., 768-dimensional embeddings with 90 layers for the 0.5B-parameter model).
+Throughout the text, the 140M-, 0.5B-, and 1.5B-parameter models are referred to as small Avey, medium Avey, and large Avey, respectively.
 # Params Embed. # Layers Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
@@
 2560 30 20.23 26.62 45.16 43.91 29.2 69.10 39.82 52.09 43.70
-Table 8: Avey’s performance under different sequence lengths, $N$, split sizes, $S$, and top-$k$ values.
-All models (a total of 69) were trained on 10B tokens using 140 million parameters.
-The sweet spot in terms of downstream task performance was at $N=512$, $S=64$, and $k=7$; hence, it was adopted as Avey’s default configuration.
-$N$ $S$ $k$ Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
+The objective of this study is to determine whether a narrower embedding dimension with a greater model depth yields better or worse performance than a wider embedding dimension with fewer layers.
+The study utilizes the experimental setup described in Section 3.2.1 and leverages the findings of Sections 3.2.13.2.23.2.3, and 3.2.4.
+Consequently, it employs $\mathrm{ReLU}^{2}$ as the activation function in the enricher, no activation in the contextualizer, an expansion factor of $4\times$, and a tail size of $50\%$.
+We begin by evaluating a 140M-parameter Avey model (referred to as small Avey) with three embedding dimensions, 512, 768, and 1024.
+To maintain a constant parameter budget, these dimensions yielded 40, 20, and 11 layers, respectively.
+Results are summarized in Table 7.
+As shown, the configuration with a 768-dimensional embedding and 20 layers outperforms the other two.
+We next evaluate a 500M-parameter Avey model (denoted as medium Avey) with embedding dimensions of 768, 1024, and 1536.
+To keep the total parameter count fixed, these settings resulted in 90, 54, and 24 layers, respectively.
+As reported in Table 7, the configuration with a 768-dimensional embedding and 90 layers achieves the strongest performance among the three.
+Lastly, we examine a 1.5B-parameter Avey model (referred to as large Avey) with embedding dimensions of 1536, 2048, and 2560.
+Again, to maintain a constant parameter budget, these settings yielded 80, 48, and 30 layers, respectively.
+As illustrated in Table 7, the configuration with a 2048-dimensional embedding and 48 layers delivers the best performance among the three.
+The above results suggest a trend, whereby wider embedding dimensions (e.g., 1024 in small Avey; 1563 in medium Avey; and 2560 in large Avey) paired with shallower architectures (e.g., 11 layers in small Avey; 24 layers in medium Avey; and 30 layers in large Avey) tend to underperform deeper models (e.g., 40 and 20 layers in small Avey; 90 and 54 layers in medium Avey; and 80 and 48 layers in large Avey) with narrower embeddings (e.g., 512 and 768 in small Avey; 768 and 1024 in medium Avey; and 1536 and 2048 in large Avey).
+As such, in all our subsequent experiments, we adopt deeper models with narrower embedding dimensions, namely, the best performing small Avey, medium Avey, and large Avey in Table 7.
+Interestingly, Table 7 indicates that some benchmarks benefit more from increased model capacity than others.
+For example, the commonsense reasoning task HellaSwag shows gains of $20.36\%$ and $28.5\%$ under the 0.5B- and 1.5B-parameter models, respectively, relative to the 140M-parameter baseline.
+In contrast, the question–answering benchmark SIQA exhibits only a modest improvement of $2.4\%$ for both the 0.5B- and 1.5B-parameter models, suggesting lower sensitivity to model size.
+3.2.6 What are the Best Sequence Length, Split Size, and Top-$k$ Values?
+Table 8: Avey’s performance under varying sequence lengths (NN), split sizes ($S$), and top-$k$ values.
+The study trains 69 models with 140M parameters on 10B tokens.
+The strongest downstream performance is achieved at $N=512$, $S=64$, and $k=7$.
+NN $S$ $k$ Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
 256 16 1 43.59 21.16 37.67 30.86 27.20 64.20 37.72 51.85 38.66
@@
 512 1 28.24 24.40 39.39 32.98 28.20 65.13 37.72 50.59 39.77
-The objective of this study is to determine whether a narrower embedding dimension with a greater model depth yields better or worse performance than a wider embedding dimension with fewer layers.
-The study utilizes the experimental setup described in Appendix B and leverages the findings presented in Appendices BCD, and E.
-Consequently, it utilizes ReLU2 as an activation function in the enricher, no activation function in the contextualizer, an expansion factor of $4x$, and a tail size of 50%.
-To begin with, we evaluated a small Avey model (referred to as small Avey) with 140 million parameters, using three different embedding dimensions, 512, 768, and 1024.
-These configurations resulted in 40, 20, and 11 layers, respectively, to maintain a constant parameter count.
-Table 7 summarizes the results.
-As illustrated, the configuration with an embedding dimension of 768 and a layer count of 20 outperformed the other two configurations.
-Afterwards, we assessed a larger Avey model with 500 million parameters (referred to as medium Avey), using three different embedding dimensions, 768, 1024, and 1536.
-These configurations resulted in 90, 54, and 24 layers, respectively, while keeping the total parameter count constant.
-As shown in Table 7, the setup with an embedding dimension of 768 and 90 layers yielded the best performance among the three tested ones.
-Finally, we examined an even larger Avey model with 1.5 billion parameters (referred to as large Avey), using three different embedding dimensions, 1536, 2048, and 2560.
-To maintain a constant parameter count across configurations, these dimensions corresponded to 80, 48, and 30 layers, respectively.
-As portrayed in Table 7, the configuration with an embedding dimension of 2048 and 48 layers delivered the best performance among the three considered configurations.
-The above results suggest a trend, whereby wider embedding dimensions (e.g., 1024 in small Avey; 1563 in medium Avey; and 2560 in large Avey) paired with shallower architectures (e.g., 11 layers in small Avey; 24 layers in medium Avey; and 30 layers in large Avey) tend to underperform deeper models (e.g., 40 and 20 layers in small Avey; 90 and 54 layers in medium Avey; and 80 and 48 layers in large Avey) with narrower embeddings (e.g., 512 and 768 in small Avey; 768 and 1024 in medium Avey; and 1536 and 2048 in large Avey).
-As such, in all our experiments discussed in Sections 3.3 and 3.4, we employed deeper models with narrower embedding dimensions, namely, the best performing small Avey, medium Avey, and large Avey in Table 7.
-Interestingly, Table 7 also highlights that certain benchmarks benefit more from increased model capacity than others.
-For instance, a commonsense reasoning benchmark like HellaSwag demonstrates performance improvements of 20.36% and 28.5% under 0.5B-parameter and 1.5B-parameter models, respectively, compared to a 140M-parameter baseline.
-In contrast, a question-answering benchmark such as SIQA exhibits only a modest gain of 2.4% under both 0.5B-parameter and 1.5B-parameter models relative to the 140M-parameter baseline, suggesting less sensitivity to model size.
-Appendix G What are the Best Sequence Length, Split Size, and Top-$k$ Values?
-We now analyze how Avey’s perplexity and overall performance are affected by variations in three key parameters, the ranker’s top-$k$ selected splits, the split size $S$, and the sequence length $N$.
-We consider three values for $N$, $256$, $512$, and $1024$.
-For each $N$, the split size $S$ is grown geometrically, starting from $16$ and doubling at each step, up to the maximum permissible value $N/2$.
-Subsequently, for any $N$ and $S$, the number of top-ranked splits $k$ can range from $1$ (i.e., contextualizing the current split with one additional relevant split) up to $N/S-1$.
-To tame the quantity of experiments, we increase $k$ arithmetically from $1$ to a maximum of $15$, whenever possible, using an increment of $2$.
-Finally, we use the experimental configurations recommended in Appendices BCDE, and F.
+We now analyze how Avey’s perplexity and overall performance vary with three key parameters, the ranker’s top-$k$ selected splits, the split size $S$, and the sequence length $N$.
+We consider $N\in\{256,512,1024\}$.
+For each $N$, the split size $S$ grows geometrically from $16$, doubling at each step up to the maximum permissible value $N/2$.
+Subsequently, for any $(N,S)$, the number of top-ranked splits $k$ ranges from $1$ (i.e., contextualizing the current split with one additional relevant split) up to N/S1N/S-1.
+To keep the experiment count tractable, we vary $k$ arithmetically from $1$ to at most $15$, in increments of $2$ whenever feasible.
+All runs use the best configurations established in Sections 3.2.1, 3.2.2, 3.2.3, 3.2.4, and 3.2.5.
 As shown in Table 8, Avey’s perplexity is highest when both $S$ and $k$ are very small (e.g., $S=16$ and $k=1$).
-While a small $S$ can help filter out irrelevant embeddings and denoise contextualization, pairing it with a very small $k$ can deprive the contextualizer of sufficient context to build expressive representations [^9] Recall from Section 2.2.2 that the context width $C$ is defined as $C=S(k+1)$..
-To expand the context (i.e., increase its width $S(k+1)$) and enrich the resulting embeddings, either $S$ or $k$ can be increased.
-For example, as we increased $k$ under $N=256$ and $S=16$, perplexity decreased and benchmark performance improved.
-However, a larger context does not always yield better outcomes, especially when involving a high proportion of irrelevant embeddings.
-This behavior was observed when $k$ was increased beyond $3$ under $N=1024$ and $S=128$, resulting in higher perplexity and diminished downstream accuracy.
-In contrast to $S$ and $k$, the sequence length $N$ determines the size of the candidate pool from which the ranker selects the top-$k$ splits for any current split during training.
-A larger $N$ allows the ranker to reach farther back in the sequence history, potentially retrieving more relevant splits and lowering perplexity.
-For example, increasing $N$ from $512$ to $1024$, while holding $S=16$ and $k=15$ constant, reduced perplexity by $5.4\%$.
-Nevertheless, lower perplexity does not always translate into improved downstream performance.
-For instance, with $N=512$ and $S=16$, $k=15$ yielded the lowest perplexity, yet $k=5$ achieved higher benchmark accuracy.
-As noted in Appendix B, the loss remains only a proxy for overall modeling capability and does not always exactly predict downstream task performance.
-As demonstrated in Table 8, the best empirical performance was obtained with $N=512$, $S=64$, and $k=7$, which was, accordingly, adopted as Avey’s default configuration.
-Appendix H RMSNorm or LayerNorm
-Table 9: Avey’s performance with RMSNorm and LayerNorm.
-A model with 153 million parameters was trained on 10B tokens using both the neural processor and ranker with the best configuration from Table 8.
+Although a small $S$ can help filter out irrelevant embeddings and denoise contextualization, pairing it with a very small $k$ can starve the contextualizer of sufficient signal to form expressive representations [^11] Recall from Section 2.2.2 that the context width $C$ is defined as $C=S(k+1)$..
+To expand the effective context (i.e., increase $S(k+1)$) and enrich the resulting representations, one can increase either $S$ or $k$.
+For example, under $N=256$ and $S=16$, increasing $k$ monotonically reduces perplexity and improves benchmark accuracy.
+However, a larger context does not invariably yield better outcomes, especially when it admits a higher proportion of irrelevant embeddings.
+This effect appears when $k>3$ for $N=1024$ and $S=128$, where perplexity rises and downstream accuracy degrades.
+Unlike $S$ and $k$, the sequence length $N$ controls the size of the candidate pool from which the ranker selects top-$k$ splits for the current split during training.
+A larger $N$ allows the ranker to reach further back in the sequence history, potentially retrieving more relevant splits and lowering perplexity.
+For example, increasing $N$ from $512$ to $1024$ while holding $S=16$ and $k=15$ fixed reduces perplexity by $5.4\%$.
+Nevertheless, lower perplexity does not always translate into higher downstream accuracy.
+With $N=512$ and $S=16$, for instance, $k=15$ attains the lowest perplexity, yet $k=5$ yields better benchmark accuracy.
+As noted in Section 3.2.1, loss remains a proxy for overall modeling capability and does not always precisely predict downstream performance.
+As shown in Table 8, the strongest empirical results are obtained with $N=512$, $S=64$, and $k=7$, which we accordingly adopt as Avey’s default configuration.
+3.2.7 RMSNorm or LayerNorm
+Table 9: Avey’s performance with RMSNorm versus LayerNorm.
+The study trains a 153M-parameter model on 10B tokens using the full architecture (neural processor + ranker) with the best configuration from Table 8.
 Normalization Method Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
@@
 LayerNorm 30.93 23.55 39.65 33.24 28.8 65.07 38.28 49.57 39.74
-In all previous runs across the appendices, we used RMSNorm [108] as a normalization technique.
-We now test Avey with another standard normalization method, namely, LayerNorm [1].
-To begin with, we note that Avey normalizes input embeddings before each layer (as illustrated in Fig.
-3) and output embeddings once after the final layer and prior to token prediction.
-We evaluate each normalization technique using Avey’s complete architecture, including its neural processor and ranker.
-The ranker is configured according to the best-performing setting identified in Appendix G (i.e., sequence length $N=512$, split size $S=64$, and number of top-ranked splits $k=7$).
-The neural processor employs the optimal configurations reported in Appendices BCDE, and F.
-As before, we trained the resulting model with 153 million parameters on 10 billion tokens from the Fineweb dataset.
-Table 9 summarizes the results.
-As shown, RMSNorm slightly outperforms LayerNorm on average, and is therefore adopted in Avey and used consistently across all the experiments presented in Sections 3.3 and 3.4.
-Appendix I Learning Rate: To Decay or Not to Decay?
-Table 10: Avey’s performance with two types of schedules, constant learning rate (LR) and cosine decay, starting from different peak learning rates.
-All models involved the neural processor and ranker, and were trained with 153 million parameters on 10B tokens, using both the neural processor and ranker with the best configuration from Table 8.
+In all preceding studies, we employed RMSNorm (zhang2019rmsnorm) as the normalization method.
+We now compare Avey against an alternative standard, that is, LayerNorm (ba2016layer).
+Recall that Avey applies pre-layer normalization to the input embeddings at every layer (as illustrated in Fig.
+3) and a single post-layer normalization to the output embeddings after the final layer, immediately before token prediction.
+We evaluate each normalization method using Avey’s full architecture, comprising both the neural processor and the ranker.
+The ranker is fixed to the best-performing setting from Section 3.2.6 (i.e., sequence length $N=512$, split size $S=64$, and top-$k=7$).
+The neural processor adopts the optimal configurations identified in Sections 3.2.1, 3.2.2, 3.2.3, 3.2.4, and 3.2.5.
+As before, we train a 153M-parameter model on 10B tokens from the FineWeb dataset.
+Results are summarized in Table 9.
+On average, RMSNorm slightly outperforms LayerNorm and is subsequently used in Avey by default.
+3.2.8 Learning Rate: To Decay or Not to Decay?
+Table 10: Avey’s performance under two learning-rate schedules, constant and cosine decay, across varying peak learning rates.
+All models use the full architecture (neural processor + ranker) with the best configuration from Table 8, and are trained with 153M parameters on 10B tokens.
 Schedule LR Perplexity ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
@@
 1e-03 25.00 23.21 41.12 34.76 27.0 65.67 38.38 50.75 40.13
-In all previous runs across the appendices, we used a constant learning rate of $1e{-3}$.
-In this study, we evaluate the performance of Avey under varying maximum learning rates and learning rate schedules.
-Specifically, we compare two schedules, cosine decay and constant learning rate (which can be viewed as cosine decay with an infinite cycle length, effectively eliminating any decay).
-The study adopts the experimental configurations suggested in Appendices BCDE, and F.
-In addition, it employs Avey’s complete architecture, including its neural processor and ranker, with the best configuration from Table 8.
-As shown in Table 10, cosine decay consistently achieves lower losses than the constant schedule across the tested learning rates.
-This observation aligns with findings from the Chinchilla paper [35], which indicates that when the cosine cycle length significantly exceeds the total number of training steps (by at least 25%), model performance tends to deteriorate.
-Notably, the longest cycle length arises when the schedule is constant.
-In contrast, setting it to approximately match the training duration yields the best final loss [35].
-To this end, we adopt a cosine decay schedule with a peak learning rate of $1e{-3}$ for Avey, especially that it provides the lowest loss across all the runs.
-We note, however, that Table 10 also shows that the lowest loss does not correspond to the best downstream task performance.
-For instance, the loss of 3.308 under the constant learning rate schedule resulted in a slightly better benchmark performance than the lower loss of 3.218 under the cosine decay schedule, both using the same peak learning rate.
-While constant learning rates can be effective for short or exploratory runs (this study uses only 10 billion tokens), it is generally the case that, as the number of training tokens increases, the learning rate must decrease to allow the optimizer to settle into a lower-loss region [103].
-Hence, schedules with decay are typically favored for longer or large-scale training runs [35, 4].
-Appendix J Scaling Laws
+In all prior experiments, we used a constant learning rate of 1e31e{-3}.
+In this study, we assess Avey’s performance under varying peak learning rates and learning-rate schedules.
+Specifically, we compare a cosine-decay schedule against a constant schedule (i.e., cosine decay with effectively infinite cycle length, yielding no decay).
+The study adopts the experimental configurations established in Sections 3.2.1, 3.2.2, 3.2.3, 3.2.4, and 3.2.5.
+In addition, we utilize Avey’s full architecture, encompassing the neural processor and ranker, with the best configuration from Table 8.
+As shown in Table 10, cosine decay consistently achieves lower loss than a constant schedule across the learning rates tested.
+This trend aligns with the Chinchilla findings (hoffmann2022training), which report that when the cosine cycle length substantially exceeds the total number of training steps (by at least $25\%$), performance degrades.
+Notably, a constant schedule corresponds to an effectively infinite cycle length.
+In contrast, setting the cycle length to approximately match the training duration yields the best final loss (hoffmann2022training).
+To this end, we adopt a cosine-decay schedule with a peak learning rate of 1e31e{-3} for Avey, as it yields the lowest training loss across all the runs.
+We note, however, that Table 10 also shows a mismatch between loss and downstream accuracy.
+For example, a higher loss of $3.308$ under the constant schedule produced slightly better benchmark performance than a lower loss of $3.218$ under cosine decay at the same peak rate.
+While constant learning rates can be effective for short or exploratory runs (this study uses 10B tokens), it is generally the case that, as the number of training tokens increases, the learning rate should decrease to allow the optimizer to settle into a lower-loss region (you2019does).
+Consequently, schedules with decay are typically preferred for longer or large-scale training (hoffmann2022training; bergsma2025straight).
+3.3 Ablation Study
+Table 11: Ablation results comparing Avey variants, with individual components removed or substituted.
+Model Variant Perplexity ARC-C ARC-E Hella OBQA PIQA SIQA Wino Average
+Avey full (all features) 30.00 25.17 39.90 33.59 28.8 65.56 37.62 51.62 40.32
+Avey without dynamic parameterization 34.31 25.00 40.66 32.99 28.8 65.34 36.64 50.51 39.99
+Avey without bypassing 32.55 22.61 38.38 32.31 28.0 64.20 38.28 52.09 39.41
+Avey without embedding expansion 39.94 22.44 37.92 28.75 25.4 62.40 38.64 52.01 38.22
+Avey without weighting selected splits 31.17 22.78 38.55 33.25 28.0 65.89 37.82 52.09 39.77
+Avey without the ranker 29.48 23.72 38.59 32.52 28.0 63.66 37.67 53.20 39.62
+Avey with self-attention in place of neural proc. 31.39 22.61 39.27 31.99 28.0 64.58 38.33 51.38 39.45
+In this section, we conduct a series of ablations on Avey’s core components, building on the best configurations identified in Section 3.2.
+All experiments use Avey’s full architecture, comprising the neural processor and ranker, with the small model (153M parameters) as the baseline.
+We train this model on 10B tokens from the FineWeb dataset, following the training methodology and hyperparameters detailed in Section 3.1.
+We begin by examining the effect of dynamic parameterization on both perplexity and downstream benchmark performance.
+As described in Section 2.2.2, dynamic parameterization enables each neuron in the contextualizer to compute a cosine similarity between its input embedding and the embeddings associated with other neurons, weight those embeddings accordingly, and aggregate them via a learned weighted sum.
+This mechanism induces selectivity (gu2023mamba) in the neural processor, making its parameterization input-dependent.
+Table 11 summarizes the results.
+Disabling this component increases perplexity by $14.3\%$ and reduces average downstream performance by $0.8\%$, highlighting its importance.
+Second, we assess the impact of partial-embedding bypassing introduced in Section 2.2.1.
+This mechanism forwards a fraction of each expanded embedding directly to the fuser, preserving raw, distinctive features that can promote more diverse representations.
+As reported in Table 11, removing this component increases perplexity by $8.5\%$ and reduces average downstream performance by $2.2\%$, underscoring its significance.
+Third, we set the enricher’s expansion factor to $1$, effectively disabling embedding expansion.
+As illustrated in Table 11, this change increases perplexity by $33.1\%$ and reduces average downstream performance by $5.2\%$, corroborating the critical role of embedding expansion in the model’s effectiveness.
+Fourth, we remove the weighting of each selected split by its corresponding normalized MaxSim score, thereby preventing the contextualizer from scaling each split’s contribution during contextualization.
+As depicted in Table 11, this adjustment increases perplexity by $3.8\%$ and reduces average downstream performance by $1.37\%$, indicating the importance of this technique.
+Fifth, we evaluate Avey without the ranker to assess its impact on downstream performance beyond its primary role of enabling effective extrapolation outside the trained context window.
+As shown in Table 11, the ranker improves the neural processor’s performance, primarily by enhancing contextualization through more meaningful cross-token interactions.
+We note, however, a slight increase in loss ($0.5\%$), which again highlights (as in Section 3.2.1) the discrepancy between the objectives of pertaining and downstream tasks.
+Finally, we replace Avey’s neural processor with a self-attention module to assess the relative contribution of each component to overall performance, given that both are designed to facilitate cross-token interactions.
+As illustrated in Table 11, this substitution increases perplexity by $4.6\%$ and reduces average downstream performance by $2.1\%$, underscoring the centrality of the neural processor and suggesting that self-attention is less effective within Avey’s architectural framework.
+3.4 Short-Range Benchmark Results
+Table 12: Zero-shot performance across multiple NLP tasks.
+Model ARC-C ARC-E HellaSwag OBQA PIQA SIQA Winogrande Average
+Avey-153M 24.37 42.33 39.36 31.40 68.37 39.13 51.28 42.32
+Transformer++-152M 23.63 43.17 39.32 29.80 67.01 38.89 50.22 41.72
+Mamba-144M 24.17 43.53 40.55 30.40 68.32 39.41 52.72 42.73
+RWKV-7-168M 24.17 43.01 41.55 29.67 68.72 39.17 51.09 42.48
+Avey-496M 27.50 48.95 51.82 32.47 72.49 40.15 54.38 46.82
+Transformer++-488M 26.73 48.09 52.66 31.73 72.13 39.93 55.25 46.65
+Mamba-500M 28.64 51.02 54.15 34.47 73.03 40.84 55.49 48.23
+RWKV-7-501M 27.13 49.37 54.54 36.27 73.58 39.40 55.72 48.00
+Avey-1.52B 31.26 56.55 61.42 36.80 75.61 42.00 57.06 51.53
+Transformer++-1.5B 30.00 56.29 63.87 38.00 76.01 42.24 61.38 52.54
+Mamba-1.4B 32.13 57.74 63.74 36.85 76.19 42.00 60.40 52.72
+RWKV-7-1.5B 32.94 59.05 64.43 37.13 76.84 41.71 60.06 53.17
+In this section, we evaluate Avey on standard autoregressive language-modeling benchmarks and compare it against Transformer++, Mamba, and RWKV-7 across three parameter scales, small, medium, and large (see Table 1).
+We adopt a suite of widely used zero-shot downstream tasks, as outlined in Section 3.1.
+Results are summarized in Table 12.
+For the small models, Avey, Mamba, and RWKV-7 outperform Transformer++ by average margins of $1.43\%$, $2.41\%$, and $1.82\%$, respectively.
+Additionally, Mamba and RWKV-7 slightly exceed Avey by $0.90\%$ and $0.30\%$.
+For the medium models, Avey, Mamba, and RWKV-7 again surpass Transformer++ by $0.30\%$, $3.40\%$, and $2.90\%$, respectively.
+For the large models, Avey underperforms Transformer++ by $1.90\%$, while Mamba and RWKV-7 edge out Avey by $0.71\%$ and $1.19\%$, respectively.
+The results above assume a fixed training budget of 100B tokens.
+To analyze Avey’s scaling behavior with respect to model size, we conduct additional experiments following the Chinchilla scaling laws (hoffmann2022training), which recommend increasing the number of training tokens in proportion to model size.
+Accordingly, we adjust the training steps and token budgets to align with these prescriptions.
+All results are reported in Appendix A, which also details the model configurations (e.g., number of layers and embedding dimensions), optimization settings (e.g., learning rates), training schedules, and total tokens for all evaluated models.
+Our methodology largely follows gu2023mamba, with minor modifications to satisfy parameter-budget constraints.
+Results in Appendix A show that Avey exhibits the strongest scaling behavior, followed by Transformer++, Mamba, and RWKV-7, in that order.
+3.5 Long-Range Benchmark Results
+Figure 5: Performance comparison between Transformer++, Mamba, RWKV-7, and Avey on S-NIAH-1 and S-NIAH-2.
+The x-axis denotes the lengths of haystacks (i.e., documents with distractor texts, varying from 2k to 64k tokens).
+All models use 0.5B parameters.
+Similar results are shown in Appendix C for other model sizes.
+We now evaluate Avey, Transformer++, Mamba, and RWKV-7 on benchmarks targeting long-range dependencies.
+Specifically, we use the standard Single Needle-In-A-Haystack (S-NIAH) suite from RULER (hsieh2024ruler), as described in Section 3.1.
+The suite includes multiple variants, notably S-NIAH-1 and S-NIAH-2.
+S-NIAH-1 requires retrieving the value associated with a given key (the pass-key) from a distractor context (the haystack) containing many key–value pairs.
+The pass-key (or the needle) appears exactly once, and the model must recover its corresponding value regardless of position within the haystack.
+S-NIAH-2 follows the same structure but is more challenging, whereby the target value is numeric (e.g., a random 9-digit number), demanding exact recall in which any single-digit error is scored as incorrect.
+Consequently, this variant tests precise extraction of structured numerical information from long contexts.
+Fig.
+5 demonstrates the results for Avey, Transformer++, Mamba, and RWKV-7 on both S-NIAH-1 and S-NIAH-2.
+As noted in Section 3.1, Transformer++, Mamba, and RWKV-7 were all trained with a 2,048-token context window.
+Transformer++ performs strongly on both benchmarks while the haystack length remains within this window.
+Once the haystack length exceeds the window width, the accuracy of Transformer++ collapses.
+Mamba and RWKV-7 generalize somewhat beyond their training windows, but their performance also degrades substantially as the haystack grows.
+In contrast, Avey maintains high accuracy on both benchmarks despite being trained with a 512-token context window.
+For example, on S-NIAH-2 with a 64k-token haystack, Avey outperforms Mamba and RWKV-7 by average margins of $85.25\%$ and $23.6\%$, respectively.
+On S-NIAH-1 under the same 64k-token setting, Avey attains $97.8\%$ accuracy, whereas Mamba and RWKV-7 drop to $0\%$ and $0.8\%$, respectively.
+Interestingly, Avey’s performance often improves as the haystack length increases, highlighting strong extrapolative capability.
+One plausible explanation is that as the haystack (or the sequence length $N$) grows, the candidate pool from which the ranker selects top-$k$ splits also expands.
+As discussed in Section 3.2.6, a larger $N$ enables the ranker to retrieve more relevant splits while discarding less informative ones, improving the quality of contextualization and potentially boosting accuracy.
+This interpretation is consistent with findings in Section 3.3, where introducing the ranker yielded measurable gains.
+Notably, embeddings containing a needle (whether in S-NIAH-1 or S-NIAH-2) are not processed in isolation but are contextualized alongside other embeddings.
+As such, better ranker-driven context selection can result in more precise value recall.
+That said, whether this fully explains Avey’s increasing accuracy with longer haystacks remains an open question, and further interpretability studies are needed to better understand the underlying drivers of this behavior.
+4 Related Work
+Recurrent Neural Networks (RNNs) (elman1990finding; rumelhart1986learning) are designed to process sequential data by capturing temporal dependencies, making them well-suited for tasks where input order is essential.
+However, their cyclical nature limits their potential for parallel computation and exposes them to vanishing and exploding gradient problems.
+As a result, they typically struggle to effectively learn long-range dependencies.
+While architectures like Long Short-Term Memory (LSTM) (hochreiter1997long) and Gated Recurrent Units (GRU) (cho2014learning) mitigate these gradient-related issues, they remain slow to optimize and challenging to scale due to retaining RNN’s core recurrent and sequential structure.
+In contrast, the Transformer (vaswani2017attention) employs a self-attention mechanism to process each sequence of tokens simultaneously.
+More precisely, it promotes two key design principles: (1) a recurrent-free architecture, which enables parallel computation of token embeddings, while still capturing their order through positional encodings, and (2) a multi-head self-attention approach, which facilitates cross-token interactions to further enrich the expressiveness of embeddings.
+These innovations make the Transformer highly effective, as well as parallelizable and efficient to train.
+However, they also cause its computational and memory requirements to scale quadratically with sequence length, making it expensive and less efficient for very long sequences.
+To address the Transformer’s quadratic computation and memory costs, a wide array of approaches have been proposed, including linear attention (kitaev2020reformer; katharopoulos2020transformers; choromanski2020rethinking; peng2021random; zhai2021attention), sparse or local attention (yuan2025native; child2019generating; parmar2018image), context compression (rae2019compressive; wang2020linformer; sukhbaatar2019adaptive; roy2021efficient), and modified attention computations (tay2021synthesizer; wu2019pay; tay2020sparse), to mention just a few.
+Notably, the Attention-Free Transformer (AFT) (zhai2021attention) offers a linear drop-in replacement for the quadratic self-attention mechanism.
+In particular, it weights key and value vectors using learned positional biases and integrates them with query vectors via element-wise multiplication.
+As such, it eliminates the need to compute and store the costly attention matrix while still preserving global query-value interactions—without requiring architectural modifications or additional tuning.
+Furthermore, AFT introduces variants such as AFT-local and AFT-conv, which leverage local attention patterns to reduce parameter count and further improve computational and memory efficiency.
+RWKV-4 (peng2023rwkv) (the first 3 versions were experimental (li2024survey)) capitalizes on AFT and suggests combining the strengths of both Transformers and RNNs.
+To elaborate, unlike Transformers and akin to RNNs, it does not process each input token solely based on its own embedding, but rather as a weighted sum of its embedding and that of the preceding one.
+To the contrary of traditional RNNs and similar to Transformers, it adopts self-attention, but an extended version of it, namely, that of AFT.
+This hybrid approach allows RWKV-4 to maintain some of the modeling capabilities of RNNs, while benefiting from the parallelization and scalability features of Transformers.
+More precisely, RWKV-4 extends AFT in two distinct ways: (1) it introduces an additional parameter to handle each current token independently, and (2) it implements a per-time-step decay mechanism that selectively removes older content from the recurrent hidden state in a data-dependent manner.
+This decay mechanism addresses a central limitation of linear attention, which pertains to the lack of a systematic way to discard outdated information (schlag2021linear; yang2023gated).
+Architecturally, RWKV-4 consists of homogeneous stacked residual blocks, each encompassing two units, a time-mixing and a channel-mixing ones.
+The time-mixing unit applies linear attention across tokens, while the channel-mixing unit integrates each element of the current token’s embedding with its corresponding element from the preceding token embedding, leveraging the output of the time-mixing unit.
+RWKV-5 (peng2024eagle) enhances RWKV-4’s architecture and learning decay mechanism by replacing traditional vector-valued states with more expressive multi-head, matrix-valued ones.
+Moreover, it reconfigures receptive states, incorporates supplementary gating mechanisms, and dynamically learns the linear interpolation between the current and preceding token embeddings instead of relying on pre-defined hyperparameters.
+RWKV-6 (peng2024eagle) promotes a new application of low-rank adaptation functions (hu2022lora; li2024survey).
+Specifically, it makes the linear interpolation between the current and preceding tokens data-dependent to improve the selectivity of the model in retaining and discarding information.
+Additionally, it replaces the static decay mechanism with a dynamic one, allowing each element in the decay vector to fluctuate independently over time in response to the input.
+The decay strategies of RWKV-4, RWKV-5, and RWKV-6 still cannot remove values stored at specific keys.
+DeltaNet (schlag2021linear) overcomes this drawback by partially replacing the values stored at current keys with equivalent new values, enabling models to erase outdated memories and include up-to-date ones on a per-key basis.
+However, it only allows a fixed scalar fraction of a value to be replaced from a state via an in-context learning rate parameter, thus demonstrating rigidity in adapting to varying data contexts.
+RWKV-7 (peng2025rwkv) builds upon the principles of DeltaNet and introduces a vector-valued in-context learning rate instead of a scalar-valued one.
+This allows selective replacement of state data on a channel-wise basis.
+Furthermore, RWKV-7 employs a vector-valued decay mechanism and uses additional low-rank projections to optimize the trade-off between the number of parameters, computational efficiency, and downstream performance.
+Lastly, it incorporates Value Residual Learning (zhou2024value), which improves the propagation of initial local information via utilizing a residual connection between the value vectors of the current layer and those of the first layer prior to the attention operation, resulting in enhanced language modeling performance.
+Most recently, RWKV-X (hou2025rwkv) proposed combining the strengths of RWKV and sparse attention, drawing inspiration from Mixture of Block Attention (MoBA) (lu2025moba).
+In particular, RWKV-X restricts each query to attend only to a small, relevant subset of the input, thus reducing computational cost and facilitating the modeling of longer-range dependencies.
+More precisely, rather than allowing each token to attend to every other token in the sequence (as in traditional self-attention), it constrains each token’s attention to a limited subset (hence, making it sparse), while maintaining the coupling between the input sequence and context window.
+Similar to RWKV, RetNet (sun2023retentive) adopts linear attention and promotes a hybrid approach that blends Transformer- and RNN-like representations, yet with a decay-based memory unit.
+Specifically, it divides the input sequence into chunks, wherein the Transformer-like parallel representation is applied.
+Additionally, it enables propagating information sequentially across chunks using the RNN-like representation.
+Lastly, it uses multiple attention heads, each governed by a distinct decay rate, and replaces LayerNorm (ba2016layer) with GroupNorm (wu2018group).
+Although linear attention has been proposed as a promising alternative to quadratic softmax attention (katharopoulos2020transformers; choromanski2020rethinking; kasai2021finetuning; peng2021random), existing implementations of it are in practice slower than optimized versions of softmax attention (dao2022flashattention; dao2023flashattention; yang2023gated).
+From an accuracy standpoint, linear attention generally underperforms conventional softmax attention, sometimes by a significant margin in language modeling (kasai2021finetuning; yang2023gated).
+To this end, and in light of the exponentially growing complexity associated with overcoming the limitations of Transformer-based architectures, there has been a renewed interest in RNN-based alternatives in recent years.
+Notably, Structured State Space Sequence (S4) models (gu2021efficiently; gu2021combining), inspired by the classical state space models (SSMs) (kalman1960new), have emerged as a promising paradigm for sequence modeling.
+These models describe the temporal evolution of a system using differential equations, offering a continuous-time formulation of dynamics, and can be viewed as generalized versions of RNNs.
+An SSM as a concept has a broad meaning, which simply refers to the notion of any recurrent process with a latent state (gu2023mamba).
+From this perspective, the RNN-like linear attention model proposed and formulated by (katharopoulos2020transformers) can be interpreted as a degenerate linear SSM.
+Interestingly, this justifies the usage of a decay factor in RetNet and RWKV, especially that a decay term (or a forget gate) has been shown to be crucial in RNNs (hochreiter1997long; van2018unreasonable; cho2014learning).
+Numerous variants of SSMs (gu2021efficiently; gu2022parameterization; gupta2022diagonal; li2024makes; ma2022mega; orvieto2023resurrecting; smith2022simplified) have demonstrated strong performance across a range of domains, including audio and vision (goel2022s; nguyen2022s4nd; saon2023diagonal).
+Nonetheless, these variants have struggled with language modeling, often lagging behind Transformers by several points in perplexity (gu2021efficiently).
+From an efficiency standpoint, however, SSMs have shown encouraging results in language modeling.
+For instance, S4 (gu2021efficiently; gu2021combining), a prominent SSM, converts the continuous-time state update equation of SSMs into a discrete form, hence, enabling parallel sequence modeling.
+Moreover, it utilizes the HiPPO (High-Order Polynomial Projection Operator) initialization (gu2020hippo), which alleviates the vanishing gradient problem and facilitates processing longer sequences.
+Another example of SSMs is H3 (fu2022hungry), which improves language modeling by allowing both, the recall of earlier tokens and token-wise comparisons within a sequence.
+H3 extends S4 by suggesting a state-passing algorithm that enhances computational efficiency on modern hardware.
+This advancement reduces the hardware-related barriers that have traditionally limited the scalability of SSM-based architectures.
+Hyena (poli2023hyena) capitalizes on H3 by replacing its S4 layer with an MLP-parameterized global convolution (romero2102ckconv).
+S5 (smith2022simplified) proposes using parallel scan (martin2017parallelizing) to parallelize S4.
+Liquid S4 (hasani2022liquid) augments S4 with an input-dependent state transition matrix, computed convolutionally in the frequency domain (which is computationally efficient) and mapped back to the time domain using an inverse Fourier transformation.
+SGConv (li2024makes), LongConv (fu2023simple), MultiresConv (shi2023sequence), and Toeplitz Neural Network (qin2023toeplitz) all focus on the convolutional representation of S4 as well, aiming to enhance its efficiency (gu2023mamba).
+Most recently, Mamba (gu2023mamba) introduced a new class of SSMs known as selective SSMs, specifically designed to improve the performance of language modeling tasks.
+Mamba addresses a key limitation in SSMs, namely, their inability to selectively process inputs in an input-dependent manner (i.e., focus on or ignore specific parts of the input sequence).
+Consequently, it makes the SSM parameters input-dependent, but introduces a technical challenge since traditional SSMs are inherently designed to be time- and input-invariant to ensure computational efficiency.
+To overcome this challenge, Mamba proposes a hardware-efficient parallel scan (or prefix sum) algorithm (blelloch1990prefix), which enables recurrent-style computation without explicitly materializing the expanded state.
+This design precludes costly I/O operations across GPU memory hierarchies and accelerates both, training and inference.
+Tri Dao and Albert Gu (dao2024transformers) argue that various approaches to operating SSMs can be reframed as matrix multiplication algorithms involving a specific class of structured matrices known as semiseparable matrices.
+They further leverage the language of tensor contractions to prove the recurrent formulation of linear attention as proposed by (katharopoulos2020transformers), before generalizing it to a new family of structured masked attention (SMA).
+Subsequently, Tri Dao and Albert Gu demonstrated that SSMs and Transformers are fundamentally connected, governed by the mathematical framework of semiseparable matrices and SMA.
+Additionally, they developed a rich state space duality (SSD) framework of theoretical connections between SSMs and various forms of attention.
+This framework facilitated the design of Mamba-2, an extended version of Mamba, which aims to improve its efficiency (not performance).
+Mamba-2 utilizes a scalar data-dependent gating mechanism (like the ones proposed by (peng2021random; sun2023retentive; beck2024xlstm)), which enables transforming its recurrent structure into a matrix-multiply form, thus allowing for efficient execution on tensor cores and better support for larger hidden state sizes.
+The strategy of the SSD framework mirrors that of linear attention (katharopoulos2020transformers), which established a connection between autoregressive attention mechanisms and linear RNNs via showing an equivalence between "dual forms" of quadratic kernelized attention and a specific type of linear recurrence.
+Conceptually, the SSD framework seeks to transfer algorithmic and systems-level optimizations originally developed for Transformers to the realm of SSMs.
+Its overarching goal is to enable the development of architectures that outperform Transformers, while scaling more efficiently with sequence length.
+Finally, several works, including Tolstikhin et al.
+(tolstikhin2021mlp), Melas-Kyriazi (melas2021you), Touvron et al.
+(touvron2022resmlp), and Ding et al.
+(ding2021repmlp), among others, have questioned the necessity of self-attention, particularly in the context of Vision Transformers.
+In contrast, Liu et al.
+(liu2021pay) introduced gMLP, an MLP-based alternative to BERT-style Transformers (devlin2019bert) that (partially) eliminates self-attention but ultimately underperforms average performance on downstream NLP tasks.
+gMLP encompasses channel (hidden) and spatial (cross-token) projections with multiplicative gating and static parameterization.
+Its gating mechanism is reminiscent of Gated Linear Units (GLUs) (dauphin2017language; shazeer2020glu; wu2019pay), as well as earlier architectures such as Highway Networks (srivastava2015highway) and LSTM-RNNs (hochreiter1997long).
+A key distinction, however, is that gMLP applies gating on the spatially projected dimension and not the hidden one.
+The gated embedding-wise neural network in Avey’s contextualizer draws inspiration from gMLP.
+Unlike all previously mentioned models, Avey abandons self-attention and recurrence, introducing a new architecture composed of a ranker and a dynamically parameterized neural processor.
+The ranker identifies the most relevant tokens for contextualization, while the neural processor contextualizes them data-dependently.
+This design decouples sequence length from context width, enabling efficient processing of arbitrarily long sequences without diminishing the influence of distant yet important tokens.
+At the core of Avey’s architecture is a weighted-selective-split interaction mechanism, which filters out irrelevant tokens beyond the context window and enables direct interactions only with relevant ones, thus preserving their influence irrespective of sequence length.
+In addition, Avey employs a partial-embedding bypassing technique that retains a portion of each token’s raw, distinctive features before fusing them with its contextualized ones through a neural network.
+This technique boosts the performance of Avey (as shown in Appendix 3.3) and might help mitigate issues such as entropy collapse (zhai2023stabilizing) and over-smoothing (zhou2021deepvit; shi2022revisiting), especially at large-scale, when the depth of the model is increased significantly.
+5 Conclusion
+In this paper, we introduce Avey, a new foundational architecture for autoregressive language modeling.
+Unlike traditional models, Avey relies on neither recurrence nor attention.
+Instead, it employs a neural mechanism to enrich and contextualize embeddings, coupled with a ranker that enables flexible and effective handling of arbitrarily long sequences despite being trained with a small context window.
+We hope this work lays the groundwork for future research and inspires a promising new direction for scalable, effective language modeling.
+Acknowledgments
+We gratefully acknowledge the Ministry of Communications and Information Technology (MCIT) in Qatar and NVIDIA for providing the computational resources that made many of the experiments reported herein possible.
+We also thank Microsoft Qatar for technical support on the GPU infrastructure provided by MCIT.
+Appendix A Scaling Laws
 Table 13: Model configurations used in the scaling law experiments.
@@
 Transformer++-1.5B 32 (16) 1664 1.25e-03 20B
-Mamba-153M 28 768 1.00e-03 2B
-Mamba-496M 42 1280 1.00e-03 7B
-Mamba-1.5B 52 2048 1.00e-03 20B
-RWKV-7-152M 12 768 3.00e-03 2B
-RWKV-7-488M 30 1024 1.50e-03 7B
-RWKV-7-1.5B 24 2048 1.25e-03 20B
+Mamba-153M 28 768 3.00e-03 2B
+Mamba-496M 42 1280 1.50e-03 7B
+Mamba-1.5B 52 2048 1.00e-04 20B
+RWKV-7-152M 12 768 6.00e-04 2B
+RWKV-7-488M 30 1024 4.00e-04 7B
+RWKV-7-1.5B 24 2048 4.00e-04 20B
 Figure 6: Scaling law results, comparing how perplexity decreases as compute increases, assuming three model sizes of 150M, 500M, and 1.5B parameters and a proportional increase in the number of training tokens with model size, following the Chinchilla scaling laws.
@@
 For Avey, we use the full architecture, including both the ranker and neural processor.
-All models are trained at three different sizes, as defined in Appendix A and summarized in Table 11.
-To ensure compute-optimal scaling, we proportionally increase the number of training tokens with model size, following the Chinchilla scaling laws [35].
-Specifically, and in line with the methodology from [25], we use 2B, 7B, and 20B tokens to train models with approximately 150M, 500M, and 1.5B parameters, respectively.
+All models are trained at three different sizes, as defined in Section 3.1 and summarized in Table 13.
+To ensure compute-optimal scaling, we proportionally increase the number of training tokens with model size, following the Chinchilla scaling laws [hoffmann2022training].
+Specifically, and in line with the methodology from [gu2023mamba], we use 2B, 7B, and 20B tokens to train models with approximately 150M, 500M, and 1.5B parameters, respectively.
 Lastly, we employ the same batch size across all the models to control for variability in the number of gradient update steps, especially because of training with a limited number of tokens.
@@
 For example, the small model (150M parameters) with sequence length $N=1024$, split size $S=16$, and top-$k=15$ achieves much lower perplexity as shown in Table 8.
-The configuration used in this experiment– and in Sections 3.3 and 3.4, was selected based on its strong downstream benchmark performance, rather than optimal perplexity.
-As discussed in Appendix B, perplexity serves as a useful proxy for modeling capability, but does not always align perfectly with downstream task accuracy., it improves more rapidly than the other models, indicating strong scaling behavior and greater benefit from additional compute.
+The configuration used in this experiment– and in Sections 3.4 and 3.5, was selected based on its strong downstream benchmark performance, rather than optimal perplexity.
+As discussed in Appendix 3.2.1, perplexity serves as a useful proxy for modeling capability, but does not always align perfectly with downstream task accuracy., it improves more rapidly than the other models, indicating strong scaling behavior and greater benefit from additional compute.
 Following Avey, Transformer++ demonstrates the next-best scaling trend, outpacing Mamba and RWKV-7.
@@
 Finally, RWKV-7 performs well at low compute but shows the flattest scaling curve, suggesting it gains the least from additional training compute.
-Appendix K Ablation Study
-Table 12: Ablation results comparing Avey variants, with individual components removed or replaced.
-Model Variant ppl ARC-C ARC-E Hella OBQA PIQA SIQA Wino Average
-Avey full (all features) 30.00 25.17 39.90 33.59 28.8 65.56 37.62 51.62 40.32
-Avey without dynamic parameterization 34.31 25.00 40.66 32.99 28.8 65.34 36.64 50.51 39.99
-Avey without bypassing 32.55 22.61 38.38 32.31 28.0 64.20 38.28 52.09 39.41
-Avey without embedding expansion 39.94 22.44 37.92 28.75 25.4 62.40 38.64 52.01 38.22
-Avey without weighting selected splits 31.17 22.78 38.55 33.25 28.0 65.89 37.82 52.09 39.77
-Avey without the ranker 29.48 23.72 38.59 32.52 28.0 63.66 37.67 53.20 39.62
-Avey with self-attention in place of neural proc. 31.39 22.61 39.27 31.99 28.0 64.58 38.33 51.38 39.45
-In this study, we perform a series of ablation experiments on the core components of Avey, leveraging the best configurations identified in Appendices BCDEFH, and I.
-All experiments are conducted using Avey’s complete architecture, comprising both the ranker and neural processor, with the small model variant (153 million parameters) as the baseline.
-For this study, we trained this model on 10 billion tokens from the FineWeb dataset, using the training methodology and hyperparameters detailed in Appendix A.
-We begin by examining the effect of dynamic parameterization on both perplexity and downstream benchmark performance.
-As described in Section 2.2.2, dynamic parameterization allows each neuron in the contextualizer network to compute a cosine similarity between its input embedding and the embeddings of all other neurons, weight those embeddings accordingly, and aggregate them via a learned weighted sum.
-This mechanism induces selectivity, as defined in [25], into the neural processor, thereby making its parameterization dynamic (or input-dependent).
-Table 12 summarizes the results.
-Disabling this component results in a 14.3% increase in perplexity and a 0.8% drop in average performance, highlighting its importance.
-Second, we evaluate the impact of the partial-embedding bypassing technique introduced in Section 2.2.1.
-This method involves forwarding a portion of each expanded embedding directly to the fuser, allowing raw, distinctive features to be preserved and potentially serve in promoting more diverse representations.
-As shown in Table 12, removing this mechanism results in an 8.5% increase in perplexity and a 2.2% drop in average performance, underscoring its significance.
-Third, we set the expansion factor in the enricher to 1, effectively disabling the expansion of input embeddings.
-As illustrated in Table 12, this modification results in a 33.1% increase in perplexity and a 5.2% drop in average performance, corroborating the critical role of embedding expansion in the model’s effectiveness.
-Fourth, we remove the weighting of each selected split by its corresponding normalized MaxSim score, thereby preventing the contextualizer from scaling each split’s contribution during contextualization.
-As depicted in Table 12, this adjustment leads to a 3.8% increase in perplexity and a 1.37% drop in average performance, indicating the importance of this technique.
-Fifth, we evaluate Avey without the ranker to assess its impact on downstream task performance, beyond its primary role of enabling effective extrapolation past the trained context window.
-As shown in Table 12, the ranker does indeed enhance the neural processor’s performance, primarily by improving the quality of contextualization through more meaningful cross-token interactions.
-We note, however, the slight increase in loss (by 0.5%), which again highlights (as in Appendix B) the discrepancy between the objectives of pertaining and downstream tasks.
-Finally, we replace Avey’s neural processor with self-attention to assess the relative contribution of each component to Avey’s overall performance, given that both are designed to pursue cross-token interactions.
-As illustrated in Table 12, this substitution leads to a 4.6% increase in perplexity and a 2.1% decline in average performance, underscoring the significance of the neural processor and suggesting that self-attention is less effective within Avey’s architectural framework.
-Appendix L Additional Short-Range Benchmark Results
+Appendix B Additional Short-Range Benchmark Results
 Table 14: Performance of all models across short-range benchmarks at 90B, 95B, and 100B training tokens.
@@
 53.12), the meaningful overlap in their confidence intervals implies that the difference between them is not statistically significant, and either model could outperform the other depending on minor experimental factors.
-Appendix M Additional Long-Range Benchmark Results
+Appendix C Additional Long-Range Benchmark Results
 Figure 7: Performance comparison between Transformer++, Mamba, RWKV-7, and Avey on S-NIAH-1 and S-NIAH-2.
@@
 All models use $\sim$150M parameters.
-In Section 3.4, we presented results for Avey, Transformer++, Mamba, and RWKV-7 under the medium parameter regime ($\sim$500M parameters) on the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER [38], which is designed to evaluate models’ abilities to handle long-range dependencies.
+In Section 3.5, we presented results for Avey, Transformer++, Mamba, and RWKV-7 under the medium parameter regime ($\sim$500M parameters) on the standard Single Needle-In-A-Haystack (S-NIAH) benchmark suite from RULER [hsieh2024ruler], which is designed to evaluate models’ abilities to handle long-range dependencies.
 The S-NIAH benchmark, along with two of its common variants—S-NIAH-1 and S-NIAH-2—was described in detail in Section 3.5.
@@
 In this paper, we kept Avey’s context window fixed at 512 tokens across all parameter regimes.
-All tuning experiments related to sequence length, split size, and top $k$ splits (see Section G) were conducted exclusively using the small model size.
+All tuning experiments related to sequence length, split size, and top $k$ splits (see Section 3.2.6) were conducted exclusively using the small model size.
 It is plausible that with larger capacity, Avey could more effectively leverage longer sequences by retrieving and contextualizing a larger set of relevant tokens while filtering out less informative ones, thereby enhancing contextual representations and further boosting performance.
 Investigating the relationship between sequence length and model size in Avey is an interesting direction for future work.
-Appendix N Complexity Analysis
+Appendix D Complexity Analysis
 As indicated in Sections 2.1, 2.2.1, 2.2.2, and 2.2.3, the training time complexities of the ranker, enricher, contextualizer, and fuser are $\mathcal{O}(N^{2}d)$, $\mathcal{O}(Nmd)$, $\mathcal{O}(NkCm_{t})$, and $\mathcal{O}(Nmd)$, respectively, where $N$ is the sequence length, $S$ is the split size, $d$ is the original embedding dimension, $m$ is the projected embedding dimension (with $m>d$), $m_{t}$ is the tail part of $m$ forwarded to the contextualizer, $C$ is the context width (with $C\leq N$), and $k$ is the number of splits contextualized with each current split.
@@
 However, the ranker’s analysis slightly changes, as at each time step $t$ (i.e., upon predicting a new token), the current split is compared against all previous splits.
-More precisely, at each time step $t$, the current split—denoted as split $i$ and incrementally filled as tokens are generated—is compared against all $i-1$ preceding splits, each consisting of $S$ tokens.
+More precisely, at each time step $t$, the current split—denoted as split $i$ and incrementally filled as tokens are generated—is compared against all i1i-1 preceding splits, each consisting of $S$ tokens.
 Consequently, the cost of comparing $t$ tokens in split $i$ (with $t\leq S$) against $S$ tokens in a previous split is $\mathcal{O}(t\cdot S\cdot d)$.
@@
 $\frac{\mathcal{O}(N^{2}d)}{N}=\mathcal{O}(Nd)$
-Appendix O Related Work
-Recurrent Neural Networks (RNNs) [19, 79] are designed to process sequential data by capturing temporal dependencies, making them well-suited for tasks where input order is essential.
-However, their cyclical nature limits their potential for parallel computation and exposes them to vanishing and exploding gradient problems.
-As a result, they typically struggle to effectively learn long-range dependencies.
-While architectures like Long Short-Term Memory (LSTM) [34] and Gated Recurrent Units (GRU) [8] mitigate these gradient-related issues, they remain slow to optimize and challenging to scale due to retaining RNN’s core recurrent and sequential structure.
-In contrast, the Transformer [98] employs a self-attention mechanism to process each sequence of tokens simultaneously.
-More precisely, it promotes two key design principles: (1) a recurrent-free architecture, which enables parallel computation of token embeddings, while still capturing their order through positional encodings, and (2) a multi-head self-attention approach, which facilitates cross-token interactions to further enrich the expressiveness of embeddings.
-These innovations make the Transformer highly effective, as well as parallelizable and efficient to train.
-However, they also cause its computational and memory requirements to scale quadratically with sequence length, making it expensive and less efficient for very long sequences.
-To address the Transformer’s quadratic computation and memory costs, a wide array of approaches have been proposed, including linear attention [50, 48, 9, 71, 106], sparse or local attention [104, 7, 67], context compression [74, 99, 90, 78], and modified attention computations [93, 100, 92], to mention just a few.
-Notably, the Attention-Free Transformer (AFT) [106] offers a linear drop-in replacement for the quadratic self-attention mechanism.
-In particular, it weights key and value vectors using learned positional biases and integrates them with query vectors via element-wise multiplication.
-As such, it eliminates the need to compute and store the costly attention matrix while still preserving global query-value interactions—without requiring architectural modifications or additional tuning.
-Furthermore, AFT introduces variants such as AFT-local and AFT-conv, which leverage local attention patterns to reduce parameter count and further improve computational and memory efficiency.
-RWKV-4 [68] (the first 3 versions were experimental [52]) capitalizes on AFT and suggests combining the strengths of both Transformers and RNNs.
-To elaborate, unlike Transformers and akin to RNNs, it does not process each input token solely based on its own embedding, but rather as a weighted sum of its embedding and that of the preceding one.
-To the contrary of traditional RNNs and similar to Transformers, it adopts self-attention, but an extended version of it, namely, that of AFT.
-This hybrid approach allows RWKV-4 to maintain some of the modeling capabilities of RNNs, while benefiting from the parallelization and scalability features of Transformers.
-More precisely, RWKV-4 extends AFT in two distinct ways: (1) it introduces an additional parameter to handle each current token independently, and (2) it implements a per-time-step decay mechanism that selectively removes older content from the recurrent hidden state in a data-dependent manner.
-This decay mechanism addresses a central limitation of linear attention, which pertains to the lack of a systematic way to discard outdated information [83, 102].
-Architecturally, RWKV-4 consists of homogeneous stacked residual blocks, each encompassing two units, a time-mixing and a channel-mixing ones.
-The time-mixing unit applies linear attention across tokens, while the channel-mixing unit integrates each element of the current token’s embedding with its corresponding element from the preceding token embedding, leveraging the output of the time-mixing unit.
-RWKV-5 [69] enhances RWKV-4’s architecture and learning decay mechanism by replacing traditional vector-valued states with more expressive multi-head, matrix-valued ones.
-Moreover, it reconfigures receptive states, incorporates supplementary gating mechanisms, and dynamically learns the linear interpolation between the current and preceding token embeddings instead of relying on pre-defined hyperparameters.
-RWKV-6 [69] promotes a new application of low-rank adaptation functions [39, 52].
-Specifically, it makes the linear interpolation between the current and preceding tokens data-dependent to improve the selectivity of the model in retaining and discarding information.
-Additionally, it replaces the static decay mechanism with a dynamic one, allowing each element in the decay vector to fluctuate independently over time in response to the input.
-The decay strategies of RWKV-4, RWKV-5, and RWKV-6 still cannot remove values stored at specific keys.
-DeltaNet [83] overcomes this drawback by partially replacing the values stored at current keys with equivalent new values, enabling models to erase outdated memories and include up-to-date ones on a per-key basis.
-However, it only allows a fixed scalar fraction of a value to be replaced from a state via an in-context learning rate parameter, thus demonstrating rigidity in adapting to varying data contexts.
-RWKV-7 [70] builds upon the principles of DeltaNet and introduces a vector-valued in-context learning rate instead of a scalar-valued one.
-This allows selective replacement of state data on a channel-wise basis.
-Furthermore, RWKV-7 employs a vector-valued decay mechanism and uses additional low-rank projections to optimize the trade-off between the number of parameters, computational efficiency, and downstream performance.
-Lastly, it incorporates Value Residual Learning [110], which improves the propagation of initial local information via utilizing a residual connection between the value vectors of the current layer and those of the first layer prior to the attention operation, resulting in enhanced language modeling performance.
-Most recently, RWKV-X [37] proposed combining the strengths of RWKV and sparse attention, drawing inspiration from Mixture of Block Attention (MoBA) [57].
-In particular, RWKV-X restricts each query to attend only to a small, relevant subset of the input, thus reducing computational cost and facilitating the modeling of longer-range dependencies.
-More precisely, rather than allowing each token to attend to every other token in the sequence (as in traditional self-attention), it constrains each token’s attention to a limited subset (hence, making it sparse), while maintaining the coupling between the input sequence and context window.
-Similar to RWKV, RetNet [91] adopts linear attention and promotes a hybrid approach that blends Transformer- and RNN-like representations, yet with a decay-based memory unit.
-Specifically, it divides the input sequence into chunks, wherein the Transformer-like parallel representation is applied.
-Additionally, it enables propagating information sequentially across chunks using the RNN-like representation.
-Lastly, it uses multiple attention heads, each governed by a distinct decay rate, and replaces LayerNorm [1] with GroupNorm [101].
-Although linear attention has been proposed as a promising alternative to quadratic softmax attention [48, 9, 47, 71], existing implementations of it are in practice slower than optimized versions of softmax attention [14, 12, 102].
-From an accuracy standpoint, linear attention generally underperforms conventional softmax attention, sometimes by a significant margin in language modeling [47, 102].
-To this end, and in light of the exponentially growing complexity associated with overcoming the limitations of Transformer-based architectures, there has been a renewed interest in RNN-based alternatives in recent years.
-Notably, Structured State Space Sequence (S4) models [27, 28], inspired by the classical state space models (SSMs) [44], have emerged as a promising paradigm for sequence modeling.
-These models describe the temporal evolution of a system using differential equations, offering a continuous-time formulation of dynamics, and can be viewed as generalized versions of RNNs.
-An SSM as a concept has a broad meaning, which simply refers to the notion of any recurrent process with a latent state [25].
-From this perspective, the RNN-like linear attention model proposed and formulated by [48] can be interpreted as a degenerate linear SSM.
-Interestingly, this justifies the usage of a decay factor in RetNet and RWKV, especially that a decay term (or a forget gate) has been shown to be crucial in RNNs [34, 97, 8].
-Numerous variants of SSMs [27, 29, 30, 51, 58, 65, 87] have demonstrated strong performance across a range of domains, including audio and vision [24, 63, 81].
-Nonetheless, these variants have struggled with language modeling, often lagging behind Transformers by several points in perplexity [27].
-From an efficiency standpoint, however, SSMs have shown encouraging results in language modeling.
-For instance, S4 [27, 28], a prominent SSM, converts the continuous-time state update equation of SSMs into a discrete form, hence, enabling parallel sequence modeling.
-Moreover, it utilizes the HiPPO (High-Order Polynomial Projection Operator) initialization [26], which alleviates the vanishing gradient problem and facilitates processing longer sequences.
-Another example of SSMs is H3 [20], which improves language modeling by allowing both, the recall of earlier tokens and token-wise comparisons within a sequence.
-H3 extends S4 by suggesting a state-passing algorithm that enhances computational efficiency on modern hardware.
-This advancement reduces the hardware-related barriers that have traditionally limited the scalability of SSM-based architectures.
-Hyena [72] capitalizes on H3 by replacing its S4 layer with an MLP-parameterized global convolution [77].
-S5 [87] proposes using parallel scan [59] to parallelize S4.
-Liquid S4 [31] augments S4 with an input-dependent state transition matrix, computed convolutionally in the frequency domain (which is computationally efficient) and mapped back to the time domain using an inverse Fourier transformation.
-SGConv [51], LongConv [21], MultiresConv [86], and Toeplitz Neural Network [73] all focus on the convolutional representation of S4 as well, aiming to enhance its efficiency [25].
-Most recently, Mamba [25] introduced a new class of SSMs known as selective SSMs, specifically designed to improve the performance of language modeling tasks.
-Mamba addresses a key limitation in SSMs, namely, their inability to selectively process inputs in an input-dependent manner (i.e., focus on or ignore specific parts of the input sequence).
-Consequently, it makes the SSM parameters input-dependent, but introduces a technical challenge since traditional SSMs are inherently designed to be time- and input-invariant to ensure computational efficiency.
-To overcome this challenge, Mamba proposes a hardware-efficient parallel scan (or prefix sum) algorithm [6], which enables recurrent-style computation without explicitly materializing the expanded state.
-This design precludes costly I/O operations across GPU memory hierarchies and accelerates both, training and inference.
-Tri Dao and Albert Gu [13] argue that various approaches to operating SSMs can be reframed as matrix multiplication algorithms involving a specific class of structured matrices known as semiseparable matrices.
-They further leverage the language of tensor contractions to prove the recurrent formulation of linear attention as proposed by [48], before generalizing it to a new family of structured masked attention (SMA).
-Subsequently, Tri Dao and Albert Gu demonstrated that SSMs and Transformers are fundamentally connected, governed by the mathematical framework of semiseparable matrices and SMA.
-Additionally, they developed a rich state space duality (SSD) framework of theoretical connections between SSMs and various forms of attention.
-This framework facilitated the design of Mamba-2, an extended version of Mamba, which aims to improve its efficiency (not performance).
-Mamba-2 utilizes a scalar data-dependent gating mechanism (like the ones proposed by [71, 91, 3]), which enables transforming its recurrent structure into a matrix-multiply form, thus allowing for efficient execution on tensor cores and better support for larger hidden state sizes.
-The strategy of the SSD framework mirrors that of linear attention [48], which established a connection between autoregressive attention mechanisms and linear RNNs via showing an equivalence between "dual forms" of quadratic kernelized attention and a specific type of linear recurrence.
-Conceptually, the SSD framework seeks to transfer algorithmic and systems-level optimizations originally developed for Transformers to the realm of SSMs.
-Its overarching goal is to enable the development of architectures that outperform Transformers, while scaling more efficiently with sequence length.
-Finally, several works, including Tolstikhin et al.
-[95], Melas-Kyriazi [60], Touvron et al.
-[96], and Ding et al.
-[18], among others, have questioned the necessity of self-attention, particularly in the context of Vision Transformers.
-In contrast, Liu et al.
-[53] introduced gMLP, an MLP-based alternative to BERT-style Transformers [16] that (partially) eliminates self-attention but ultimately underperforms average performance on downstream NLP tasks.
-gMLP encompasses channel (hidden) and spatial (cross-token) projections with multiplicative gating and static parameterization.
-Its gating mechanism is reminiscent of Gated Linear Units (GLUs) [15, 84, 100], as well as earlier architectures such as Highway Networks [88] and LSTM-RNNs [34].
-A key distinction, however, is that gMLP applies gating on the spatially projected dimension and not the hidden one.
-The gated embedding-wise neural network in Avey’s contextualizer draws inspiration from gMLP.
-Unlike all previously mentioned models, Avey entirely abandons self-attention and recurrence, introducing a new architecture composed of a ranker and a dynamically parameterized neural processor.
-The ranker identifies the most relevant tokens for contextualization, while the neural processor contextualizes them data-dependently.
-This design decouples sequence length from context width, enabling efficient processing of arbitrarily long sequences without diminishing the influence of distant yet important tokens.
-At the core of Avey’s architecture is a weighted-selective-split interaction mechanism, which filters out irrelevant tokens beyond the context window and enables direct interactions only with relevant ones, thus preserving their influence irrespective of sequence length.
-In addition, Avey employs a partial-embedding bypassing technique that retains a portion of each token’s raw, distinctive features before fusing them with its contextualized ones through a neural network.
-This technique boosts the performance of Avey (as shown in Appendix K) and might help mitigate issues such as entropy collapse [107] and over-smoothing [109, 85], especially at large-scale, when the depth of the model is increased significantly.
-Appendix P Limitations
-The scope of our work is limited to textual data and does not involve other modalities such as images, audio, or genomics.
-Additionally, our evaluation of Avey is restricted to standard autoregressive language modeling, benchmarking it against popular open-source architectures using both pretraining metrics (perplexity) and zero-shot evaluations on established NLP benchmarks.
-As a result, we do not investigate Avey’s ability to construct bidirectional contextualized word representations, as done in BERT [16].
-We leave this for future work.
-Finally, the paper focuses solely on effectiveness rather than efficiency.
-While we provide a complexity analysis showing that Avey exhibits quadratic training time like Transformers, our current implementation is slower.
-As such, further engineering efforts are required to optimize it.
+Appendix E Is the Ranker a RAG Component?
+The ranker is an internal component of Avey that operates within the input sequence, selecting among its splits for more effective contextualization.
+It does not query external corpora or indexes, introduces no retrieval I/O or freshness dependencies, and adds no retrieval latency.
+Its role is architectural, that is, to allocate Avey’s internal contextual budget and decouple context width from sequence length so that Avey can fully contextualize sequences far beyond its training window.
+By contrast, Retrieval-Augmented Generation (RAG) [lewis2020rag] augments a model with external (non-parametric) knowledge via a retriever, ranging from classic BM25 [robertson2009bm25] and dense passage retrieval (DPR) [karpukhin2020dpr] to system-level designs such as REALM [guu2020realm], RETRO [borgeaud2022retro], and MacRAG [lim2025macrag], among others.
+RAG aims to: (1) improve factuality by grounding outputs in retrieved documents, (2) make models updatable by reflecting new information without retraining, and (3) reduce compute for long-context tasks by moving knowledge out of weights.
+As such, the two mechanisms are orthogonal.
+The ranker allocates the model’s internal contextual budget over the given sequence, whereas RAG changes the evidence set by importing out-of-sequence content.
+They can be composed (RAG can be layered atop Avey, as it is with the Transformer) but one does not subsume the other.
+Appendix F Neural Contextualization vs. Attention
+Avey’s contextualizer is an embedding-wise neural network that dispenses with attention.
+In Appendix 3.3, we replaced it with standard self-attention and observed a $4.6\%$ increase in perplexity alongside a $2.1\%$ decline in average task performance, underscoring its central role in Avey’s architecture.
+Formally, the contextualizer is defined in Equation 2, repeated below for convenience:
+$\mathbf{c}(\mathbf{Z}_{t})=\mathbf{Z}_{tl}\ \odot\ \sigma\!\Big(\big(\mathbf{V}\ \odot\ \mathcal{N}(\mathbf{Z}_{tr})\,\mathcal{N}(\mathbf{Z}_{tr})^{\top}\big)\,\mathbf{Z}_{tr}\ +\ \mathbf{b}^{\prime}\Big).$
+Let $\mathbf{S}:=\mathcal{N}(\mathbf{Z}_{tr})\,\mathcal{N}(\mathbf{Z}_{tr})^{\top}$.
+The product $(\mathbf{V}\!~\odot\!~\mathbf{S})\mathbf{Z}_{tr}$ yields a content-dependent signal that is passed through a pointwise nonlinearity and used to gate $\mathbf{Z}_{tl}$ elementwise, producing a bounded, feature-wise modulation rather than a mixture over values.
+By contrast, self-attention computes a row-stochastic convex combination of value vectors after a Q/K split and softmax normalization.
+Evidently, our formulation departs away from both the softmax and the Q/K/V decomposition, whereby weights are neither constrained to be nonnegative nor to sum to one, and the output acts as a gate on carrier features (i.e., $\mathbf{Z}_{tl}$) rather than a convex average of value vectors.
+The contextualizer also differs fundamentally from linear attention [choromanski2021performer, katharopoulos2020transformers, wang2020linformer, beltagy2020longformer, sun2023retentive].
+Linear-attention variants obtain near-linear complexity by exploiting an associative kernel factorization that permits reordering and prefix accumulation, typically of the form $\phi(\mathbf{Q})\big(\phi(\mathbf{K})^{\top}\mathbf{V}_{\!\text{val}}\big)$.
+Equation 2 does not admit such reordering.
+In fact, the Hadamard coupling ($\mathbf{V}\!~\odot\!~\mathbf{S}$) breaks the algebraic associativity required to push multiplications across terms, and the normalization $\mathcal{N}(\cdot)$ is neither linear nor guaranteed nonnegative, precluding the kernel tricks used to approximate softmax attention with associative feature map functions (e.g., ReLU and Exp).
+Lastly, we note that the contextualizer remains quadratic (not linear) in sequence length.
+For similar reasons, Equation 2 cannot be reformulated as a finite-state RNN under an autoregressive mask.
+Let $\mathbf{S}_{t}=\mathcal{N}(\mathbf{Z}_{tr}^{\leq t})\,\mathcal{N}(\mathbf{Z}_{tr}^{\leq t})^{\top}$.
+The update at step t+1t\!+\!1 depends on the full pairwise matrix $(\mathbf{V}\!~\odot\!~\mathbf{S}_{t})$, that is, on all position-specific interactions among the past tokens after data-dependent normalization.
+Because the learned weight matrix $\mathbf{V}$ introduces position-dependent multiplicative couplings, there is no time-invariant transition $h_{t+1}=f(h_{t},x_{t+1})$ with a fixed-dimensional sufficient statistic $h_{t}$ that exactly summarizes $(\mathbf{V}\!~\odot\!~\mathbf{S}_{t})$.
+In particular, the required weights vary across positions and must be recomputed, so any streaming recurrence would either approximate by tying/averaging $\mathbf{V}$ or maintain $\mathcal{O}(t)$ state.
+Therefore, an exact finite-state RNN equivalence is unavailable.
+Empirically, $\mathbf{V}$ performs most of the heavy-lifting in Avey, while $\mathbf{S}$ primarily induces selectivity, dynamically emphasizing or suppressing interactions conditioned on the input, echoing the selectivity principle advocated in recent sequence models [gu2023mamba].
+An ablation in Appendix 3.3 shows that including $\mathbf{S}$ delivers a consistent, albeit modest, gain by making the neural processor’s parametrization input-adaptive.
+Putting everything together, these distinctions (i.e., gating rather than mixing, non-associative pairwise modulation rather than kernel-factorizable operations, and explicit quadratic interactions), explain both the theoretical departure from self-attention and linear attention and the observed empirical contribution of the contextualizer within Avey.
+Appendix G Design Rationale
+We designed Avey around clear functional roles for its core modules. Below, we outline some of the guiding intuitions and how they inform its architecture.
+Enricher:
+A substantial body of evidence indicates that much of a language model’s knowledge is stored in feed-forward sublayers and accessed through non-linear feature interactions (e.g., [geva2021transformer_ffn_memory]).
+The enricher is designed accordingly.
+It serves both as the primary repository of parametric knowledge and as a mechanism for intra-embedding interactions, enabling higher-order, non-linear composition of features within each embedding.
+This improves expressivity by allowing features to modulate and refine one another in a context-aware manner.
+Contextualizer:
+The contextualizer operates as an embedding-wise neural network such that each neuron forms a weighted sum over input embeddings with learned coefficients (see Equation 2).
+To introduce input-dependent selectivity (as in [gu2023mamba]), we augment these static weights with a cosine-similarity term that produces a second, data-driven set of weights (the two are combined multiplicatively via a Hadamard product).
+This dynamic modulation improves behaviors such as copying and induction by strengthening interactions that are semantically relevant to the current input.
+The split-and-gate structure follows established gated designs in gMLP [liu2021pay] and GLU variants [dauphin2017language, shazeer2020glu, wu2019pay].
+Partial Embedding Bypassing:
+The enricher’s output is partitioned into two streams, one is passed to the contextualizer and the other is bypassed and fed directly to the fuser.
+The bypassed part plays two complementary roles.
+First, it provides a strong residual path that preserves signal and stabilizes optimization by improving gradient flow within each Avey layer.
+Second, it supplies additional non-linear capacity in the downstream feed-forward fuser, complementing the contextualizer’s primarily linear mixing across embeddings.
+This balance between context-aware and context-invariant processing yields richer, more diverse representations.
+Fuser:
+The fuser (a position-wise feed-forward network) learns how to combine the contextualized and bypassed streams and then projects the result back to the model’s embedding dimension, ensuring compatibility with residual pathways across layers.
+As a feed-forward network (FFN), it also contributes to storing and accessing parametric knowledge learned during training, analogous to FFN roles in Transformers.
+Appendix H Limitations
+The scope of this work is limited to textual data and does not extend to other modalities such as images or audio, nor to specialized domains like genomics.
+Our evaluation of Avey focuses on standard autoregressive language modeling, benchmarking against popular open-source architectures using both pretraining metrics (perplexity) and zero-shot performance on established NLP benchmarks.
+Consequently, we do not investigate Avey’s ability to produce bidirectional contextualized representations as in BERT [devlin2019bert].
+We leave this direction to future work.
+Finally, the paper emphasizes effectiveness rather than efficiency.
+While our complexity analysis indicates quadratic training time similar to Transformers, our current implementation is slower in practice, and additional engineering is needed to optimize it.